### v1

# and the homomorphism conjecture for $\mathcal{A}_{g}$

Abstract.
The tautological $\mathbb{Q}$-subalgebra $\mathsf{R}^{*}(\mathcal{A}_{g})\subset\mathsf{CH}^{*}(\mathcal{A}_{g})$ of the Chow ring of the moduli space of principally polarized abelian varieties is generated by the Chern classes of the Hodge bundle. A canonical $\mathbb{Q}$-linear projection operator
$\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\rightarrow\mathsf{R}^{*}(\mathcal{A}_{g})$
was constructed in [9]. We present here new calculations of intersection products of the Torelli locus in $\mathcal{A}_{g}$ with the product loci
$\mathcal{A}_{r}\times\mathcal{A}_{g-r}\rightarrow\mathcal{A}_{g}$
for $r\leq 3$. The results suggest that $\mathsf{taut}$ is a $\mathbb{Q}$-algebra homomorphism (at least for special cycles). We discuss a conjectural framework for the homomorphism property of tautological projection.
Our calculations follow two independent approaches. The first is a direct study of the excess intersection geometry of the fiber product of the Torelli and product morphisms. The second approach is by recasting the geometry in terms of families Gromov-Witten classes, which are computed by a wall-crossing formula related to unramified maps [49, 51]. Both approaches are of interest beyond their use here.
We define new tautological projections of cycles on the fiber products $\mathcal{X}_{g}^{s}\to\mathcal{A}_{g}$ of the universal family. We compute these projections for a class of product cycles on $\mathcal{X}_{g}^{s}$ in terms of a determinant involving the universal theta divisors and Poincaré classes. The Torelli pullback of the morphism $\mathcal{A}_{1}\times\mathcal{A}_{5}\to\mathcal{A}_{6}$ was used in [10] to construct a nontrivial element of the Gorenstein kernel of ${\mathsf{R}}^{5}(\mathcal{M}^{\mathrm{ct}}_{6})$. Using Abel-Jacobi pullbacks of product cycles on $\mathcal{X}_{g}^{s}$ and their projections, we construct a new family of classes which we conjecture to lie in the Gorenstein kernels of the tautological rings $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g,n})$. In particular, we construct a nontrivial element of the Gorenstein kernel of $\mathsf{R}^{5}(\mathcal{M}_{5,2}^{\mathrm{ct}})$. Whether these classes generate the Gorenstein kernels of all $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ is an intriguing open question.
1. Introduction
1.1. Tautological classes on the moduli space of abelian varieties
Let $\mathcal{A}_{g}$ be the moduli space of principally polarized abelian varieties of dimension $g\geq 1$. Points of $\mathcal{A}_{g}$ correspond to pairs $(X,\theta)$ of an abelian variety $X$ together with a polarization given by an ample line bundle $\theta$ satisfying $h^{0}(X,\mathcal{O}(\theta))=1$. The moduli space $\mathcal{A}_{g}$ is an irreducible, nonsingular Deligne-Mumford stack of dimension $\binom{g+1}{2}$ equipped with a universal family and a universal zero section,
$\pi:\mathcal{X}_{g}\rightarrow\mathcal{A}_{g}\,,\ \ \ \mathsf{z}:\mathcal{A}_{g}\rightarrow\mathcal{X}_{g}\,.$
We refer the reader to [5] for a foundational treatment of the moduli of abelian varieties.
The Hodge bundle $\mathsf{z}^{*}\Omega_{\pi}=\mathbb{E}\rightarrow\mathcal{A}_{g}$ is the rank $g$ vector bundle with fibers
$\mathbb{E}|_{[(X,\theta)]}=\mathrm{H}^{0}(X,\Omega_{X})\,.$
The Chern classes $\lambda_{i}=c_{i}(\mathbb{E})$ are naturally defined in the Chow ring [^1] All Chow and cohomology rings will be taken with $\mathbb{Q}$-coefficients. of the moduli space of abelian varieties. Over the complex numbers, $\mathcal{A}_{g}$ is isomorphic to the orbifold $\mathbb{H}_{g}/\operatorname{Sp}_{2g}(\mathbb{Z})]$, where $\mathbb{H}_{g}$ is the Siegel upper half space. Since $\mathbb{H}_{g}$ is contractible,
$\mathsf{H}^{*}(\mathcal{A}_{g})\cong\mathsf{H}^{*}(\operatorname{Sp}_{2g}(\mathbb{Z}),\mathbb{Q})\,.$
By a result of Borel [6], the rational group cohomology of $\operatorname{Sp}_{2g}(\mathbb{Z})$ stabilizes to a polynomial algebra in the odd lambda classes:
$\lim_{g\to\infty}\mathsf{H}^{*}(\operatorname{Sp}_{2g}(\mathbb{Z}),\mathbb{Q})=\mathbb{Q}[\lambda_{1},\lambda_{3},\ldots]\,.$
Motivated by Borel’s calculation, van der Geer [63] defined the tautological ring
$\mathsf{R}^{*}(\mathcal{A}_{g})\subset\mathsf{CH}^{*}(\mathcal{A}_{g})$
as the subalgebra generated by $\lambda_{1},\ldots,\lambda_{g}$. The structure of $\mathsf{R}^{*}(\mathcal{A}_{g})$ is well understood:
Theorem 1 (van der Geer [63]).
  1. (i)
    The relations $c(\mathbb{E}\oplus\mathbb{E}^{\vee})=1$ and $\lambda_{g}=0$ hold in $\mathsf{R}^{*}(\mathcal{A}_{g})$ and generate the ideal of all relations among the $\lambda$ classes.
  2. (ii)
    $\mathsf{R}^{\binom{g}{2}}(\mathcal{A}_{g})\stackrel{{\scriptstyle\sim}}{{=}}\mathbb{Q}$ and is generated by the class $\prod_{i=1}^{g-1}\lambda_{i}$.
  3. (iii)
    The intersection pairing
    $\mathsf{R}^{*}(\mathcal{A}_{g})\times\mathsf{R}^{\binom{g}{2}-*}(\mathcal{A}_{g})\rightarrow\mathsf{R}^{\binom{g}{2}}(\mathcal{A}_{g})\cong\mathbb{Q}$
    is perfect.
The presentation (i) yields an isomorphism
$\mathsf{R}^{*}(\mathcal{A}_{g})\cong\mathsf{CH}^{*}(\mathsf{LG}_{g-1})\,$,
where $\mathsf{LG}_{g-1}$} is the Lagrangian Grassmannian of (g1)(g-1)-dimensional Lagrangian subspaces of $\mathbb{C}^{2g-2}.$
1.2. Tautological projection for $\mathcal{A}_{g}$
By definition, $\mathsf{R}^{*}(\mathcal{A}_{g})$ is a $\mathbb{Q}$-linear subspace of $\mathsf{CH}^{*}(\mathcal{A}_{g})$. A canonical $\mathbb{Q}$-linear tautological projection was constructed in [9]. A review of the construction is presented here.
Given $\alpha\in\mathsf{CH}^{\binom{g}{2}}(\mathcal{A}_{g})$, consider the evaluation,
$\epsilon^{\mathrm{ab}}:\mathsf{CH}^{\binom{g}{2}}(\mathcal{A}_{g})\to\mathbb{Q}\,,\ \quad\alpha\mapsto\int_{\overline{\mathcal{A}}_{g}}\overline{\alpha}\cdot\lambda_{g}\ \,$,
where $\overline{\alpha}\in\mathsf{CH}^{\binom{g}{2}}(\overline{\mathcal{A}}_{g})$ is a lift of $\alpha$ to a nonsingular toroidal compactification,
$\mathcal{A}_{g}\subset\overline{\mathcal{A}}_{g}\,,\ \ \ \overline{\alpha}|_{\mathcal{A}_{g}}=\alpha\,.$
The evaluation $\epsilon^{\mathrm{ab}}(\alpha)$ is well-defined (independent of both the choice of the toroidal compactification and the choice of the lift) by the results of [9].
In addition to the evaluation $\epsilon^{\mathrm{ab}}$, we also have a $\lambda_{g}$-pairing,
(1) $\langle\,,\,\rangle:\,\mathsf{CH}^{*}(\mathcal{A}_{g})\times\mathsf{R}^{\binom{g}{2}-*}(\mathcal{A}_{g})\,\to\,\mathbb{Q}\,,\ \quad\langle\gamma,\delta\rangle=\epsilon^{\mathrm{ab}}(\gamma\cdot\delta)=\int_{{\overline{\mathcal{A}}_{g}}}\overline{\gamma}\cdot\overline{\delta}\cdot\lambda_{g}\,$,
The results of [9] show that the $\lambda_{g}$-pairing is independent of choices and, therefore, well-defined.
The restriction of the $\lambda_{g}$-pairing to tautological classes
$\mathsf{R}^{*}(\mathcal{A}_{g})\times\mathsf{R}^{\binom{g}{2}-*}(\mathcal{A}_{g})\to\mathsf{R}^{\binom{g}{2}}(\mathcal{A}_{g})\cong\mathbb{Q}$
is nondegenerate by Theorem 1. The map to $\mathbb{Q}$ on the right is defined by $\epsilon^{\mathrm{ab}}$ and is an isomorphism because [^2] Here, $B_{2i}$ is the Bernoulli number.
(2) $\epsilon^{\mathsf{ab}}\left(\prod_{i=1}^{g-1}\lambda_{i}\right)=\int_{\overline{\mathcal{A}}_{g}}\,\prod_{i=1}^{g}\lambda_{i}=\prod_{i=1}^{g}\frac{|B_{2i}|}{4i}\,\,$,
see [63, page 71].
Definition 2 ([9]).
Let $\gamma\in\mathsf{CH}^{*}(\mathcal{A}_{g})$. The tautological projection $\mathsf{taut}(\gamma)\in\mathsf{R}^{*}(\mathcal{A}_{g})$ is the unique tautological class satisfying
$\langle\mathsf{taut}(\gamma),\delta\rangle=\langle\gamma,\delta\rangle$
for all classes $\delta\in\mathsf{R}^{*}(\mathcal{A}_{g})$. ∎
$\bullet$ If $\gamma\in\mathsf{R}^{*}(\mathcal{A}_{g})$, then $\gamma=\mathsf{taut}(\gamma)$, so we have a $\mathbb{Q}$-linear projection operator:
$\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\to\mathsf{R}^{*}(\mathcal{A}_{g})\,,\ \ \ \mathsf{taut}\circ\mathsf{taut}=\mathsf{taut}\,.$
$\bullet$ For $\gamma\in\mathsf{CH}^{*}({\mathcal{A}}_{g})$, tautological projection provides a canonical decomposition
$\gamma=\mathsf{taut}(\gamma)+(\gamma-\mathsf{taut}(\gamma))$
into purely tautological and purely non-tautological parts.
$\bullet$ If $\gamma\in\mathsf{CH}^{*}(\mathcal{A}_{g})$ and $\delta\in\mathsf{R}^{*}(\mathcal{A}_{g})$, we have
(3) $\mathsf{taut}(\gamma\cdot\delta)=\mathsf{taut}(\gamma)\cdot\delta\,.$
1.3. Geometric cycles on $\mathcal{A}_{g}$
The geometry of abelian varieties provides a collection of interesting cycles on $\mathsf{CH}^{*}(\mathcal{A}_{g})$. We will discuss three types of geometric cycles here.
Product cycles: For a partition [^3] The parts are an ordered set of positive integers. $g=g_{1}+\ldots+g_{k}$, the product map
(4) $\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}\rightarrow\mathcal{A}_{g}\,$
is proper. The image is the set of $(X,\theta)\in\mathcal{A}_{g}$ for which there exists a decomposition
(5) $(X,\theta)\cong(X_{1},\theta_{1})\times\cdots\times(X_{k},\theta_{k})$
as principally polarized abelian varieties. Let
$\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}]\in\mathsf{CH}^{*}(\mathcal{A}_{g})$
be the pushforward of the fundamental class along the morphism (4).
The tautological projections of all product cycles in $\mathcal{A}_{g}$ have been computed in [9, Theorem 6]. The cases most relevant for our study here are:
[…]
$\begin{split}\mathsf{taut}\left(\left[\mathcal{A}_{1}\times\mathcal{A}_{1}\times\mathcal{A}_{g-2}\right]\right)&=\frac{1}{36}\cdot\frac{g(g-1)}{|B_{2g}||B_{2g-2}|}\cdot\lambda_{g-1}\lambda_{g-2}\,.\end{split}$
Noether-Lefschetz cycles: By permitting the decomposition (5) to hold up to an isogeny of polarized abelian varieties, we obtain an infinite collection of cycle classes which contain the product loci $\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}]\in\mathsf{CH}^{*}(\mathcal{A}_{g})$ as special cases. Even more generally, we can consider irreducible components of the Noether-Lefschetz locus where the abelian varieties $(A,\theta)$ have Néron-Severi rank at least $k$. We will refer to the fundamental classes of any such irreducible loci as Noether-Lefschetz cycles. A classification of Noether-Lefschetz loci with generic Néron-Severi rank 2 can be found in [14].
After the product loci, the simplest examples of Noether-Lefschetz loci are:
$\operatorname{NL}_{g,d}=\{(X,\theta)\,|\,\text{ $\exists$ an elliptic subgroup }E\subset X\text{ with }\deg(\theta|_{E})=d\}\subset\mathcal{A}_{g}\,.$
A formula [^4] The product in the formula is over all primes $p$ dividing $d$. for $\mathsf{taut}([\operatorname{NL}_{g,d}])\in\mathsf{R}^{g-1}(\mathcal{A}_{g})$ was proven by Iribar López [40, Theorem 3]:
(6) $\mathsf{taut}([\operatorname{NL}_{g,d}])=\frac{d^{2g-1}g}{6|B_{2g}|}\prod_{p|d}(1-p^{2-2g})\lambda_{g-1}\,$,
for $g\geq 2$ and $d\geq 1$. A connection to modular forms can be found in a different basis. Define
$\widetilde{\operatorname{NL}}_{g,d}]=\sum_{e|d}\sigma_{1}\left(\frac{d}{e}\right)[\operatorname{NL}_{g,e}]\ \in\ \mathsf{CH}^{g-1}(\mathcal{A}_{g})\,$,
where $\sigma_{r}(n)=\sum_{d|n}d^{r}$, see [31]. Iribar López’s result (6) can then be rewritten [^5] Equation (7) is correct also for $g=1$ with the convention $\operatorname{NL}_{1,d}]=[\mathcal{A}_{1}]$. as
(7) $\mathsf{taut}\left(\frac{(-1)^{g}}{24}\lambda_{g-1}+\sum_{d=1}^{\infty}[\widetilde{\operatorname{NL}}_{g,d}]\,q^{d}\right)=\frac{(-1)^{g}}{24}E_{2g}(q)\lambda_{g-1}\,$,
where $E_{2g}(q)$ is the Eisenstein modular function of weight $2g$ in the variable $q=e^{2\pi i\tau}$. Further connections to modularity, including conjectures by Greer, Lian, Iribar López, and Pixton, can be found in the lecture [55]. The tautological projections of $\widetilde{\mathrm{NL}}_{g,d}]$ have a surprising connection to the genus 1 Gromov-Witten theory of the Hilbert scheme of points of $\mathbb{C}^{2}$ proven in [39].
Torelli cycles: A natural source of principally polarized abelian varieties comes from the geometry of curves. A connected nodal curve $C$ is said to be of compact type if the dual graph is a stable tree. The moduli space of genus $g$ curves of compact type
$\mathcal{M}_{g}^{\mathrm{ct}}\subset\overline{\mathcal{M}}_{g}$
is a nonsingular Deligne-Mumford stack of dimension 3g33g-3. The Torelli morphism
(8) $\mathsf{Tor}:\mathcal{M}_{g}^{\mathrm{ct}}\rightarrow\mathcal{A}_{g}$
sends a compact type curve $C$ to the product of the Jacobians of the irreducible components of $C$, each equipped with the canonical principal polarization given by the corresponding theta divisor. The Torelli morphism (8) is proper, and the associated pushforward of the fundamental class is the Torelli cycle,
$\mathsf{Tor}_{*}([\mathcal{M}_{g}^{\mathrm{ct}}])=[\mathcal{J}_{g}]\in\mathsf{CH}^{\binom{g-2}{2}}(\mathcal{A}_{g})\,.$
Faber observed [^6] Faber’s calculations, undertaken before the complete definition of $\mathsf{taut}$ was available, can now be interpreted as computing $\mathsf{taut}([\mathcal{J}_{g}])$. that the calculation of $\mathsf{taut}([\mathcal{J}_{g}])\in\mathsf{R}^{\binom{g-2}{2}}(\mathcal{A}_{g})$ is equivalent to the calculation of all Hodge integrals of the form
$\int_{\overline{\mathcal{M}}_{g}}\lambda_{g}\cdot\mathsf{P}(\lambda_{1},\ldots,\lambda_{g})\,$,
where $\mathsf{P}$ is an arbitrary polynomial of degree 2g32g-3 in the Chern classes of the Hodge bundle [24].
For $g\leq 8$, we have
$\mathsf{taut}(\left[\mathcal{J}_{2}\right])=1\,$,
$\mathsf{taut}(\left[\mathcal{J}_{3}\right])=2\,$,
$\mathsf{taut}(\left[\mathcal{J}_{4}\right])=16\lambda_{1}\,$,
$\mathsf{taut}(\left[\mathcal{J}_{5}\right])=144\lambda_{1}\lambda_{2}-96\lambda_{3}\,$,
$\mathsf{taut}(\left[\mathcal{J}_{6}\right])=768\lambda_{1}\lambda_{2}\lambda_{3}-2304\lambda_{2}\lambda_{4}+\frac{948096}{691}\lambda_{1}\lambda_{5}\,$,
$\mathsf{taut}(\left[\mathcal{J}_{7}\right])=1536\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}-13824\lambda_{2}\lambda_{3}\lambda_{5}+\frac{4418304}{691}\lambda_{1}\lambda_{4}\lambda_{5}+\frac{15044352}{691}\lambda_{1}\lambda_{3}\lambda_{6}-\frac{17685504}{691}\lambda_{4}\lambda_{6}\,$,
$\mathsf{taut}(\left[\mathcal{J}_{8}\right])=$ $\ \ \ \,\frac{500106387456}{2499347}\lambda_{2}\lambda_{6}\lambda_{7}-\frac{311646117888}{2499347}\lambda_{3}\lambda_{5}\lambda_{7}-\frac{203316609024}{2499347}\lambda_{1}\lambda_{2}\lambda_{5}\lambda_{7}$
$+\frac{139564449792}{2499347}\lambda_{1}\lambda_{3}\lambda_{4}\lambda_{7}-\frac{14966784}{691}\lambda_{4}\lambda_{5}\lambda_{6}+\frac{25731072}{691}\lambda_{1}\lambda_{3}\lambda_{5}\lambda_{6}$
$-\frac{12533760}{691}\lambda_{2}\lambda_{3}\lambda_{4}\lambda_{6}+\frac{552960}{691}\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}\lambda_{5}\,.$
Formulas for $\mathsf{taut}([\mathcal{J}_{g}])$ for $g\leq 11$ can be found at [60].
Since Faber’s seminal paper [24], many closed formulas for Hodge integrals have been found (especially when the integrand includes the top Chern class $\lambda_{g}$), see [19, 20, 21]. However, at present, a closed formula for $\mathsf{taut}(\left[\mathcal{J}_{g}\right])$ is not known.
1.4. The homomorphism property
Tautological projection is constructed as a $\mathbb{Q}$-linear map. The following property was introduced in [40] to study the relationship of $\mathsf{taut}$ with the multiplicative structure of $\mathsf{CH}^{*}(\mathcal{A}_{g})$.
Definition 3 (Iribar López [40]).
Let $\alpha$, $\beta\in\mathsf{CH}^{*}(\mathcal{A}_{g})$. The pair $(\alpha,\beta)$ satisfies the homomorphism property if
(9) $\mathsf{taut}(\alpha\cdot\beta)=\mathsf{taut}(\alpha)\cdot\mathsf{taut}(\beta)\in\mathsf{R}^{*}(\mathcal{A}_{g})\,.\qed$
By equation (3), if either $\alpha\in\mathsf{R}^{*}(\mathcal{A}_{g})$ or $\beta\in\mathsf{R}^{*}(\mathcal{A}_{g})$, then the homomorphism property holds for $(\alpha,\beta)$. The homomorphism property is therefore an obstruction to being tautological.
The homomorphism property has been recently proven to hold in the following nontrivial cases:
  • (i)
    If both $\alpha$ and $\beta$ are supported on the locus in $\mathcal{A}_{g}$ of non-simple abelian varieties, then the homomorphism property holds for $(\alpha,\beta)$, see [40].
  • (ii)
    The homomorphism property holds for the pair $([\mathcal{J}_{g}],[\operatorname{NL}_{g,d}])$ for all $g,d\geq 1$ by a geometric calculation of both sides of the equality (9), see [39].
  • (iii)
    If $\beta$ is supported on the Noether-Lefschetz locus of abelian varieties that decompose (up to isogeny) as a product of $m$ elliptic curves and a (gm)(g-m)-dimensional abelian variety, then the homomorphism property holds for $(\alpha,\beta)$ for all $\alpha\in\mathsf{CH}^{*}(\mathcal{A}_{g})$, see [26]. In particular, the homomorphism property holds for $(\alpha,[\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{1}\times\mathcal{A}_{g-m}])$.
In case (ii), the homomorphism property holds even though $\mathrm{NL}_{g,d}]\in\mathsf{CH}^{*}(\mathcal{A}_{g})$ is known to be non-tautological in many cases [10, 40].
We study here the homomorphism property for pairs of the form
$(\left[\mathcal{J}_{g}\right],[\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])\,.$
The main geometric difficulty is the calculation of the pullback
(10) $\mathsf{Tor}^{*}([\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])\in\mathsf{CH}^{*}(\mathcal{M}^{\mathrm{ct}}_{g})\,.$
We use two different approaches to compute (10). The first is a direct study of the excess intersection geometry of the fiber product of the Torelli and product morphisms following [10]. The second approach is by recasting the intersection geometry in terms of families Gromov-Witten classes, which are computed by a wall-crossing formula related to unramified maps [51, 49].
1.5. The Torelli fiber product
Let $\underline{g}=(g_{1},\ldots,g_{k})$ be a partition of $g$. We study the fiber product
[…]
in Sections 2 and 3. The stack $\mathcal{Z}$ parametrizes curves of compact type with decomposable Jacobians. Our first result provides a complete description of the fiber product. For the intersection calculation (10), an understanding of all intersections of the irreducible components of $\mathcal{Z}$ is required. In case $\underline{g}=(1,g-1)$, a parallel result was proven in [10].
Theorem 4.
There is a natural stratification of $\mathcal{Z}$ by nonsingular Deligne-Mumford stacks indexed by stable trees $\mathsf{T}$ with a genus assignment
$\mathsf{g}:\mathsf{V}(\mathsf{T})\rightarrow\mathbb{Z}_{\geq 0}$
and a coloring of the vertices of positive genus
$\mathsf{c}:\mathsf{V}(\mathsf{T})_{>0}\rightarrow\{1,\ldots,k\}$
subject to the following two conditions:
  • (i)
    For each $i=1,\ldots,k$, $\sum_{v\in{\mathsf{c}}^{-1}(i)}{\mathsf{g}}(v)=g_{i}$.
  • (ii)
    For each edge $e$ of $\mathsf{T}$, there is at least one path containing $e$ with endpoints given by two vertices of positive genus of different colors and all interior vertices of genus 0.
For each $(\mathsf{T},\mathsf{g},\mathsf{c})$, there is a morphism
$\mathcal{M}_{\mathsf{T}}^{\mathsf{ct}}=\prod_{v\in\mathsf{V}(\mathsf{T})}\mathcal{M}_{\mathsf{g}(v),\mathsf{n}(v)}^{\mathrm{ct}}\,\to\mathcal{Z}.$
If $\mathsf{T}$ has no vertices of genus 0, then $\mathcal{M}_{\mathsf{T}}^{\mathsf{ct}}$ corresponds to an irreducible component of $\mathcal{Z}$. Trees with genus 0 vertices correspond to strata of the intersections of the irreducible components.
We will refer to the trees in Theorem 4 as $\underline{g}$-colored extremal trees. Let $\mathsf{Tree}(\underline{g})$ be the set of $\underline{g}$-colored extremal trees. An automorphism of $\mathsf{T}\in\mathsf{Tree}(\underline{g})$ is an automorphism of the underlying tree that respects both the genus assignment and the coloring. Let $\mathsf{Aut}(\mathsf{T})$ be the automorphism group.
Theorem 5 (Excess contributions).
We have
$\mathsf{Tor}^{*}([\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])=\sum_{\mathsf{T}\in\mathsf{Tree}(\underline{g})}\frac{1}{|\mathsf{Aut}(\mathsf{T})|}\left(\xi_{\mathsf{T}}\right)_{*}(\mathsf{Cont}_{\mathsf{T}})\,\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}}),$
where $\mathsf{Cont}_{\mathsf{T}}\in\mathsf{R}^{*}(\mathcal{M}_{\mathsf{T}}^{\mathsf{ct}})$ is a recursively computed tautological class, and
$\xi_{\mathsf{T}}:\mathcal{M}_{\mathsf{T}}^{\mathsf{ct}}\rightarrow\mathcal{M}_{g}^{\mathrm{ct}}$
is the gluing map.
The definitions of the tautological rings $\mathsf{R}^{*}(\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}})$ and $\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$ will be reviewed in Section 3.2. Theorem 5 extends [10, Theorem 3] from the case $\underline{g}=(1,g-1)$ to arbitrary partitions $\underline{g}$.
A key step in the proof of Theorem 5 is the reducedness of the fiber product $\mathcal{Z}$ proven by Drakengren [18]. Reducedness is crucial here: we can then write local equations for $\mathcal{Z}$ and compute the pullback class via excess intersection theory [29] on the strata of $\mathcal{M}_{g}^{\mathrm{ct}}$. The result is a graph sum formula expressing the pullback in terms of tautological classes on the moduli of curves.
Reducedness of $\mathcal{Z}$ is a fortunate property. For example, Drakengren [17] has proven that the Torelli square in genus $4$,
[…]
has a nonreduced fiber product $\mathcal{F}$.
1.6. Wall-crossing
The wall-crossing method can be used to calculate
$\mathsf{Tor}^{*}([\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])\in\mathsf{CH}^{*}(\mathcal{M}^{\mathrm{ct}}_{g})\,.$
The calculation of $\mathsf{Tor}^{*}([\mathcal{A}_{1}\times\mathcal{A}_{g-1}])$ by wall-crossing was carried out in [49]. The wall-crossing geometry pursued here for $\mathsf{Tor}^{*}([\mathcal{A}_{r}\times\mathcal{A}_{g-r}])$ is more subtle.
Let $1\leq r\leq g-1$, and let
$\pi_{r}:\mathcal{X}_{r}\rightarrow\mathcal{A}_{r}$
be the universal family of principally polarized abelian $r$-folds. Let
$\beta_{\mathrm{min}}=\frac{\theta^{r-1}}{(r-1)!}$
be the minimal class, and let
$\mathcal{M}_{g}^{\mathrm{ct}}(\pi_{r})=\mathcal{M}^{\mathrm{ct}}_{g}(\pi_{r},\beta_{\text{min}})/\mathcal{X}_{r}$
be the space of stable maps to the fibers of $\pi_{r}$ whose domain is a curve of compact type and whose image lies in the homology class $\beta_{\text{min}}$, up to translations of the fibers.
The Jacobian of a curve $C$ of compact type splits as a product of principally polarized abelian varieties of dimension $r$ and grg-r,
$\mathsf{Jac(C)}={X}_{r}\times{Y}_{g-r}\,$,
if and only if there is a stable map $C\to X_{r}$ that represents the minimal class. The correspondence extends to families, so we obtain an isomorphism of stacks
[…]
When $r=1$, the above isomorphism is explained in [10, Proposition 21]. For higher $r$, the argument is parallel.
The space of stable maps $\mathcal{M}^{\mathrm{ct}}_{g}(\pi_{r})$ carries a 2-term perfect obstruction theory and a virtual fundamental class. Since an $r$-dimensional abelian variety has $h^{0,2}=\binom{r}{2}$, the virtual fundamental class can be reduced $h^{0,2}$ times, yielding a reduced [^7] See [7] for a study of the reduced virtual fundamental class of the moduli space of stable maps to an abelian variety. virtual fundamental class $\mathcal{M}^{\mathrm{ct}}_{g}(\pi_{r})]^{\text{red}}$. By a result of [26] extending [31, Theorem 1.2], there is an equality of cycles
(12) $\mathsf{Tor}^{!}[\mathcal{A}_{r}\times\mathcal{A}_{g-r}]=[\mathcal{M}^{\mathrm{ct}}_{g}(\pi_{r})]^{\mathrm{red}}\,\in\mathsf{CH}^{*}(\mathcal{M}_{g}^{\mathrm{ct}}(\pi_{r}))$
under the isomorphism (11).
The wall-crossing formula established in [51, Theorem 6.21] relates the virtual fundamental classes of moduli spaces of stable and unramified maps. Let
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r})=\mathcal{M}_{g_{0}}^{\mathrm{ct},\mathrm{un}}(\pi_{r},\beta_{\mathrm{min}})/\mathcal{X}_{r}\,$
be the analogously defined space of unramified maps to the fibers of $\pi_{r}$. More generally, for the statement of the wall-crossing formula, spaces of unramified maps from possibly disconnected curves and with fixed multiplicities over moving points on the target are required [^8] We use the superscript $\circ$ for connected domains and the superscript $\bullet$ for possibly disconnected domains. If no superscript is given, then the domains are understood to be connected. ,
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet},$
where $\mu^{i}=(\mu^{i}_{1},\ldots,\mu^{i}_{\ell})$ are the partitions that encode the multiplicities. Unramified maps and unramified maps with prescribed ramifications are higher dimensional generalizations of admissible covers, see [44, 51]. The precise definitions are reviewed in Section 4, and an expository treatment can be found in [54].
Over the moduli space of unramified maps, before taking the quotient by translations of the target, derivatives of maps at points associated with multiplicity profiles $\mu^{i}$ provide evaluation morphisms to $\mathbb{P}(T_{\mathcal{X}_{r}/\mathcal{A}_{r}})^{\ell(\mu^{i})}$, see [44] for details. After taking the quotient, we obtain evaluation morphisms to $\mathbb{P}(\mathbb{E}^{\vee}_{r})$, the tangent space at the zero section $\mathsf{z}$:
$\mathrm{ev}_{i}\colon\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet}\rightarrow\mathbb{P}(\mathbb{E}^{\vee}_{r})^{\ell(\mu^{i})},\quad i=1,\ldots,m\,.$
The classes responsible for the difference between the enumerative geometry of stable and unramified maps live on the moduli spaces
$\mathcal{M}^{\mathrm{ct}}_{h,\,\ell(\mu)}\times\mathbb{P}(\mathbb{E}^{\vee}_{r})^{\ell(\mu)}\,$,
and take the form
(13) $I_{h,\mu}(z)=\left.\frac{1}{|\operatorname{{\mathsf{Aut}}}(\mu)|}\,\frac{\prod^{r}_{j=1}\Lambda^{\vee}(z-\alpha_{j})}{\prod^{\ell(\mu)}_{k=1}(z-\psi_{k}-H_{k})}\right|_{z^{\geq 0}}\in\mathsf{CH}^{*}\left(\mathcal{M}^{\mathrm{ct}}_{h,\,\ell(\mu)}\times\mathbb{P}(\mathbb{E}^{\vee}_{r})^{\ell(\mu)}\right)[z]\,$,
where $\Lambda^{\vee}(t)=\sum_{i=0}^{h}c_{i}(\mathbb{E}^{\vee}_{h})t^{h-i},$ $\alpha_{j}$ are Chern roots of $\mathbb{E}_{r}$, $H_{k}$ are the hyperplane classes on the factors of $\mathbb{P}(\mathbb{E}^{\vee}_{r})^{\ell(\mu)}$, and $\psi_{k}$ are the cotangent classes on $\mathcal{M}^{\mathrm{ct}}_{h,\,\ell(\mu)}$. The following result is established in [26].
Theorem 6 (Wall-crossing).
For $1\leq r\leq g-1$, we have
$\mathcal{M}^{\mathrm{ct}}_{g}(\pi_{r})]^{\textnormal{red}}=\sum_{\mathsf{S}\in\mathsf{Star}(g)}\frac{1}{|\operatorname{{\mathsf{Aut}}}(\mathsf{S})|}\,(\xi_{\mathsf{S}})_{*}\left([\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet}]^{\textnormal{red}}\boxtimes\prod^{m}_{i=1}\mathrm{ev}^{*}_{i}I_{g_{i},\mu^{i}}(-\Psi_{i})\right),$
where the sum is taken over star-shaped graphs $\mathsf{S}$ with partition labels $\mu^{1},\ldots,\mu^{m}$ on the edges that correspond to connected domains after gluing and satisfy
$g=\sum^{m}_{i=1}(g_{i}+\ell(\mu^{i}))+g_{0}-m\,.$
The $\Psi_{i}$ are the cotangent classes on $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet}$ associated to the markings of the target. The morphism
$\xi_{\mathsf{S}}\colon\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet}\times\prod^{m}_{i=1}\mathcal{M}^{\mathrm{ct}}_{g_{i},\ell(\mu^{i})}\rightarrow\mathcal{M}^{\mathrm{ct}}_{g}(\pi_{r})$
is given by applying the forgetful morphism
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet}\rightarrow\mathcal{M}^{\mathrm{ct}}_{g_{0},\sum\ell(\mu^{i})}(\pi_{r})^{\bullet}$
followed by the gluing morphism which attaches contracted components.
A formal definition of star-shaped graphs $\mathsf{S}$ and the cotangent line classes $\Psi_{i}$ will be given in Section 4. The complexity of the moduli spaces of unramified maps which appear in the wall-crossing formula depends upon $r$:
$\bullet$ For $r=1$, the only moduli spaces of unramified maps which occur in the wall-crossing formula are
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{1}(\pi_{1},(1),\ldots,(1))^{\circ},$
where $\mu^{i}=(1)$ for all $i$.
$\bullet$ For $r=2$, the only moduli spaces of unramified maps which occur are
(14) $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{2}(\pi_{2},(1),\ldots,(1))^{\circ}\ \ \ \text{and}\ \ \ \mathcal{M}^{\mathrm{ct},\mathrm{un}}_{1}(\pi_{2},(1,1),(1),\ldots,(1))^{\bullet}\,$,
The partition $\mu=(1,1)$ arises for unramified maps to fibers of $\pi_{2}$ given by products $E_{1}\times E_{2}$ of elliptic curves: the associated domains curves are two disconnected genus 1 curves mapping to a nodal genus 2 curve consisting of two elliptic components. The point on the target corresponding to the node of the genus 2 curve has multiplicity $(1,1)$.
A result of [26] exhibits the spaces (14) as an iterated blowup of $\mathcal{M}^{\mathrm{ct}}_{2,n}$ and $\mathcal{M}^{\mathrm{ct}}_{1,n}$ along boundary strata in both the connected and disconnected cases. As a result, the moduli space is nonsingular of the expected dimension, and the reduced virtual fundamental class agrees with the fundamental class. Therefore, Theorem 6 provides an explicit method to evaluate the wall-crossing formula in terms of tautological classes on the moduli spaces of curves. A full analysis is provided in Section 4 with examples.
$\bullet$ For $r\geq 3$, only partitions of the type $\mu=(1,\ldots,1)$ can occur since the minimal curves can have at worst ordinary multiple point singularities. The analysis for $r\geq 3$ is more complex and will not be treated here.
1.7. Results about the homomorphism property
The codimension of $\mathcal{J}_{g}\subset\mathcal{A}_{g}$ is
$\mathsf{cod}_{\mathcal{J}}=\binom{g+1}{2}-(3g-3)\,.$
The codimension of the product cycle $\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}$ in $\mathcal{A}_{g}$ is
$\mathsf{cod}_{\underline{g}}=\binom{g+1}{2}-\sum_{i=1}^{k}\binom{g_{i}+1}{2}=\sum_{1\leq i<j\leq k}g_{i}g_{j}\,.$
The homomorphism property
(15) $\mathsf{taut}([\mathcal{J}_{g}]\cdot[\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}])\stackrel{{\scriptstyle?}}{{=}}\mathsf{taut}([\mathcal{J}_{g}])\cdot\mathsf{taut}([\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}])\,\in\,\mathsf{R}^{\mathsf{cod}_{\mathcal{J}}+\mathsf{cod}_{\underline{g}}}(\mathcal{A}_{g})$
is trivial (and true) if
(16) $\mathsf{cod}_{\mathcal{J}}+\mathsf{cod}_{\underline{g}}>\binom{g}{2}\,$.
Simple arithmetic then shows that the validity of the homomorphism property (15) is a non-trivial question only in the following cases:
  • (i)
    $\mathcal{A}_{1}\times\mathcal{A}_{g-1}\rightarrow\mathcal{A}_{g}$ ,
  • (ii)
    $\mathcal{A}_{2}\times\mathcal{A}_{g-2}\rightarrow\mathcal{A}_{g}$ ,
  • (iii)
    $\mathcal{A}_{3}\times\mathcal{A}_{3}\rightarrow\mathcal{A}_{6}$ ,
  • (iv)
    $\mathcal{A}_{1}\times\mathcal{A}_{1}\times\mathcal{A}_{g-2}\rightarrow\mathcal{A}_{g}$ .
Case (i) was affirmatively answered by combining [10, Theorem 4] and [40, Theorem 7], while case (iv) follows from the results of [26]. Theorems 5 and 6 are suitable to cover cases (ii) and (iii).
We have implemented the computations described in Sections 1.5 and 1.6 in the programs [37, 27] which can be evaluated with the software $\mathsf{admcycles}$ [15] designed for calculations in the tautological ring of the moduli spaces of curves. The excess intersection calculations are discussed in Sections 3.53.6, and the wall-crossing calculations are discussed in Sections 4.94.11. The outcome is the verification of the homomorphism property in the following new cases.
Theorem 7.
The homomorphism property holds for the pairs:
  • (ii)
    $([\mathcal{J}_{g}],\,[\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ on $\mathcal{A}_{g}$ for all $2\leq g\leq 8\,$,
  • (iii)
    $([\mathcal{J}_{6}],\,[\mathcal{A}_{3}\times\mathcal{A}_{3}])$ on $\mathcal{A}_{6}\,.$
The only difficulty in extending the results (ii) to higher $g$ is computational complexity. We expect the homomorphism property (ii) will hold for all $g\geq 2$.
For partitions $\underline{g}=(g_{1},\ldots,g_{k})$ that do not fall into cases (i) - (iv), inequality (16) yields
$\mathsf{cod}_{\underline{g}}>\binom{g}{2}-\mathsf{cod}_{\mathcal{J}}=2g-3\,.$
By either Theorem 5 or Theorem 6, the Torelli pullbacks of product classes lie in the tautological ring
$\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})\subset\mathsf{CH}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})\,.$
The vanishing $\mathsf{R}^{>2g-3}(\mathcal{M}_{g}^{\mathrm{ct}})=0$ of [22, 30] then implies that
$\mathsf{Tor}^{*}([\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])=0\,$,
unless $(g_{1},\ldots,g_{k})$ is one of the partitions listed in (i) - (iv). For instance, for $\mathcal{A}_{3}\times\mathcal{A}_{g-3}\rightarrow\mathcal{A}_{g}$, we have the following result related to case (iii):
$\mathsf{Tor}^{*}([\mathcal{A}_{3}\times\mathcal{A}_{g-3}])=0\in\mathsf{CH}^{3g-9}(\mathcal{M}_{g}^{\mathrm{ct}})$
for $g\geq 7$. After pushing forward under $\mathsf{Tor}$, we obtain the vanishing:
$\mathcal{J}_{g}]\cdot[\mathcal{A}_{3}\times\mathcal{A}_{g-3}]=0\in\mathsf{CH}^{\binom{g+1}{2}-6}(\mathcal{A}_{g}),\quad g\geq 7\,.$
The same reasoning also yields the following sporadic vanishings in $\mathsf{CH}^{*}(\mathcal{A}_{g})$:
[…]
1.8. Conjectures
The widest interpretation of our calculations is that the homomorphism property holds in general.
Conjecture 8 (Strong form).
The $\mathbb{Q}$-linear map $\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\rightarrow\mathsf{R}^{*}(\mathcal{A}_{g})$ is a homomorphism of $\mathbb{Q}$-algebras.
An explanation for Conjecture 8 could be the existence of a compact subvariety
$\iota:V\hookrightarrow\mathcal{A}_{g}$
of codimension $g$ for which $\mathsf{H}^{*}(V)$ is generated by pullbacks of tautological classes on $\mathcal{A}_{g}$. By the ampleness of $\lambda_{1}$ and the Gorenstein property of $\mathsf{R}^{*}(\mathcal{A}_{g})$, we would then have
$\mathsf{H}^{*}(V)\stackrel{{\scriptstyle\sim}}{{=}}\mathsf{R}^{*}(\mathcal{A}_{g})$
as $\mathbb{Q}$-algebras. Moreover, $\mathsf{taut}$ would then be realized as the pullback
$\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\stackrel{{\scriptstyle\iota^{*}}}{{\rightarrow}}\mathsf{H}^{*}(V)\stackrel{{\scriptstyle\sim}}{{=}}\mathsf{R}^{*}(\mathcal{A}_{g})\,$
and, therefore, would be a homomorphism of $\mathbb{Q}$-algebras.
For $g\geq 3$, there are no such subvarieties $V\subset\mathcal{A}_{g}$ over the complex numbers [43]. However, in characteristic $p$, the locus $V_{0}\subset\mathcal{A}_{g}$ of abelian varieties with $p$-rank 0 provides a compact subvariety of the correct codimension (which represents a multiple of $\lambda_{g}$), see [52, 63]. Unfortunately, the Chow groups of $V_{0}$ have not yet been determined. Nevertheless, the existence of $V_{0}$ in characteristic $p$ may be viewed as motivation for the following weakening of Conjecture 8:
Conjecture 9 ($\overline{\mathbb{Q}}$ form).
The homomorphism property holds for all classes $\alpha,\beta\in\mathsf{CH}^{*}(\mathcal{A}_{g})$ defined over $\overline{\mathbb{Q}}$.
Both Conjectures 8 and 9 are rather bold interpretations of the current data. Another possibility is that the homomorphism property holds in the restricted case where at least one of the classes $\alpha$ or $\beta$ is a Noether-Lefschetz cycle or, more generally, a special cycle  [^9] See [47] for a discussion of special cycles in the context of $\mathcal{A}_{g}$..
1.9. Stable curves and toroidal compactifications of $\mathcal{A}_{g}$
Let $\mathcal{A}_{g}\subset\overline{\mathcal{A}}_{g}$ be a toroidal compactification to which the Torelli map can be lifted:
[…]
The two standard examples are the perfect cone compactification and the second Voronoi compactification, see [2] for the definitions. A natural question is to understand the Torelli fiber product
[…]
where, in the lower left corner, a compactification of the product that maps to $\overline{\mathcal{A}}_{g}$ is taken. In the case of the perfect cone compactification, the product of the compactifications maps to $\overline{\mathcal{A}}_{g}$, as proven in [61], so we can take
$\overline{\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}}=\overline{\mathcal{A}}_{g_{1}}\times\cdots\times\overline{\mathcal{A}}_{g_{k}}\,.$
If we further assume that the image $\mathsf{Tor}(\overline{\mathcal{M}}_{g})$ lies in the nonsingular locus of $\overline{\mathcal{A}}_{g}$, a property that holds for both perfect cone and second Voronoi [46, Section 5.2], then the following question is well-defined.
Question 10.
Compute $\mathsf{Tor}^{*}([\,\overline{\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}}\,])\in\mathsf{CH}^{*}(\overline{\mathcal{M}}_{g})\,.$
At the moment, Question 10 is open. In principle, both excess intersection theory and the wall-crossing approach can be applied. For the former, the scheme theoretic structure of the fiber product (17) over the divisor $\Delta_{0}\subset\overline{\mathcal{M}}_{g}$ would have to be determined. The wall-crossing approach has an advantage in that the families Gromov-Witten theory is well-defined for the family
$\overline{\pi}_{r}:\overline{\mathcal{X}}_{r}\rightarrow\overline{\mathcal{A}}_{r}\ ,$
since $\overline{\pi}_{r}$ is log smooth. Wall-crossing then provides a formula for the families Gromov-Witten class in terms of simpler unramified contributions. However, the comparison (12) of the families Gromov-Witten class to the intersection theory is not immediate over the compactification. If we restrict our attention to the partial compactification of $\mathcal{A}_{g}$ determined by the torus rank 1 locus [^10] See [48] for the definitions. $\mathcal{A}_{g}^{\leq 1}\subset\overline{\mathcal{A}}_{g}$, we obtain the fiber diagram
[…]
where $\mathcal{M}_{g}^{\leq 1}$ denotes the moduli space of stable curves with at most 1 loop in the dual graph. The wall-crossing method [26] provides a complete calculation of
$\mathsf{Tor}^{*}\left(\left[\,\overline{\mathcal{A}_{1}\times\mathcal{A}_{g-1}}\big|_{\mathcal{A}^{\leq 1}_{g}}\,\right]\right)\in\mathsf{R}^{*}({\mathcal{M}}^{\leq 1}_{g})\,.$
In [10], the kernel in $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{6})$ of the $\lambda_{6}$-pairing was constructed geometrically by the class
$\mathsf{Tor}^{*}\Big([\mathcal{A}_{1}\times\mathcal{A}_{5}]-\mathsf{taut}([\mathcal{A}_{1}\times\mathcal{A}_{5}])\Big)\in\mathsf{R}^{5}(\mathcal{M}_{6}^{\mathrm{ct}})\,.$
Using the geometry of the torus rank $1$ locus $\mathcal{A}_{6}^{\leq 1}\subset\overline{\mathcal{A}}_{6}$, a parallel construction holds for the 1-dimensional kernel of the $\lambda_{5}$-pairing on $\mathsf{R}^{*}(\mathcal{M}_{5,2}^{\mathrm{ct}})$.
Theorem 11 (Feusi-Iribar López-Nesterov [26]).
Let $\iota:\mathcal{M}_{5,2}^{\mathrm{ct}}\rightarrow\mathcal{M}_{6}^{\leq 1}$ be the boundary map. Then, the class [^11] The definition of the class $\mathsf{taut}^{\leq 1}\left([\overline{\mathcal{A}_{1}\times\mathcal{A}_{g-1}}\big|_{\mathcal{A}_{g}^{\leq 1}}]\right)\in\mathsf{CH}^{g-1}(\mathcal{A}_{g}^{\leq 1})$ extending $\mathsf{taut}([\mathcal{A}_{1}\times\mathcal{A}_{g-1}])\in\mathsf{CH}^{g-1}(\mathcal{A}_{g})$ over the boundary is given in [26].
$\iota^{*}{\mathsf{Tor}}^{*}\left(\left[\overline{\mathcal{A}_{1}\times\mathcal{A}_{5}}|_{\mathcal{A}^{\leq 1}_{6}}\right]-\mathsf{taut}^{\leq 1}\left(\left[\overline{\mathcal{A}_{1}\times\mathcal{A}_{5}}|_{\mathcal{A}^{\leq 1}_{6}}\right]\right)\right)\in\mathsf{R}^{5}(\mathcal{M}_{5,2}^{\mathsf{ct}})$
is the generator of the 1-dimensional kernel of the $\lambda_{5}$-pairing on $\mathsf{R}^{*}(\mathcal{M}_{5,2}^{\mathrm{ct}})$.
A different approach to the kernel of the $\lambda_{5}$-pairing on $\mathsf{R}^{*}(\mathcal{M}_{5,2}^{\mathrm{ct}})$ via the geometry of the universal family $\mathcal{X}_{5}\rightarrow\mathcal{A}_{5}$ will be presented in Section 1.10.
1.10. The geometry of the universal fiber product $\mathcal{X}^{s}_{g}$
A difficulty in studying Question 10 and extending Theorem 11 to higher torus rank loci is the complexity of the boundary geometry of toroidal compactifications $\mathcal{A}_{g}\subset\overline{\mathcal{A}}_{g}$. To generate elements of the Gorenstein kernel of $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g-s,2s})$ for general $s$, a different approach can be pursued using the $s$-fold fiber product
$\pi^{s}_{g}:\mathcal{X}^{s}_{g}\rightarrow\mathcal{A}_{g}\,$
of the universal abelian variety $\pi_{g}:\mathcal{X}_{g}\rightarrow\mathcal{A}_{g}$. The universal fiber product geometry stays over $\mathcal{A}_{g}$, so the intricacies of the boundary of $\overline{\mathcal{A}}_{g}$ are avoided.
There is a natural proper morphism $\mathcal{A}_{1}\times\mathcal{X}_{g-1}\rightarrow\mathcal{X}_{g}$ defined [^12] We have dropped the principal polarizations here for notational convenience. by
$({E},(Y,p\in Y))\mapsto({E}\times Y,(0,p)\in{E}\times Y)\,.$
Hence, we also have a morphism $\mathcal{A}_{1}\times\mathcal{X}_{g-1}\rightarrow\mathcal{A}_{g}$ which factors through
$\mathcal{A}_{1}\times\mathcal{A}_{g-1}\rightarrow\mathcal{A}_{g}\,.$
Definition 12.
The generalized product locus is
(18) $\operatorname{PR}_{g,s}=\underbrace{(\mathcal{A}_{1}\times\mathcal{X}_{g-1})\times_{\mathcal{A}_{g}}\cdots\times_{\mathcal{A}_{g}}(\mathcal{A}_{1}\times\mathcal{X}_{g-1})}_{s}\rightarrow\mathcal{X}_{g}^{s}\,.\qed$
An immediate property of the definition is the factorization
$\operatorname{PR}_{g,s}\rightarrow\mathcal{X}^{s}_{g}|_{\mathcal{A}_{1}\times\mathcal{A}_{g-1}}\rightarrow\mathcal{X}_{g}^{s}\,$,
where both morphisms are proper. Let
$\operatorname{PR}_{g,s}]\in\mathsf{CH}^{g-1+s}(\mathcal{X}_{g}^{s})$
denote the pushforward of the fundamental class of $\operatorname{PR}_{g,s}$. If $s=0$, $\mathcal{X}_{g}^{0}=\mathcal{A}_{g}$ and
$\operatorname{PR}_{g,0}]=[\mathcal{A}_{1}\times\mathcal{A}_{g-1}]\in\mathsf{CH}^{g-1}(\mathcal{A}_{g})\,.$
The algebra of tautological classes $\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\subset\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})$ is generated over $\mathsf{R}^{*}(\mathcal{A}_{g})$ by the classes
$\big\{\,\theta_{i}\in\mathsf{CH}^{1}(\mathcal{X}_{g}^{s})\,\big\}_{1\leq i\leq s}\ \ \ \text{and}\ \ \ \big\{\,\eta_{ij}\in\mathsf{CH}^{1}(\mathcal{X}_{g}^{s})\,\big\}_{1\leq i<j\leq s}\,$,
where $\theta_{i}$ is the pullback of the universal $\theta$ class [^13] The $\theta$ class is symmetric and normalized: the restriction of $\theta$ to the zero section of $\pi:\mathcal{X}_{g}\rightarrow\mathcal{A}_{g}$ is required to vanish. from the $i^{th}$ factor $\mathcal{X}_{g}\rightarrow\mathcal{A}_{g}$ of the universal fiber product and $\eta_{ij}$ is the pullback of the first Chern class of the universal Poincaré bundle [^14] The Poincaré bundle is normalized by the standard convention: the restriction is required to be trivial on the two zero sections. By symmetry, $\eta_{ij}=\eta_{ji}$, see Section 5. from the factors $i$ and $j$ of the universal fiber product. A projection operator
$\mathsf{taut}^{s}:\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})\rightarrow\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$
is defined in Section 5.4, and the following projection is computed in Section 5.6.
Theorem 13.
Let $\kappa_{g,s}=\prod_{k=0}^{s-1}(g+\frac{k}{2})$. Then,
$\mathsf{taut}^{s}([\operatorname{PR}_{g,s}])=\frac{g}{6\kappa_{g,s}|B_{2g}|}\,\det\begin{pmatrix}\theta_{1}&\eta_{12}/2&\ldots&\eta_{1s}/2\\ \eta_{12}/2&\theta_{2}&\ldots&\eta_{2s}/2\\ \vdots&\vdots&\ddots&\vdots\\ \eta_{1s}/2&\eta_{2s}/2&\ldots&\theta_{s}\end{pmatrix}\cdot\lambda_{g-1}\,\in\mathsf{R}^{*}(\mathcal{X}^{s}_{g})\,.$
Definition 14.
The Abel-Jacobi map $\operatorname{aj}:\mathcal{M}_{g,2s}^{\mathrm{ct}}\rightarrow\mathcal{X}_{g}^{s}$ is defined by
(19) $\operatorname{aj}(C,p_{1},\ldots,p_{2s})=(\mathsf{Jac}(C),\mathcal{O}_{C}(p_{1}-p_{2}),\mathcal{O}_{C}(p_{3}-p_{4}),\ldots,\mathcal{O}_{C}(p_{2s-1}-p_{2s}))\,.\qed$
The Abel-Jacobi pullback of the class
$\Delta_{g,s}=[\operatorname{PR}_{g,s}]-\mathsf{taut}^{s}([\operatorname{PR}_{g,s}])\in\mathsf{CH}^{g-1+s}(\mathcal{X}_{g}^{s})$
is proven in Section 6.3 to be tautological,
$\operatorname{aj}^{*}(\Delta_{g,s})\in\mathsf{R}^{g-1+s}(\mathcal{M}^{\mathrm{ct}}_{g,2s})\,.$
Definition 15.
The Gorenstein kernel, $\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g,n})$, is the kernel of the $\lambda_{g}$-pairing,
(20) $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\times\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\rightarrow\mathbb{Q}\,,\ \ \ \langle\alpha,\beta\rangle=\int_{\overline{\mathcal{M}}_{g,n}}\overline{\alpha}\cdot\overline{\beta}\cdot\lambda_{g}\,.\qed$
Conjecture 16.
Let $g\geq 1$ and $s\geq 0$. The class
$\operatorname{aj}^{*}(\Delta_{g,s})\in\mathsf{R}^{g-1+s}(\mathcal{M}^{\mathrm{ct}}_{g,2s})$
lies in the Gorenstein kernel of $\mathsf{R}^{g-1+s}(\mathcal{M}^{\mathrm{ct}}_{g,2s})$ and is nonzero for all but finitely many pairs of the form $(g\geq 2,s)$.
For $s=1$, the classes $\operatorname{aj}^{*}(\Delta_{g-1,1})\in\mathsf{R}^{g-1}(\mathcal{M}_{g-1,2}^{\mathsf{ct}})$ and
$\iota^{*}{\mathsf{Tor}}^{*}\left(\left[\overline{\mathcal{A}_{1}\times\mathcal{A}_{g-1}}|_{\mathcal{A}^{\leq 1}_{g}}\right]-\mathsf{taut}^{\leq 1}\left(\left[\overline{\mathcal{A}_{1}\times\mathcal{A}_{{g-1}}}|_{\mathcal{A}^{\leq 1}_{g}}\right]\right)\right)\in\mathsf{R}^{g-1}(\mathcal{M}_{g-1,2}^{\mathsf{ct}})$
are proven to be equal in [26]. Independent of the above equality and Theorem 11, we can calculate $\operatorname{aj}^{*}(\Delta_{5,1})$ directly to obtain:
Theorem 11.
The class $\operatorname{aj}^{*}(\Delta_{5,1})\in\mathsf{R}^{5}(\mathcal{M}_{5,2}^{\mathrm{ct}})$ is the generator of the 1-dimensional kernel of the $\lambda_{5}$-pairing on $\mathsf{R}^{*}(\mathcal{M}_{5,2}^{\mathrm{ct}})$.
The construction of $\operatorname{aj}^{*}(\Delta_{g,s})$ for $s>1$ may be viewed as probing the higher torus rank loci of toroidal compactifications. That $\operatorname{aj}^{*}(\Delta_{g,s})$ lies in the Gorenstein kernel was proven for $s=0$ in [10] and is proven for $s=1$ in Theorem 47, Section 6. The non-vanishing part of Conjecture 16 is open for all $s\geq 0$.
1.11. Gorenstein kernel
The system of Gorenstein kernels $\big\{\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\big\}_{g,n}$ satisfies several properties that follow directly from the definitions:
  1. (i)
    $\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g,n})$ is an ideal,
  2. (ii)
    $\pi_{i*}:\mathsf{K}_{g,n}\rightarrow\mathsf{K}_{g,n-1}$, where $\pi_{i}:\mathcal{M}_{g,n}^{\mathrm{ct}}\rightarrow\mathcal{M}_{g,n-1}^{\mathrm{ct}}$ is the map forgetting the $i^{th}$ marking,
  3. (iii)
    $\pi^{*}_{i}:\mathsf{K}_{g,n-1}\rightarrow\mathsf{K}_{g,n}$,
  4. (iv)
    $\xi_{*}:{\mathsf{K}}_{g_{1},n_{1}+1}\otimes\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1})\rightarrow{\mathsf{K}}_{g_{1}+g_{2},n_{1}+n_{2}}$, where
    $\xi:\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1}\times\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1}\rightarrow\mathcal{M}^{\mathrm{ct}}_{g_{1}+g_{2},n_{1}+n_{2}}$
    is the gluing map.
To state the compatibility with respect to pullback along the gluing map $\xi$, we first define the correspondence map
$\eta:\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1}+g_{2},n_{1}+n_{2}})\otimes\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})\,\,,\ \ \ \eta(\alpha\otimes\beta)=\pi_{1*}\left(\xi^{*}(\alpha)\cup\pi_{2}^{*}(\beta)\right)\,$,
where $\pi_{i}:\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1}\times\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1}\rightarrow\mathcal{M}^{\mathrm{ct}}_{g_{i},n_{i}+1}$ are the projections and
$\pi_{1*}:\mathsf{R}^{*}(\mathcal{M}_{g_{1},n_{1}+1}^{\mathrm{ct}})\times\mathsf{R}^{*}(\mathcal{M}_{g_{2},n_{2}+1}^{\mathrm{ct}})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})$
is defined by the $\lambda_{g_{2}}$-evaluation on the second factor. Then, we have:
  1. (v)
    $\eta:\mathsf{K}_{g_{1}+g_{2},n_{1}+n_{2}}\otimes\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1})\rightarrow{\mathsf{K}}_{g_{1},n_{1}+1}$.
Conjecture 16 motivates a natural question.
Question 17.
Is $\big\{\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\big\}_{g,n}$ the smallest system of ideals that satisfies properties (i)-(v) and the further condition $\operatorname{aj}^{*}(\Delta_{g,s})\in\mathsf{K}_{g,2s}$ ?
In other words, do the elements $\operatorname{aj}^{*}(\Delta_{g,s})\in\mathsf{K}_{g,2s}$ generate all the Gorenstein kernels of $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ using properties (i)-(v)? Our evidence is limited: the assertion is true for $\mathsf{K}_{6,0}$, $\mathsf{K}_{7,0}$, and $\mathsf{K}_{5,2}$. These cases correspond to [10, Theorem 5], [10, Proposition 40], and Theorem 11 above, respectively, and depend upon the calculations of the Gorenstein kernels $\mathsf{K}_{g,n}$ for low values of $(g,n)$ in [8]. A conjecture, incorporating further Abel-Jacobi pullbacks, is proposed in Section 6.4.
1.12. Acknowledgments
The calculation of $\mathsf{Tor}^{*}[\mathcal{A}_{2}\times\mathcal{A}_{g-2}]\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$ with the goal of determining whether the homomorphism property holds for $([\mathcal{J}_{g}],[\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ was a group project at ETH Zürich in the Fall term of 2025 with both theoretical and computational sides. We dedicate our paper to Carel Faber: his work [24] is the starting point of several of the themes developed here, and he pioneered the computational exploration of the tautological ring of the moduli space of curves.
We thank David Holmes for discussions about $\mathcal{A}_{2}\times\mathcal{A}_{g-2}\rightarrow\mathcal{A}_{g}$ on the Strand Noordwijk aan Zee in 2023, Gerard van der Geer for discussions about the homomorphism property in Leiden in November 2025, and Aaron Pixton for discussions in Ann Arbor in December 2025 about Conjecture 16 and Section 6. We also benefitted from conversations about the moduli space of abelian varieties with Sam Grushevsky, Sam Molcho, and Dan Petersen.
S.C. was supported by the SNF Ambizione grant PZ00P2-223473. L.D. was supported by SNF-200020-219369. A.I.L. was supported by SNF-200020-219369. D.N. was supported by a Hermann-Weyl-instructorship at the Forschungsinstitut für Mathematik at ETH Zürich. R.P. was supported by SNF-200020-219369 and SwissMAP. J.S. was supported by a SwissMAP fellowship.
2. Fiber product with the Torelli map
2.1. Overview
Let $\underline{g}=(g_{1},\dots,g_{k})$ be a partition of $g$ with $g_{i}\geq 1$. Consider the fiber diagram
$\mathcal{Z}$$\mathcal{M}_{g}^{\mathrm{ct}}$$\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}}$$\mathcal{A}_{g}\,.$$\scriptstyle{\mathsf{Tor}}$
We describe here the local structure of the stack $\mathcal{Z}$. Our presentation follows [10] where the case $\underline{g}=(1,g-1)$ was treated. The structure result for $\mathcal{Z}$ for all partitions $\underline{g}$ is proven by Drakengren [18].
2.2. Proof of Theorem 4
The starting point of the analysis of the stratification of the fiber product $\mathcal{Z}$ by $\underline{g}$-colored extremal trees is the following result [13, Corollary 3.23].
Proposition 18.
A principally polarized abelian variety $(X,\theta)$ decomposes uniquely, up to reordering, as a product of indecomposable principally polarized abelian varieties.
As an immediate consequence, we can describe when the Jacobian of a compact type curve lies in a product locus.
Corollary 19.
Let $C$ be a compact type curve of genus $g$ with Jacobian isomorphic (as principally polarized abelian varieties) to a product
$\mathsf{Jac}(C)\cong X_{g_{1}}\times\dots\times X_{g_{k}}\text{ with }(X_{g_{i}},\theta_{i})\in\mathcal{A}_{g_{i}}\,.$
Then there is a partition of $\mathrm{Irr}(C)_{>0}$, the set of irreducible components of $C$ of positive genus, into disjoint nonempty subsets
$\mathrm{Irr}(C)_{>0}=S_{1}\sqcup\ldots\sqcup S_{k}$
such that, for all $1\leq i\leq k$, we have
$\prod_{D\in S_{i}}\mathsf{Jac}(D)\cong X_{g_{i}}\,.$
In particular, for all $1\leq i\leq k$, we have $\sum_{D\in S_{i}}g(D)=g_{i}$ .
Using Corollary 19, we stratify the fiber product $\mathcal{Z}=\mathsf{Tor}^{-1}(\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}})$. The strata are indexed by $\underline{g}$-colored extremal trees $\mathsf{T}$ as defined in the statement of Theorem 4 of Section 1.5. For each such tree $\mathsf{T}$, we have
$\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}}=\prod_{v\in\mathsf{V}(\mathsf{T})}\mathcal{M}_{\mathsf{g}(v),\mathsf{n}(v)}^{\mathrm{ct}}\,$,
where $\mathsf{n}(v)$ is the valence of $v$. Using the coloring $\mathsf{c}$, we obtain a canonical Torelli map
$\mathsf{Tor}_{\mathsf{T}}:\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}}\rightarrow\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}}\,,\quad C\mapsto\mathsf{Jac}(C_{1})\times\dots\times\mathsf{Jac}(C_{k})\,$,
where the subcurve $C_{i}$ is the union of all positive-genus components of $C$ whose corresponding vertices are assigned the color $i$. Let
$\xi_{\mathsf{T}}:\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}}\rightarrow\mathcal{M}_{g}^{\mathrm{ct}}$
be the gluing map associated to the underlying tree $\mathsf{T}$. Since $\xi_{\mathsf{T}}$ and $\mathsf{Tor}_{\mathsf{T}}$ agree after mapping to $\mathcal{A}_{g}$, we obtain a canonical map
$\epsilon_{\mathsf{T}}:\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}}\rightarrow\mathcal{Z}\,.$
The images of these maps stratify the fiber product $\mathcal{Z}$.
The next definition allows us to describe incidences between the various strata constructed above.
Definition 20.
Let $\mathsf{T},\mathsf{T}^{\prime}$ be two $\underline{g}$-colored extremal trees with color assignments
$\mathsf{c}:\mathsf{V}(\mathsf{T})_{>0}\to\{1,\ldots,k\}\,,\quad\mathsf{c}^{\prime}:\mathsf{V}(\mathsf{T}^{\prime})_{>0}\to\{1,\ldots,k\}\,.$
A $\mathsf{T}^{\prime}$-structure on $\mathsf{T}$ is given by a partition of the vertex set of $\mathsf{T}$,
$\mathsf{V}({\mathsf{T}})=\mathcal{V}_{1}\cup\ldots\cup\mathcal{V}_{m}\,$,
together with a surjective map
$\phi:\mathsf{V}(\mathsf{T})\rightarrow\mathsf{V}(\mathsf{T}^{\prime})$
satisfying the following properties:
  1. (i)
    For each $1\leq i\leq m$, the vertex subset $\mathcal{V}_{i}$ determines a connected nonempty subtree of $\mathsf{T}$.
  2. (ii)
    The assignment $\phi$ sends all vertices in $\mathcal{V}_{i}$ to the same vertex $\phi(\mathcal{V}_{i})\in\mathsf{V}(\mathsf{T}^{\prime})$. If $i\neq j$, then
    $\phi(\mathcal{V}_{i})\neq\phi(\mathcal{V}_{j})\,.$
  3. (iii)
    We have
    $\sum_{v\in\mathcal{V}_{i}}\mathsf{g}(v)=\mathsf{g}(\phi(\mathcal{V}_{i}))\,.$
    Moreover, all positive genus vertices $v\in\mathcal{V}_{i}$ (if any) have the same color and in this case
    $\mathsf{c}(v)=\mathsf{c}^{\prime}(\phi(\mathcal{V}_{i}))\,.$
  4. (iv)
    An edge $e^{\prime}\in\mathsf{E}(\mathsf{T}^{\prime})$ connects the vertices $v^{\prime},w^{\prime}\in\mathsf{V}(\mathsf{T^{\prime}})$ with $v^{\prime}=\phi(\mathcal{V}_{i})$ and $w^{\prime}=\phi(\mathcal{V}_{j})$ if and only if there exists an edge $e\in\mathsf{E}(\mathsf{T})$ which connects a vertex of $\mathcal{V}_{i}$ to a vertex of $\mathcal{V}_{j}$.∎
Item (ii) implies that the sets $\mathcal{V}_{1},\ldots,\mathcal{V}_{m}$ are uniquely determined as the fibers of the surjection $\phi$. We note that if $\mathsf{T}$ has a $\mathsf{T}^{\prime}$-structure, then we must have $|\mathsf{V}(\mathsf{T}^{\prime})|\leq|\mathsf{V}(\mathsf{T})|$, with equality if and only if $\phi$ is an isomorphism preserving the genus and color assignments.
We define $\mathsf{T}$ to be a nontrivial degeneration of $\mathsf{T}^{\prime}$ if $\mathsf{T}$ carries a nontrivial $\mathsf{T}^{\prime}$-structure. We also refer to $\mathsf{T}^{\prime}$ as a smoothing of $\mathsf{T}$. For each degeneration $\mathsf{T}$ of $\mathsf{T}^{\prime}$, the assignment $\phi$ determines a gluing map
$\xi_{\mathsf{T}^{\prime},\mathsf{T}}:\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}}\rightarrow\mathcal{M}_{\mathsf{T}^{\prime}}^{\mathrm{ct}}\,.$
An irreducible $\underline{g}$-colored extremal tree $\mathsf{I}$ is a $\underline{g}$-colored extremal tree without genus 0 vertices. Irreducible trees $\mathsf{I}$ admit no nontrivial smoothings and index the irreducible components of the fibered product $\mathcal{Z}$. Let ${\mathsf{I}}_{1},\ldots,{\mathsf{I}}_{\ell}$ be irreducible $\underline{g}$-colored extremal trees. The intersection of the images of $\mathcal{M}^{\mathrm{ct}}_{\mathsf{I}_{1}},\ldots,\mathcal{M}^{\mathrm{ct}}_{\mathsf{I}_{\ell}}$ in $\mathcal{Z}$ is the union of the images of $\mathcal{M}^{\mathrm{ct}}_{\mathsf{T}}$ taken over all $\underline{g}$-colored extremal trees $\mathsf{T}$ which admit $\mathsf{I}_{1},\ldots,\mathsf{I}_{\ell}$-structures. ∎
The strict stratum $\mathcal{M}^{\mathrm{ct},\star}_{\mathsf{T}}$ associated to a $\underline{g}$-colored extremal tree $\mathsf{T}$ is the complement of the images of all gluing maps $\xi_{\mathsf{T}^{\prime},\mathsf{T}}$ for all $\underline{g}$-colored nontrivial degenerations $\mathsf{T}^{\prime}$ of $\mathsf{T}$. Of course, the strict strata also cover $\mathcal{Z}.$
2.3. Local equations
We now describe the stack structure on the fibered product $\mathcal{Z}$. The map
$\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}}\rightarrow\mathcal{A}_{g}$
is not an embedding, but the differential is an injection on tangent spaces. Therefore, locally in the analytic topology, $\mathcal{Z}$ is a closed substack of $\mathcal{M}_{g}^{\mathrm{ct}}$. Let $x\in\mathcal{Z}$ be a point in the strict stratum $\mathcal{M}^{\mathrm{ct},\star}_{\mathsf{T}}$. We take a small open neighborhood $W$ of $x$, which we may assume is the versal deformation space of $x$. We have a map
$\nu:W\rightarrow\prod_{e\in\mathsf{E}(\mathsf{T})}\mathbb{C}_{e}$
to the deformation space of the nodes of the underlying curve $C$. Here, $\mathbb{C}_{e}$ is the one-dimensional versal deformation space of $C$ corresponding to a node, and hence an edge $e$ of $\mathsf{T}$.
Let $z_{e}$ be the standard coordinate for each factor $\mathbb{C}_{e}$. We write local equations near $x$ using these coordinates.
Definition 21.
Let $v_{1},v_{2}\in\mathsf{V}(\mathsf{T})_{>0}$ such that $\mathsf{c}(v_{1})\neq\mathsf{c}(v_{2})$. Let $\mathsf{P}$ be the minimal path of edges connecting $v_{1}$ to $v_{2}$. We say that $\mathsf{P}$ is critical if the only vertices of positive genus on the path are $v_{1}$ and $v_{2}$. ∎
To each critical path $\mathsf{P}_{\textrm{crit}},$ we associate the monomial
$\mathsf{Mon}(\mathsf{P}_{\mathrm{crit}})=\prod_{e}z_{e}\,$,
where the product is taken over all edges appearing in $\mathsf{P}_{\textrm{crit}}.$ We have the following generalization of [10, Proposition 29].
Theorem 22 (Drakengren [18]).
The local equations for $\mathcal{Z}$ near the point $x$ in the strict stratum indexed by the $\underline{g}$-colored extremal tree $\mathsf{T}$ are given by pullback from $\prod_{e\in\mathsf{E}(\mathsf{T})}\mathbb{C}_{e}$ of the monomial set
$\{\mathsf{Mon}(\mathsf{P}_{\mathrm{crit}}):\mathsf{P}_{\mathrm{crit}}\text{ is a critical path}\}\subset\mathbb{C}[\{z_{e}\}_{e\in\mathsf{E}(\mathsf{T})}]\,.$
In particular, $\mathcal{Z}$ is a reduced Deligne-Mumford stack.
In [10], the local structure of $\mathcal{Z}$ was determined in case $\underline{g}=(1,g-1)$ by interpreting $\mathcal{Z}$ as a moduli space of stable maps to a moving elliptic curve. The deformation theory of stable maps was then used to prove reducedness. For the general case, an interpretation in terms of stable maps is not available. Drakengren [18] directly studies the Torelli map near the boundary of $\mathcal{M}_{g}^{\mathrm{ct}}$ using the results of [38].
3. Excess intersection calculation
3.1. Overview
We prove a graph sum formula for
(21) $\mathsf{Tor}^{*}([\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}}])\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$
using the method of [10, Section 5]. The calculations here generalize those of [10], which concern only the case $\underline{g}=(1,g-1)$.
3.2. Tautological classes
Before studying the pullback (21), we review first a few standard definitions regarding the tautological ring of the moduli space of curves $\overline{\mathcal{M}}_{g,n}$. The tautological classes of $\mathcal{M}_{g,n}^{\mathrm{ct}}$ are obtained by restriction
$\mathsf{R}^{*}(\overline{\mathcal{M}}_{g,n})\to\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\to 0\,.$
A detailed discussion of the tautological rings of $\overline{\mathcal{M}}_{g,n}$ can be found in [23, 57, 56]. The following tautological classes will appear in our calculations:
  • the cotangent classes $\psi_{1},\ldots,\psi_{n}$ given as
    $\psi_{1}=c_{1}(\mathbb{L}_{1}),\ldots,\psi_{n}=c_{1}(\mathbb{L}_{n})\,$,
    where the fiber of $\mathbb{L}_{i}$ over $(C,x_{1},\ldots,x_{n})$ equals $T^{*}_{x_{i}}C\,$,
  • $\lambda$-classes obtained as Chern classes of the Hodge bundle $\mathbb{E}_{g}\to\mathcal{M}_{g,n}^{\mathrm{ct}}$,
    $\lambda_{i}=c_{i}(\mathbb{E}_{g})\,$,
    pulled back from $\mathcal{A}_{g}$ under the Torelli map,
  • the $\kappa$-classes,
    $\kappa_{i}=\pi_{*}\left(c_{1}(\omega_{\pi}(s_{1}+\ldots+s_{n}))^{i+1}\right)$
    where $s_{1},\ldots,s_{n}$ are the universal sections of the universal curve over $\mathcal{M}_{g,n}^{\mathrm{ct}}.$
The tautological ring $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{\mathsf{T}})\subset\mathsf{CH}^{*}(\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}})$ for any $\underline{g}$-colored extremal tree is generated by the pullbacks of the tautological rings of the factors of the product
$\mathcal{M}_{\mathsf{T}}^{\mathsf{ct}}=\prod_{v\in\mathsf{V}(\mathsf{T})}\mathcal{M}_{\mathsf{g}(v),\mathsf{n}(v)}^{\mathrm{ct}}\,.$
Pushforward under the gluing map $\xi_{\mathsf{T}}:\mathcal{M}_{\mathsf{T}}^{\mathsf{ct}}\to\mathcal{M}_{g}^{\mathrm{ct}}$ sends tautological classes on $\mathcal{M}_{\mathsf{T}}^{\mathsf{ct}}$ to tautological classes on $\mathcal{M}_{g}^{\mathrm{ct}}.$
3.3. Proof of Theorem 5
We will express the class (21) as a sum of contributions $\mathsf{Cont}_{\mathsf{T}}$ supported on $\mathcal{M}^{\mathrm{ct}}_{\mathsf{T}}$ for each $\underline{g}$-colored extremal tree $\mathsf{T}$. Only trees $\mathsf{T}$ with at most
$d=\mathsf{cod}_{\underline{g}}=\operatorname{codim}_{\mathcal{A}_{g}}(\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}})=\sum_{1\leq i<j\leq k}g_{i}g_{j}$
edges contribute. The corresponding contributions are computed inductively, with the base cases provided by the irreducible $\underline{g}$-colored extremal trees.
Each contribution is expressed in terms of the Chern classes of the normal bundle of the product
$\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}}\to\mathcal{A}_{g}$
and the Chern classes of the normal bundles of the strata of $\mathcal{Z}$ indexed by the $\underline{g}$-colored extremal trees. These contributions can be found using excess residual intersections as in [29, Chapter IX]. The exact residual terms are universal expressions depending only on the normal bundle data, so we can compute them in a convenient local model. The local equations in Theorem 22 are used for the local model calculations.
Let $c$ be the number of critical paths between all possible unordered pairs of positive genus vertices in $\mathsf{T}$ that have distinct colors. A simple check shows that $c\leq d$. Let $\mathcal{N}$ be a rank $d$ vector bundle on $\prod_{e\in\mathsf{E}(\mathsf{T})}\mathbb{C}_{e}$ of the form
$\mathcal{N}=\mathcal{O}^{c}\oplus L_{1}\oplus\dots\oplus L_{d-c}\,$,
where the $L_{i}$ are arbitrary torus equivariant line bundles on $\prod_{e\in\mathsf{E}(\mathsf{T})}\mathbb{C}_{e}$. Set
$s=(\mathsf{Mon}(\mathsf{P}^{1}_{\mathrm{crit}}),\dots,\mathsf{Mon}(\mathsf{P}^{c}_{\mathrm{crit}}),0,\dots,0)\,.$
Let $\mathsf{Z}(s)\subset\prod_{e\in\mathsf{E}(\mathsf{T})}\mathbb{C}_{e}$ be the zero locus of the section $s$. The zero locus $\mathsf{Z}(s)$ has precisely the equations of our local model by Theorem 22 and is independent of the ordering of the monomials in $s$. We may calculate the contribution of $\mathsf{T}$ in the excess intersection problem for $c_{d}(\mathcal{N})$ determined by $\mathsf{Z}(s)$. We start with the equation
$c_{d}(\mathcal{N})=(\ell_{1}\cdots\ell_{d-c})\prod_{\mathsf{P}_{\mathrm{crit}}}(\sum_{e\in\mathsf{P}_{\mathrm{crit}}}z_{e})\,$,
where the $\ell_{i}$ are the equivariant first Chern classes of the line bundles $L_{i}$.
  • (i)
    Consider first the case where the $\underline{g}$-colored extremal tree $\mathsf{I}$ has no genus 0 vertices. Then, $\mathsf{I}$ corresponds to an irreducible component of $\mathcal{Z}$, and every edge of $\mathsf{I}$ is a critical path. The contribution $\mathsf{Cont}_{\mathsf{I}}$ can be computed by the usual excess intersection formula:
    (22) $\mathsf{Cont}_{\mathsf{I}}=\left[\frac{c(\mathcal{N})}{\prod_{e\in\mathsf{E}(\mathsf{I})}\,(1+z_{e})\,}\right]_{d-c}\,.$
    The subscript indicates that only the part of degree dcd-c is taken. The pushforward $\iota_{{\mathsf{I}}*}$ to the ambient torus-equivariant $\prod_{e\in\mathsf{E}(\mathsf{T})}\mathbb{C}_{e}$ is found by multiplying by the top Chern class of the normal bundle, which equals $z_{1}\cdots z_{c}$. Therefore,
    $\iota_{\mathsf{I}*}\mathsf{Cont}_{\mathsf{I}}=z_{1}\cdots z_{c}\left[\frac{c(\mathcal{N})}{\prod_{e\in\mathsf{E}(\mathsf{I})}\,(1+z_{e})\,}\right]_{d-c}\,.$
  • (ii)
    Next, let $\mathsf{T}$ be an arbitrary $\underline{g}$-colored extremal tree. By induction, we can assume we have computed $\mathsf{Cont}_{\mathsf{T}^{\prime}}$ for all smoothings $\mathsf{T}^{\prime}$ of $\mathsf{T}$. We set
    (23) $\mathsf{Cont}_{\mathsf{T}}\cdot\prod_{e\in\mathsf{E}(\mathsf{T})}z_{e}=c_{d}(\mathcal{N})-\sum_{\mathsf{T}^{\prime}}\iota_{\mathsf{T}^{\prime}*}\mathsf{Cont}_{\mathsf{T}^{\prime}}\,.$
    Equation (23) determines $\mathsf{Cont}_{\mathsf{T}}$.
The formula for $\mathsf{Cont}_{\mathsf{T}}$ in terms of tautological classes is then obtained via substitution of variables:
  • $\bullet$
    we replace each edge variable $z_{e}$ by the normal factor corresponding to the smoothing of the edge $e$ (the sum of tangent lines corresponding to the two half-edges of $e$),
  • $\bullet$
    we replace the Chern classes of $\mathcal{N}$ by the Chern classes of the normal bundle of the product
    $\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}}\to\mathcal{A}_{g}\,.$
In the end, $\mathsf{Cont}_{\mathsf{T}}$ is expressed in terms of tautological $\psi$ and $\lambda$ classes obtained from the moduli of curves. ∎
3.4. Examples
We present full calculations of the contributions of Theorem 5 in a few cases.
Example 23.
Let $\mathsf{I}$ be an irreducible $\underline{g}$-colored extremal tree, and let
$\mathsf{V}(\mathsf{I})=S_{1}\sqcup\ldots\sqcup S_{k}$
be the partition of the vertices determined by the $k$ colors. Then, $\mathcal{M}_{\mathsf{I}}^{\mathrm{ct}}$ determines an irreducible component of $\mathcal{Z}=\mathsf{Tor}^{-1}(\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}})$ with contribution calculated by (22). The normal bundle to the product $\mathcal{A}_{g_{1}}\times\dots\times\mathcal{A}_{g_{k}}\rightarrow\mathcal{A}_{g}$ is
$\operatorname{Sym}^{2}(\mathbb{E}_{g_{1}}^{\vee}\boxplus\dots\boxplus\mathbb{E}_{g_{k}}^{\vee})-\sum_{i=1}^{k}\operatorname{Sym}^{2}(\mathbb{E}_{g_{i}}^{\vee})=\sum_{i<j}\mathbb{E}^{\vee}_{g_{i}}\boxtimes\mathbb{E}^{\vee}_{g_{j}}\,.$
When pulled back to $\mathcal{M}^{\mathrm{ct}}_{\mathsf{I}}$, the normal bundle splits as
$\sum_{i<j}\left(\bigoplus_{v\in S_{i}}\mathbb{E}^{\vee}_{\mathsf{g}(v)}\right)\boxtimes\left(\bigoplus_{w\in S_{j}}\mathbb{E}^{\vee}_{\mathsf{g}(w)}\right)\,.$
The normal bundle of $\mathcal{M}^{\mathrm{ct}}_{\mathsf{I}}$ in $\mathcal{M}_{g}^{\mathrm{ct}}$ is the sum of factors corresponding to the smoothings of each node. Therefore, the contribution of $\mathsf{I}$ is
(24) $\mathsf{Cont}_{\mathsf{I}}=\left[\frac{\prod_{i<j}\prod_{v\in S_{i}}\prod_{w\in S_{j}}c\left(\mathbb{E}^{\vee}_{\mathsf{g}(v)}\boxtimes\mathbb{E}^{\vee}_{\mathsf{g}(w)}\right)}{\prod_{e\in\mathsf{E}(\mathsf{I})}(1-\psi_{e}^{\prime}-\psi^{\prime\prime}_{e})}\right]_{d-c}\,$,
where $\psi^{\prime}_{e},\psi^{\prime\prime}_{e}$ are the cotangent classes at the node associated to $e$. ∎
For the partition $\underline{g}=(1,g-1)$, several explicit computations involving trees $\mathsf{T}$ which are not irreducible are detailed in [10, Examples 30-35]. We record here a new example.
Example 24.
Consider the partition $\underline{g}=(2,4)$ for $g=6$. Then, $d=8$. We use two colors, blue and green, to paint the positive genus vertices. Consider the following $\underline{g}$-colored extremal tree $\mathsf{T}$:
114
Here, the gray vertex of $\mathsf{T}$ has genus 0. We decorate the edge incident to the vertex of genus $4$ by $z_{3}$, while $z_{1}$, $z_{2}$ correspond to the remaining two edges. The tree $\mathsf{T}$ admits two smoothings $\mathsf{R}$ and $\mathsf{S}$ respectively:
42141
The smoothings $\mathsf{R}$ and $\mathsf{S}$ correspond to irreducible components of $\mathsf{Tor}^{-1}(\mathcal{A}_{2}\times\mathcal{A}_{4})$. Therefore, we have
$\iota_{\mathsf{R}*}\mathsf{Cont}_{\mathsf{R}}=z_{3}\left[\frac{c(\mathcal{N})}{1+z_{3}}\right]_{7},\quad\iota_{\mathsf{S}*}\mathsf{Cont}_{\mathsf{S}}=z_{1}z_{2}\left[\frac{c(\mathcal{N})}{(1+z_{1})(1+z_{2})\,}\right]_{6}\,$,
where
$c(\mathcal{N})=(1+\ell_{1})\cdots(1+\ell_{6})(1+z_{1}+z_{3})(1+z_{2}+z_{3})\,.$
By (23), we have
$\mathsf{Cont}_{\mathsf{T}}\cdot z_{1}z_{2}z_{3}=(z_{1}+z_{3})(z_{2}+z_{3})(\ell_{1}\cdots\ell_{6})-\iota_{\mathsf{R}*}\mathsf{Cont}_{\mathsf{R}}-\iota_{\mathsf{S}*}\mathsf{Cont}_{\mathsf{S}}\,.$
After solving for $\mathsf{Cont}_{\mathsf{T}}$, we obtain:
$\mathsf{Cont}_{\mathsf{T}}=$ $-3c_{5}(\mathcal{N})+(4z_{1}+4z_{2}+6z_{3})\,c_{4}(\mathcal{N})-(5z_{1}^{2}+5z_{1}z_{2}+10z_{1}z_{3}+5z_{2}^{2}+10z_{2}z_{3}+10z_{3}^{2})\,c_{3}(\mathcal{N})$
$+$ $(6z_{1}^{3}+6z_{1}^{2}z_{2}+15z_{1}^{2}z_{3}+6z_{1}z_{2}^{2}+15z_{1}z_{2}z_{3}+20z_{1}z_{3}^{2}+6z_{2}^{3}+15z_{2}^{2}z_{3}+20z_{2}z_{3}^{2}+15z_{3}^{3})\,c_{2}(\mathcal{N})$
- $(7z_{1}^{4}+7z_{1}^{3}z_{2}+21z_{1}^{3}z_{3}+7z_{1}^{2}z_{2}^{2}+21z_{1}^{2}z_{2}z_{3}+35z_{1}^{2}z_{3}^{2}+7z_{1}z_{2}^{3}$
$+21z_{1}z_{2}^{2}z_{3}+35z_{1}z_{2}z_{3}^{2}+35z_{1}z_{3}^{3}+7z_{2}^{4}+21z_{2}^{3}z_{3}+35z_{2}^{2}z_{3}^{2}+35z_{2}z_{3}^{3}+21z_{3}^{4})\,c_{1}(\mathcal{N})$
$+$ $(8z_{1}^{5}+8z_{1}^{4}z_{2}+28z_{1}^{4}z_{3}+8z_{1}^{3}z_{2}^{2}+28z_{1}^{3}z_{2}z_{3}+56z_{1}^{3}z_{3}^{2}+8z_{1}^{2}z_{2}^{3}+28z_{1}^{2}z_{2}^{2}z_{3}+56z_{1}^{2}z_{2}z_{3}^{2}$
$+70z_{1}^{2}z_{3}^{3}+8z_{1}z_{2}^{4}+28z_{1}z_{2}^{3}z_{3}+56z_{1}z_{2}^{2}z_{3}^{2}+70z_{1}z_{2}z_{3}^{3}+56z_{1}z_{3}^{4}+8z_{2}^{5}+28z_{2}^{4}z_{3}+56z_{2}^{3}z_{3}^{2}$
$+70z_{2}^{2}z_{3}^{3}+56z_{2}z_{3}^{4}+28z_{3}^{5})\,.\qed$
3.5. Computer implementation
Theorem 7 concerns partitions with $2$ parts $\underline{g}=(h,g-h)$, specifically for $\underline{g}$ in the set
(25) $\{\,(2,2)\,,(2,3)\,,(2,4)\,,(3,3)\,,(2,5)\,,(2,6)\,\}\,.$
For these cases, the number of $\underline{g}$-colored extremal trees with at most $d$ edges is
9,37,153,210,622,2569,9\,,37\,,153\,,210\,,622\,,2569\,,
respectively. A computer calculation yields the required Torelli pullbacks in $\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}}).$
We have implemented the computation of $\mathsf{Cont}_{\mathsf{T}}$ using (23) in the Julia package $\mathsf{TorelliTrees.jl}$ [37]. Given the partition $\underline{g}$, we first enumerate all $\underline{g}$-colored extremal trees up to isomorphism in increasing order of the number of vertices. The smoothings are obtained by first computing the minimal smoothings: the smoothing that can not be obtained by applying two or more non-trivial smoothings. Each edge $\{v_{1},v_{2}\}\in\mathsf{E}(\mathsf{T})$ with $\mathsf{g}(v_{1})\mathsf{g}(v_{2})=0$ determines a unique minimal smoothing. All other smoothings are obtained as iterations of minimal smoothings. We then apply (23) to obtain $\textsf{Cont}_{\mathsf{T}}$ for each $\mathsf{T}$.
Finally, we substitute for $\{z_{e}\}_{e\in\mathsf{E}(\mathsf{T})}$ and the Chern classes of $\mathcal{N}$ as explained above. The output generated is $\mathsf{admcycles}$ code, which, when executed, provides an expression for
$\mathsf{Tor}^{*}([\mathcal{A}_{h}\times\mathcal{A}_{g-h}])\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$
in terms of tautological classes.
To illustrate, the $9$ graphs with at most $4$ edges which appear for the partition $\underline{g}=(2,2)$ are shown below. There are two colors, blue and green, and as usual, the gray vertices have genus 0. The Torelli fiber product has $4$ irreducible components corresponding to the following graphs:
222211221111221111111111
The remaining $5$ graphs correspond to the intersections of these components which appear in the calculation.
11011221101122110111111111101111111101111
3.6. Proof of Theorem 7
For the $6$ partitions $\underline{g}=(h,g-h)$ listed in (25), the procedure described in Section 3.5 confirms that
(26) $\mathsf{Tor}^{*}\left([\mathcal{A}_{h}\times\mathcal{A}_{g-h}]\right)=\mathsf{Tor}^{*}\left(\mathsf{taut}([\mathcal{A}_{h}\times\mathcal{A}_{g-h}])\right)\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})\,.$
The homomorphism property stated in Theorem 7 then follows. Indeed, pushing forward (26) under $\mathsf{Tor}$, we obtain
$\mathcal{A}_{h}\times\mathcal{A}_{g-h}]\cdot[\mathcal{J}_{g}]=\mathsf{taut}([\mathcal{A}_{h}\times\mathcal{A}_{g-h}])\cdot[\mathcal{J}_{g}]\in\mathsf{CH}^{*}(\mathcal{A}_{g})\,.$
Applying the operator $\mathsf{taut}$ and using (3), we conclude
$\mathsf{taut}([\mathcal{A}_{h}\times\mathcal{A}_{g-h}]\cdot[\mathcal{J}_{g}])=\mathsf{\mathsf{taut}}(\mathsf{taut}([\mathcal{A}_{h}\times\mathcal{A}_{g-h}])\cdot[\mathcal{J}_{g}])=\mathsf{taut}([\mathcal{A}_{h}\times\mathcal{A}_{g-h}])\cdot\mathsf{taut}([\mathcal{J}_{g}])\,$,
as required. ∎
Motivated by the calculations for $g\leq 8$, we expect the equality
(27) $\mathsf{Tor}^{*}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])\stackrel{{\scriptstyle?}}{{=}}\mathsf{Tor}^{*}(\mathsf{taut}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}]))\in\mathsf{R}^{2g-4}(\mathcal{M}_{g}^{\mathrm{ct}})\,$
to hold for all $g\geq 2$. The homomorphism property for the pair $([\mathcal{A}_{2}\times\mathcal{A}_{g-2}],[\mathcal{J}_{g}])$ is then implied.
Equality (27) is equivalent to an identity for Hodge integrals. The $\lambda_{g}$-pairing
$\mathsf{R}^{1}(\mathcal{M}_{g}^{\mathrm{ct}})\times\mathsf{R}^{2g-4}(\mathcal{M}_{g}^{\mathrm{ct}})\to\mathsf{R}^{2g-3}(\mathcal{M}_{g}^{\mathrm{ct}})\stackrel{{\scriptstyle\epsilon}}{{\longrightarrow}}\mathbb{Q}$
is perfect by [8, Theorem 4]. Furthermore, $\mathsf{R}^{1}(\mathcal{M}_{g}^{\mathrm{ct}})$ is generated by the boundary divisors and the class $\kappa_{1}$, see [1, Theorem 1]. The proof of [10, Theorem 4] shows that both sides of (27) intersect the boundary divisors trivially. The key ingredient here is the assertion that all product loci in $\mathcal{A}_{g}$ intersect trivially [10, Proposition 17], so in particular
$\mathcal{A}_{2}\times\mathcal{A}_{g-2}]\cdot[\mathcal{A}_{r}\times\mathcal{A}_{g-r}]=0\in\mathsf{CH}^{*}(\mathcal{A}_{g})\,.$
Thus, equation (27) is equivalent to
(28) $\kappa_{1}\cdot\mathsf{Tor}^{*}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])=\kappa_{1}\cdot\mathsf{Tor}^{*}(\mathsf{taut}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}]))\in\mathsf{R}^{2g-3}(\mathcal{M}_{g}^{\mathrm{ct}})\,.$
By applying Theorem 5 and the $\lambda_{g}$-evaluation $\epsilon$, (28) can be equivalently rewritten as an identity involving Hodge integrals:
(29) $\sum_{\mathsf{T}\in\mathsf{Tree}(2,g-2)}\frac{1}{|\mathsf{Aut}(\mathsf{T})|}\int_{{\overline{\mathcal{M}}_{\mathsf{T}}}}\xi_{\mathsf{T}}^{*}(\kappa_{1}\cdot\lambda_{g})\cdot\mathsf{Cont}_{\mathsf{T}}=\int_{\overline{\mathcal{M}}_{g}}\kappa_{1}\cdot\lambda_{g}\cdot\mathsf{taut}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])\,$,
where $\mathsf{Cont}_{\mathsf{T}}$ denotes the canonical extension to the compactified space
$\overline{\mathcal{M}}_{\mathsf{T}}=\prod_{v\in\mathsf{V}(\mathsf{T})}\overline{\mathcal{M}}_{\mathsf{g}(v),\mathsf{n}(v)}\,.$
Using the formula for $\mathsf{taut}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ in Section 1.3, the right side of (29) becomes
$\frac{1}{360}\cdot\frac{g(g-1)}{|B_{2g}||B_{2g-2}|}\int_{\overline{\mathcal{M}}_{g}}\kappa_{1}\cdot\lambda_{g}\cdot\lambda_{g-1}\lambda_{g-3}=\frac{g(g-1)}{1440}\cdot\frac{1}{(2g-2)!}\,$,
where the Hodge integral above is computed by [20, Theorem 3], in particular equation (50) therein. The evaluation of the left hand side of (29) for all $g\geq 2$ is an open question.
By similar reasoning, the homomorphism property for the pair $([\mathcal{A}_{1}\times\mathcal{A}_{1}\times\mathcal{A}_{g-2}],\,[\mathcal{J}_{g}])$, established in [26], is equivalent to the identity
(30) $\sum_{\mathsf{T}\in\mathsf{Tree}(1,1,g-2)}\frac{1}{|\mathsf{Aut}(\mathsf{T})|}\int_{{\overline{\mathcal{M}}_{\mathsf{T}}}}\xi_{\mathsf{T}}^{*}\lambda_{g}\cdot\mathsf{Cont}_{\mathsf{T}}=\frac{1}{288}\cdot\frac{1}{(2g-2)!}\,$,
for all $g\geq 2$. The connection (if any) between identities (29) and (30) is not clear.
4. Wall-crossing calculation
4.1. Overview
After reviewing Fulton-MacPherson degenerations and the definition of unramified maps, we present the calculation of
(31) $\mathsf{Tor}^{*}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$
using the wall-crossing formula of Theorem 6. The families Gromov-Witten class $\mathcal{M}^{\mathrm{ct}}_{g}(\pi_{2})]^{\textnormal{red}}$, computed by wall-crossing, is pushed-forward along the morphism $\mathcal{M}^{\mathrm{ct}}_{g}(\pi_{2})\rightarrow\mathcal{M}_{g}^{\mathrm{ct}}$ to obtain $\mathsf{Tor}^{*}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$. As noted in Section 3.6, the homomorphism property for the pair $([\mathcal{J}_{g}],[\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ is implied by the equality
(32) $\mathsf{Tor}^{*}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])=\mathsf{Tor}^{*}(\mathsf{taut}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}]))\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})\,.$
We have checked (32) for $2\leq g\leq 8$ by this approach.
4.2. Fulton-MacPherson degenerations
Stable maps from curves to a target variety allow for nodal degenerations of the domain curves, but do not degenerate the target. Unramified maps [44], however, permit the target to degenerate. Before defining unramified maps, we recall the definition of the degenerations of the targets which arise.
Let $X$ be a nonsingular projective variety of dimension $r$, which for us will be an abelian variety. The configuration space of $n$ distinct labeled points in $X$ can be compactified by means of the Fulton-MacPherson space $\textrm{FM}_{n}(X)$ constructed in [28]. The space $\textrm{FM}_{n}(X)$ is obtained as an iterated blowup of the product $X^{n}$ along various diagonals in a specified order, encoding how the points of $X$ collide. There is a universal family
(33) $\mathcal{W}\to\textrm{FM}_{n}(X)\,$,
constructed as an iterated blowup of $\textrm{FM}_{n}(X)\times X,$ which carries a projection
$\rho_{n}:\mathcal{W}\to X\,.$
Fulton-MacPherson degenerations
$\rho:W\to X$
are obtained by taking $W$ to be a fiber of the universal family (33), for arbitrary $n$, and letting $\rho$ denote the restriction of $\rho_{n}$ to the fiber $W$. An example of a Fulton-MacPherson degeneration is
(34) $W=\mathrm{Bl}_{0}(X)\cup_{E=D}\mathbb{P}^{r},$
which is given by the degeneration to the normal cone of the point $0\in X$. Here, $D$ is the divisor at infinity in $\mathbb{P}^{r}\cong\mathbb{P}(T_{0}X\oplus\mathbb{C})$ and $E\cong\mathbb{P}(T_{0}X)$ is the exceptional divisor of the blowup of $X$. In general, $W$ is an iterated degeneration to the normal cones of regular points.
4.3. Unramified maps
Let $(\rho:W\rightarrow X,x_{1},\ldots,x_{m})$ denote a marked Fulton-MacPherson degeneration of an abelian variety $X$. We require that the markings $x_{1},\ldots,x_{m}$ are pairwise distinct and are contained in the nonsingular locus of $W$.
For a principally polarized abelian variety $(X,\theta)$, we are interested in the minimal curve class
$\beta_{\min}=\frac{\theta^{r-1}}{(r-1)!}\in H_{2}(X,\mathbb{Z})\,.$
We give the definition of unramified maps from a nodal curve $C$ to $W$ in the class $\beta_{\min}$ with prescribed multiplicity profiles at the markings $x_{i}$.
Definition 25.
Let $(\rho:W\rightarrow X,x_{1},\ldots,x_{m})$ be a marked Fulton-MacPherson degeneration of $X$. Let $C$ be a curve with (at worst) nodal singularities. A map $f:C\rightarrow W$ has curve class $\beta_{\min}\in H_{2}(X,\mathbb{Z})$ if
(35) $(\rho\circ f)_{*}[C]=\beta_{\min}\,.\qed$
Definition 26 ([44], [51]).
A map $f\colon C\rightarrow W$ is unramified if
  • (i)
    $f$ is a logarithmic morphism  [^15] We use the canonical logarithmic structures coming from the singularities of $C$ and $W$, respectively. with nowhere vanishing logarithmic derivative
    $df^{\log}\colon T^{\log}_{C}\rightarrow f^{*}T^{\log}_{W}\,$,
  • (ii)
    the group of automorphisms
    $\{(\gamma_{1},\gamma_{2})\in\operatorname{{\mathsf{Aut}}}(C)\times\operatorname{{\mathsf{Aut}}}(W,x_{1},\ldots,x_{m})\mid f\circ\gamma_{1}=\gamma_{2}\circ f\}$
    is finite. [^16] For a general curve class $\beta$, a more intricate stability condition has to be imposed: see [44, Definition 3.1.1 (4)].
Fix ordered partitions $(\mu^{1},\ldots,\mu^{m})$, and write
[…]
for their parts. Let $\{p_{ij}\}$ be distinct nonsingular points on $C$, for $1\leq i\leq m$, $1\leq j\leq\ell(\mu^{i}).$
Definition 27.
The map $f:(C,\{p_{ij}\})\to(W,\{x_{i}\})$ has multiplicity $\mu^{i}$ at the marked points $x_{i}\in W$ provided
$f^{-1}(x_{i})=\sum^{\ell(\mu^{i})}_{j=1}\mu^{i}_{j}\,p_{ij}\,$,
for $1\leq i\leq m$. We continue to require conditions (i) and (ii) of Definition 26, modified so the logarithmic structure and the automorphisms of the domain also take into account the markings $\{p_{ij}\}$. ∎
Let $\pi_{r}:\mathcal{X}_{r}\to\mathcal{A}_{r}$ denote the universal principally polarized abelian variety. Write
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet}$
for the moduli space of unramified maps parameterizing:
  • marked Fulton-MacPherson degenerations
    $\rho:W\to X,\,x_{1},\ldots,x_{m}$
    of arbitrary principally polarized abelian varieties $(X,\theta)\in\mathcal{A}_{r}$,
  • unramified maps
    $f:(C,\{p_{ij}\})\to(W,\{x_{i}\})$
    having curve class $\beta_{\min}\in H_{2}(X,\mathbb{Z})$ and multiplicity $\mu^{i}$ at the marked points $x_{i}\in W$, $1\leq i\leq m$.
The domain curves $C$ are assumed to be (at worst) nodal, of compact type, possibly disconnected, and of genus $g(C)=g_{0}$. The maps are considered up to automorphisms of Fulton-MacPherson degenerations, which include translations of abelian varieties and affine automorphisms of projective spaces. The markings of the domain are not explicitly indicated in the notation, for ease of reading. The analogously defined moduli space with connected domain curves $C$ will be denoted $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\circ}$.
Because of the minimality of the curve class, we will only need multiplicities of the form
$\mu^{i}=(1,\ldots,1)\,.$
For other curve classes $\beta$, other multiplicities may also arise.
4.4. Definition of $\psi$-classes
Let $\mathcal{FM}_{m}(\pi_{r})$ be the moduli space of Fulton-MacPherson degenerations of abelian varieties with $m$ markings [^17] The markings are pairwise distinct and contained in the nonsingular locus of the Fulton-MacPherson degeneration. The marked degenerations are allowed to have an infinite group of automorphisms.. There exists a natural forgetful morphism
(36) $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet}\rightarrow\mathcal{FM}_{m}(\pi_{r})\,.$
For a subset $I\subset\{1,\ldots,m\}$, consider the divisor
$\mathcal{D}_{I}\subset\mathcal{FM}_{m}(\pi_{r})\,$,
whose generic element parameterizes a Fulton-MacPherson degeneration of the form (34) where the marked points labeled by $I$ are contained on the bubble $\mathbb{P}^{r}$. The descendant class associated to the $i^{th}$ marking $x_{i}$ is defined by
$\Psi_{i}=\sum_{i\in I}\mathcal{D}_{I}\,.$
Pulling back under the forgetful morphism (36), we obtain the class
$\Psi_{i}\in\mathsf{CH}^{1}(\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet})\,.$
Up to certain normalization terms, these $\Psi$-classes can also be defined via vector bundles associated with the cotangent spaces at the markings, see [50, Section 3.2] for more details. The same definition applies to $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\circ}$.
4.5. Star-shaped graphs
A star-shaped graph $\mathsf{S}$ is a rooted graph whose non-root vertices are connected to the root by one edge. We allow the star-shaped graphs to have genus labels on vertices and partition labels on edges, as depicted in Figure 1. In the figure, the non-root vertices are gray. The automorphism group $\operatorname{{\mathsf{Aut}}}(\mathsf{S})$ of a graph $\mathsf{S}$ is defined to be the automorphism group of the set of pairs $\{(g_{1},\mu^{1}),\dots,(g_{m},\mu^{m})\}$.
$\mu^{1}$$\mu^{2}$$\mu^{m}$$g_{0}$$g_{1}$$g_{2}$$g_{m}$
Figure 1. Star-shaped graph with partition labels on edges.
4.6. Blowup description for $r=2$
The wall-crossing formula in Theorem 6 is a sum of contributions indexed by star-shaped graphs. For $r=2$, only two types of moduli spaces appear in the wall-crossing terms:
(37) $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{2}(\pi_{2},\underbrace{(1),\ldots,(1)}_{k})^{\circ}\ \ \ \text{and}\ \ \ \mathcal{M}^{\mathrm{ct},\mathrm{un}}_{1}(\pi_{2},(1,1),\underbrace{(1),\ldots,(1)}_{k})^{\bullet}\,.$
Indeed, for non-product abelian surfaces, the curves in the minimal class $\beta_{\min}$ are smooth genus 2 curves, while for product abelian surfaces $E_{1}\times E_{2}$, the minimal curves are nodal genus 2 curves of compact type. The moduli spaces (37) can be described as iterated blowups of
$\mathcal{M}^{\mathrm{ct}}_{2,k}\ \ \ \text{and}\ \ \ \mathcal{M}_{1,2+k}^{\bullet}=\bigcup_{k_{1}+k_{2}=k}\mathcal{M}^{\mathrm{ct}}_{1,p_{1}+k_{1}}\times\mathcal{M}^{\mathrm{ct}}_{1,p_{2}+k_{2}}\,$,
respectively, where $p_{i}$ are distinguished markings corresponding to the multiplicity $(1,1)$ point.
Let us start with a description of the blowup centers for the first type of moduli spaces of unramified maps. Given an integer $1\leq\ell\leq k$, we define
$Z_{\ell}\subset\mathcal{M}^{\mathrm{ct}}_{2,k}$
to be the locus with generic element parameterizing a nodal genus $2$ curve with a rational bridge that connects two genus $1$ components and contains $\ell$ marked points. For $\ell=1$, such a curve is depicted on the left side of Figure 2. The locus $Z_{\ell}$ is of pure codimension 2 with possibly several non-disjoint nonsingular irreducible components (corresponding to all possible ways that the points can be distributed among the genus 1 components and the rational bridge).
We blow up $\mathcal{M}_{2,k}^{\mathrm{ct}}$ along the $Z_{\ell}$ in the following order:
$\widetilde{\mathcal{M}}_{2,k}=\mathrm{Bl}_{\widetilde{Z}_{1}}\mathrm{Bl}_{\widetilde{Z}_{2}}\ldots\mathrm{Bl}_{Z_{k}}\mathcal{M}^{\mathrm{ct}}_{2,k}\,.$
More precisely, we start by blowing up $Z_{k}$ and then proceed inductively: at the $i^{th}$ stage, we blow up the proper transform $\widetilde{Z}_{k-i}$ of $Z_{k-i}$ for $i=0,\dots,k-1$. After the blowups at stage i1i-1, the proper transform $\widetilde{Z}_{k-i}$ is nonsingular, since all of the smooth irreducible components become disjoint, as explained in [26]. In particular, the space $\widetilde{\mathcal{M}}_{2,k}$ is nonsingular.
Similarly, given an integer $1\leq\ell\leq k$, let
$V_{\ell}\subset\mathcal{M}_{1,2+k}^{\bullet}$
be the locus with generic element parameterizing a pair of genus 1 curves with the distinguished markings $p_{1}$ and $p_{2}$ and $\ell$ other markings on rational tails [^18] In other words, there are $\ell+2$ marked points including $p_{1}$ and $p_{2}$ on rational tails of the two elliptic curves with at least $2$ on each.. As above, we define the iterated blowup along loci $V_{\ell}$,
$\widetilde{\mathcal{M}}^{\bullet}_{1,2+k}=\mathrm{Bl}_{\widetilde{V}_{1}}\mathrm{Bl}_{\widetilde{V}_{2}}\ldots\mathrm{Bl}_{V_{k}}\mathcal{M}^{\bullet}_{1,2+k}\,.$
The moduli space $\widetilde{\mathcal{M}}^{\bullet}_{1,2+k}$ is also nonsingular (for the same reason).
Theorem 28 ([26]).
There exist natural isomorphisms
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{2}(\pi_{2},\underbrace{(1),\ldots,(1)}_{k})^{\circ}\cong\widetilde{\mathcal{M}}_{2,k}\quad\text{and}\quad\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{1}(\pi_{2},(1,1),\underbrace{(1),\ldots,(1)}_{k})^{\bullet}\cong\widetilde{\mathcal{M}}^{\bullet}_{1,2+k},$
such that the reduced virtual fundamental classes coincide with the standard fundamental classes. Moreover, each stage of the iterated blowup admits a moduli-theoretic interpretation in terms of maps to marked Fulton-MacPherson degenerations of abelian surfaces.
4.7. Geometric interpretation of the blowup
To explain how these blowups arise through the moduli problem of unramified maps, let us consider $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{2}(\pi_{2},(1))^{\circ}$. By associating to the marked point $x_{1}\in W$ the fiber $f^{-1}(x_{1})\in C$, we obtain an identification of spaces
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{2}(\pi_{2},(1))^{\circ}\cong\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{2,1}(\pi_{2})^{\circ},$
where the latter is the moduli space of unramified maps with marked points on the domain curve. If the domain $C$ is a nonsingular genus $2$ curve, then $C$ embeds into a non-product abelian surface $X=\textsf{Jac}(C)$ via the Abel-Jacobi map. If $C$ is a nodal genus 2 curve with the marked point on a rational bridge, then the Abel-Jacobi map contracts the rational bridge to a point in $E_{1}\times E_{2}$. The resulting map is not unramified because the differential vanishes at all points on the rational bridge. Hence, the target $E_{1}\times E_{2}$ must sprout a $\mathbb{P}^{2}$-bubble, over which the rational bridge embeds as a quadric. The geometry is illustrated in Figure 2.
$Bl_{0}X$
$\mathbb{P}^{2}$
$g=0$
$g=1$
$g=1$
Figure 2. An example of an unramified map from a nodal genus $2$ curve to a Fulton-MacPherson degeneration of an abelian surface.
The intersection points of the quadric with the hyperplane at infinity are determined, and the $\mathbb{P}^{2}$-bubble is considered up to translations and dilations $\mathbb{G}^{2}_{a}\rtimes\mathbb{G}_{m}$. Overall, the moduli of such quadrics up to these automorphisms of $\mathbb{P}^{2}$ is a point. However, there is an extra degree of freedom associated with choosing a point on the quadric to which the marked point on the rational bridge is mapped – there is a $\mathbb{P}^{1}$-moduli of such choices. As a result, we must blow up $\mathcal{M}^{\mathrm{ct}}_{2,1}$ along the locus $Z_{1}$. With more marked points, there can be more Fulton-MacPherson expansions, which are realized via the iterated blowups of $Z_{\ell}$.
4.8. Blowup description of classes
Using Theorem 28, we can describe all the classes appearing in the wall-crossing formula of Theorem 6 in terms of tautological and exceptional classes. More precisely, the classes that require a separate description are the cotangent line classes $\Psi_{i}$ and the hyperplane classes $H$. The Chern roots $\alpha_{j}$ can be identified with the pullbacks of the Chern roots of the Hodge bundles from $\mathcal{M}^{\mathrm{ct}}_{2}$ and $\mathcal{M}_{1,2}^{\bullet}$.
Let us start with $\widetilde{\mathcal{M}}_{2,k}$. For a non-empty subset $I\subset\{1,\ldots,k\}$, let $E_{I}\subset\widetilde{\mathcal{M}}_{2,k}$ denote the exceptional divisor [^19] The divisors $E_{I}$ may be reducible. corresponding to the locus
$Z_{I}\subset Z_{|I|},$
with generic element parameterizing a nodal genus 2 curve with a rational bridge that connects two genus 1 components and contains marked points labeled by the set $I$. Let
$E_{i}=\sum_{i\in I}E_{I}.$
Proposition 29 ([26]).
With respect to the isomorphism
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{2}(\pi_{2},\smash{\underbrace{(1),\ldots,(1)}_{k}})^{\circ}\cong\widetilde{\mathcal{M}}_{2,k}$
of Theorem 28, for $1\leq i\leq k$, we have:
$\Psi_{i}$ $=\psi_{i}-\epsilon_{i}^{*}\psi_{1}+E_{i}\,$,
$\operatorname{ev}_{i}^{*}H$ $=\epsilon_{i}^{*}\psi_{1}-E_{i}\,$,
$\operatorname{ev}_{i}^{*}\alpha_{j}$ $=\alpha_{j}\,$,
where $\epsilon_{i}\colon\mathcal{M}^{\mathrm{ct}}_{2,k}\longrightarrow\mathcal{M}^{\mathrm{ct}}_{2,1}$ is the morphism which forgets all but the $i^{th}$ marking. The $\psi$ classes are pullbacks of their counterparts on $\mathcal{M}_{2,k}^{\mathrm{ct}}$.
We now proceed with $\widetilde{\mathcal{M}}^{\bullet}_{1,2+k}$. Let $E$ be the total exceptional divisor of $\widetilde{\mathcal{M}}^{\bullet}_{1,2+k}\rightarrow\mathcal{M}^{\bullet}_{1,2+k}$, so $E$ is the union of all exceptional divisors of the iterated blowups.
Proposition 30 ([26]).
With respect to the isomorphism
$\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{1}(\pi_{2},(1,1),\smash{\underbrace{(1),\ldots,(1)}_{k}})^{\bullet}\cong\widetilde{\mathcal{M}}^{\bullet}_{1,2+k}$
of Theorem 28, we have:
$\Psi_{1}$ $=\psi_{p_{1}}+\psi_{p_{2}}-E,$
$\Psi_{1+i}$ $=\psi_{i}\,,\ \ \ 1\leq i\leq k,$
$\operatorname{ev}_{i}^{*}H$ $=0\,,\ \ \ 1\leq i\leq k+1,$
$\operatorname{ev}_{i}^{*}\alpha_{j}=\alpha_{j}$ $=0\,,\ \ \ 1\leq i\leq k+1,$
where the $\psi$ classes are pullbacks of their counterparts on $\mathcal{M}^{\bullet}_{1,2+k}$.
The main computational challenge of the wall-crossing method for $r=2$ is to compute the pushforwards of powers of exceptional classes. Below, we will provide expressions for all relevant powers of exceptional classes for $g=4$ and $g=5$.
4.9. Calculations for $g=4$
The graphs involved in the wall-crossing formula of Theorem 6 for $g=4$ are depicted in Figure 3. Exceptional classes appear only for the first two graphs, which correspond to blowups of $\mathcal{M}^{\mathrm{ct}}_{2,1}$ and $\mathcal{M}^{\mathrm{ct}}_{2,2}$ respectively. The last two graphs have too few markings for exceptional loci to occur: there must be at least $4$ markings on the root vertex for a locus $V_{\ell}\subset\mathcal{M}^{\bullet}_{1,2+k}$ to be non-empty. The pushforwards of powers of exceptional divisors can be computed directly through the description of the centers of blowups and their normal bundles.
$(12)(1^{2})1122(12)(1^{2})(1)(1)11$$11$
Figure 3. Graphs for $r=2$ and $g=4$, where $(1^{2})$ denotes the partition $(1,1)$.
The first graph corresponds to the blowup $\tau\colon\widetilde{\mathcal{M}}_{2,1}\rightarrow\mathcal{M}^{\mathrm{ct}}_{2,1}$ along $Z_{1}$. The normal bundle of $Z_{1}\subset\mathcal{M}^{\mathrm{ct}}_{2,1}$ is trivial, therefore we have
$\tau_{*}(E^{2}_{1})=-Z_{1}\,,\quad\tau_{*}(E_{1}^{\ell})=0\ \ \ \text{for}\ \ \ \ell\neq 2\,.$
Here, $E_{1}$ is the unique exceptional divisor.
The second graph corresponds to the iterated blowup $\tau\colon\widetilde{\mathcal{M}}_{2,2}\rightarrow\mathcal{M}^{\mathrm{ct}}_{2,2}$ along $Z_{2}$ and $Z_{1}$. Inspecting equation (13), we see that the $I$-functions $I_{1,(1)}(z)$ appearing in Theorem 6 have $z$-degree equal to $1$ in this case. Hence at most the first powers of $E_{1}$ and $E_{2}$ contribute, and
$\tau_{*}(E_{i})=0\,,\ \ \ \tau_{*}(E_{1}\cdot E_{2})=\tau_{*}(E_{\{1,2\}}^{2})=-Z_{\{1,2\}}\,$,
where $Z_{\{1,2\}}\subset\mathcal{M}^{\mathrm{ct}}_{2,2}$ is the locus $Z_{I}$ for $I=\{1,2\}$.
The above cases cover all of the exceptional contributions to the wall-crossing formula for $g=4$. The remainder of the calculation lies in the tautological rings
$\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{2,k})\ \ \ \text{and}\ \ \ \mathsf{R}^{*}(\mathcal{M}^{\bullet}_{1,2+k})\,$
and can be computed using admcycles. We find that the pushforward of $\mathcal{M}^{\mathrm{ct}}_{4}(\pi_{2})]^{\textnormal{red}}$ along
$\mathcal{M}^{\mathrm{ct}}_{4}(\pi_{2})\rightarrow\mathcal{M}_{4}^{\mathrm{ct}}$
equals $\mathsf{Tor}^{*}(\mathsf{taut}([\mathcal{A}_{2}\times\mathcal{A}_{2}]))=42\lambda_{1}\lambda_{3}\in\mathsf{R}^{*}(\mathcal{M}_{4}^{\mathrm{ct}})$.
4.10. Calculations for $g=5$
The graphs involved in the wall-crossing formula of Theorem 6 for $g=5$ are depicted in Figure 4. Although the complexity of the graphs does not increase much, the number of exceptional classes involved in the calculation jumps. The exceptional classes appear only for the first three graphs (the graphs without $(1,1)$-edges). For the sixth graph corresponding to $\widetilde{\mathcal{M}}^{\bullet}_{1,2+2}$, the $z$-degree of $I$-function $I_{1,(1,1)}(z)$ is 0, so no substitution by $\Psi_{1}$ is required.
(1)(1)2233(1)(1)(1)(1)2222110111111(1)(1)(1)(1)(1)(1)22111111
[…]
Figure 4. Graphs for $r=2$ and $g=5$, where $(1^{2})$ denotes the partition $(1,1)$.
The first graph corresponds to $\widetilde{\mathcal{M}}_{2,1}$, and the same analysis applies as in the case of $g=4$. For the second graph, we analyze the moduli space $\widetilde{\mathcal{M}}_{2,2}$. Since the degrees in $z$ of the corresponding $I$-functions in Theorem 6 are $3$ and $1$ respectively, we must only compute pushforwards of monomials whose degree in $E_{1}$ is at most 3 and whose degree in $E_{2}$ is at most 1. An analysis of the normal bundles of the proper transforms as well as the intersections of the irreducible components of $Z_{1}\cup Z_{2}=Z_{\{1,2\}}\cup Z_{\{1\}}\cup Z_{\{2\}}$ shows:
$\tau_{*}(E_{i})$ $=0\,$, $\qquad\tau_{*}(E_{1}\cdot E_{2})$ $=-Z_{\{1,2\}}\,$,
$\tau_{*}(E_{1}^{2})$ $=-Z_{\{1\}}-Z_{\{1,2\}}\,$, $\qquad\tau_{*}(E_{1}^{2}\!\cdot E_{2})$ $=-\psi_{q_{1}|Z_{\{1\}}}+(\psi_{q_{1}}+\psi_{q_{2}})_{|Z_{\{1,2\}}}\,$,
$\tau_{*}(E_{1}^{3})$ $=-\psi_{q_{1}|Z_{\{1\}}}+(\psi_{q_{1}}+\psi_{q_{2}})_{|Z_{\{1,2\}}}\,$, $\qquad\tau_{*}(E_{1}^{3}\!\cdot E_{2})$ $=0\,$,
where $\psi_{q_{1}}$ is the $\psi$-class on $Z_{\{1\}}$ associated to the node between the rational bridge and the genus $1$ curve containing the second marking and $\psi_{q_{1}}$ and $\psi_{q_{2}}$ are the $\psi$-classes on $Z_{\{1,2\}}$ associated to the two nodes between the rational bridge and the genus $1$ curves.
For $\widetilde{\mathcal{M}}_{2,3}$, at most the first powers of each exceptional class $E_{i}$ appear:
$\tau_{*}(E_{i})$ $=0\,$,
$\tau_{*}(E_{i}\cdot E_{j})$ $=-Z_{\{i,j\}}-Z_{\{1,2,3\}}\,$,
$\tau_{*}(E_{1}\cdot E_{2}\cdot E_{3})$ $=-\psi_{q_{1}|Z_{\{1,2\}}}-\psi_{q_{1}|Z_{\{1,3\}}}-\psi_{q_{1}|Z_{\{2,3\}}}+(\psi_{q_{1}}+\psi_{q_{2}})_{|Z_{\{1,2,3\}}},$
where $\psi$-classes $\psi_{q_{i}}$ are defined in the same way as in the case of two markings. Although the powers are not high, the number of exceptional components increases considerably for three markings: $Z_{1}\cup Z_{2}\cup Z_{3}$ has 10 irreducible components. In general, the blowup centers
$Z_{1}\cup\ldots\cup Z_{k}$
have $(3^{k}-2^{k}-1)/2+1$ components, giving rise to the same number of exceptional components on the blowup. The explosion of exceptional components is the main source of computational complexity for the wall-crossing method.
After an admcycles calculation, we find that the pushforward of $\mathcal{M}^{\mathrm{ct}}_{5}(\pi_{2})]^{\textnormal{red}}$ along
$\mathcal{M}^{\mathrm{ct}}_{5}(\pi_{2})\rightarrow\mathcal{M}_{5}^{\mathrm{ct}}$
equals $\mathsf{Tor}^{*}(\mathsf{taut}([\mathcal{A}_{2}\times\mathcal{A}_{3}]))=22\lambda_{2}\lambda_{4}\in\mathsf{R}^{*}(\mathcal{M}_{5}^{\mathrm{ct}})$.
4.11. Higher genus
The calculation of the contributions obtained by pushing forward exceptional divisors on iterated blowups of $\widetilde{\mathcal{M}}_{2,k}$ and $\widetilde{\mathcal{M}}^{\bullet}_{1,2+k}$ is computationally challenging. The natural setting for these calculations is the logarithmic Chow ring: the algebra of piecewise polynomials on the moduli space of tropical curves. The required logarithmic intersection calculus has been implemented in logtaut in admcycles. Using the wall-crossing formulas, we have verified the homomorphism property for $([\mathcal{J}_{g}],[\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ for $2\leq g\leq 8$ using the code [27].
The algorithm scales linearly with the number of star-shaped graphs and the number of components of the blowup centers. As such, the total run time grows super-exponentially in $g$, and the computation is feasible only for small values of $g$.
5. Tautological projections on $\mathcal{X}_{g}^{s}\rightarrow\mathcal{A}_{g}$
5.1. Overview
We start by defining the tautological ring $\mathsf{R}^{*}(X^{s})$ associated to a self-product of a fixed principally polarized abelian variety $(X,\theta)$ together with a corresponding tautological projection operator
$\mathsf{taut}^{s}_{X}:\mathsf{CH}^{*}(X^{s})\rightarrow\mathsf{R}^{*}(X^{s})\,.$
The construction naturally extends to the fiber product [^20] In Section 1.10, the universal family and the fiber product were denoted by $\pi_{g}$ and $\pi_{g}^{s}$. We drop the subscript $g$ here for notational convenience.
$\pi^{s}:\mathcal{X}_{g}^{s}=\underbrace{\mathcal{X}_{g}\times_{\mathcal{A}_{g}}\cdots\times_{\mathcal{A}_{g}}\mathcal{X}_{g}}_{s}\rightarrow\mathcal{A}_{g}$
of the universal family of abelian varieties $\pi:\mathcal{X}_{g}\rightarrow\mathcal{A}_{g}$. We then derive closed-form expressions for the tautological projections of natural product cycles on $\mathcal{X}_{g}^{s}$.
5.2. The tautological ring of products of abelian varieties
Let $(X,\theta)$ be a principally polarized abelian variety of dimension $g$. There is a canonical [^21] Starting with an arbitrary lift $\widetilde{\theta}$ of the polarization to Chow, the symmetric lift is given by $\theta=\frac{1}{2}(\widetilde{\theta}+(-1)^{*}\widetilde{\theta})\in\mathsf{CH}^{1}(X)\,.$ The symmetric lift $\theta$ is independent of the choice of $\widetilde{\theta}$. We denote both the principal polarization and the symmetric Chow class lift by $\theta$. symmetric rational Chow class $\theta\in\mathsf{CH}^{1}(X)$ of the principal polarization $\theta\in\mathsf{H}^{2}(X,\mathbb{Z})$ to a divisor class of $X$. Let $X^{s}$ be the $s$-fold product for $s\geq 1$. Let
$\theta_{i}\in\mathsf{CH}^{1}(X^{s})\,,\quad 1\leq i\leq s\,$,
be the pullbacks of the theta classes on each of the $s$ factors. Let $\mathcal{P}_{ij}$ be the Poincaré line bundle for each pair of factors $(i,j)$, trivialized along both zero sections:
$\mathcal{P}_{ij}=\pi_{ij}^{*}(\mathcal{P})\,,\quad\pi_{ij}:X^{s}\longrightarrow X\times X\cong X\times X^{\vee}\,,\quad 1\leq i,j\leq s\,,\quad i\neq j\,.$
We define the classes
$\eta_{ij}=c_{1}(\mathcal{P}_{ij})\in\mathsf{CH}^{1}(X^{s})\,,\quad 1\leq i,j\leq s\,,\quad i\neq j\,.$
Our conventions imply $\eta_{ij}=\eta_{ji}$, so we will often take $i<j$. For convenience, let $\eta_{ii}=\theta_{i}$.
Definition 31.
The tautological ring of $X^{s}$ is the subring $\mathsf{R}^{*}(X^{s})\subset\mathsf{CH}^{*}(X^{s})$ generated by the classes $\theta_{i}$ for $1\leq i\leq s$ and the classes $\eta_{ij}$ for $1\leq i<j\leq s.$
The ring $\mathsf{R}^{*}(X^{s})$ is explicitly described in Proposition 33 below. The result is likely known to the experts, see [32, Section 6], but we give a detailed presentation here.
We study $\mathsf{R}^{*}(X^{s})\subset\mathsf{CH}^{*}(X^{s})$ by considering the images of tautological classes under the cycle class map
(38) $\mathsf{CH}^{*}(X^{s})\rightarrow\mathsf{H}^{2*}(X^{s},\mathbb{Q})\,.$
Let $V\cong\mathbb{Q}^{2g}$ be the standard symplectic vector space. We identify $\mathsf{H}^{1}(X,\mathbb{Q})\cong V$, and, therefore obtain an identification
$\mathsf{H}^{*}(X^{s})\cong\wedge^{*}(V\otimes\mathbb{Q}^{s})\,.$
In symplectic coordinates $\{x_{1},y_{1},\ldots,x_{g},y_{g}\}$ on $X$, we can write
(39) $\theta$ $=\sum_{i=1}^{g}dx_{i}\wedge dy_{i}\in\wedge^{2}V\cong\mathsf{H}^{2}(X,\mathbb{Q})\,$,
$\eta$ $=c_{1}(\mathcal{P})=\sum_{i=1}^{g}dx_{i}^{(1)}\wedge dy_{i}^{(2)}-dy_{i}^{(1)}\wedge dx_{i}^{(2)}\in\wedge^{2}(V\otimes\mathbb{Q}^{2})\cong\mathsf{H}^{2}(X\times X,\mathbb{Q})\,.$
The superscripts in the formula for $\eta$ indicate pullbacks from the factors. These elements are invariant under the action of the symplectic group $\operatorname{Sp}(V)$ on $\wedge^{2}V$ and $\wedge^{2}(V\otimes\mathbb{Q}^{2})$ respectively.
More generally, let
$\mathfrak{I}_{g,s}=\left(\wedge^{*}(V\otimes\mathbb{Q}^{s})\right)^{\operatorname{Sp}(V)}$
be the $\operatorname{Sp}(V)$-invariant subring of $\wedge^{*}(V\otimes\mathbb{Q}^{s})$. On tautological classes, the cycle map (38) then factors through
$\mathsf{R}^{*}(X^{s})\rightarrow\mathfrak{I}_{g,s}\,.$
The structure of the ring of invariants $\mathfrak{I}_{g,s}$ can be made explicit. We fix a basis $e_{1},\ldots,e_{s}$ of $\mathbb{Q}^{s}$. We consider the induced basis $e_{I}$ of the symmetric algebra $\operatorname{Sym}^{\bullet}(\mathbb{Q}^{s})$ indexed by non-decreasing strings $I$ of integers chosen from the set $\{1,\ldots,s\}$. For each $k\geq 1$, there is a canonical embedding
$\iota_{k}:\operatorname{Sym}^{2k}(\mathbb{Q}^{s})\hookrightarrow\operatorname{Sym}^{k}(\operatorname{Sym}^{2}\mathbb{Q}^{s})$
constructed as follows. For each non-decreasing string $I$ of length $2k$, we set
$\iota_{k}(e_{I})=\sum_{(J_{1},\ldots,J_{k})}e_{J_{1}}\ldots e_{J_{k}}\,$,
where the sum is taken over all possible partitions $(J_{1},\ldots,J_{k})$ of $I$ with $k$ parts, each consisting of two elements of the string II. [^22] The sum has $\frac{1}{k!}\binom{2k}{2,\ldots,2}=(2k-1)!!$ terms. The following result is proven using classical invariant theory.
Theorem 32 ([62, Theorem 3.4]).
The ring of invariants has the following presentation
$\mathfrak{I}_{g,s}\cong\frac{\operatorname{Sym}^{\bullet}(\operatorname{Sym}^{2}\mathbb{Q}^{s})}{\langle\operatorname{im}(\iota_{g+1})\rangle}\,.$
The isomorphism in the theorem is induced by
$\wedge^{2}(V\otimes\mathbb{Q}^{s})\to\text{Sym}^{2}\mathbb{Q}^{s},\quad(v\otimes e_{i})\wedge(w\otimes e_{j})\mapsto\omega(v,w)e_{ij}\,$,
where $\omega$ is the symplectic form on $V$. Using (39), it follows that the cycle map (38) is given on generators by
(40) $\eta_{ij}\to 2e_{ij},\quad 1\leq i<j\leq s,\quad\theta_{i}\to e_{ii},\quad 1\leq i\leq s\,.$
In particular, $\mathsf{R}^{*}(X^{s})$ surjects onto $\mathfrak{I}_{g,s}$. In fact, more is true.
Proposition 33.
The cycle map gives an isomorphism
$\mathsf{R}^{*}(X^{s})\cong\mathfrak{I}_{g,s}\,.$
Therefore, the tautological ring $\mathsf{R}^{*}(X^{s})$ is independent (as a $\mathbb{Q}$-algebra) of the choice of the abelian variety $(X,\theta)$.
Proof.
We only need to show that the cycle map is injective. If $a_{1},\ldots,a_{s}$ are formal variables, and $I$ is a non-decreasing string of length $2g+2$ chosen from the set $\{1,2,\ldots,s\}$, then the coefficient of $a_{I}=\prod_{i\in I}a_{i}$ in the expression
$\left(\sum_{1\leq i,j\leq s}a_{i}a_{j}e_{ij}\right)^{g+1}$
is precisely $\frac{(2g+2)!}{|\text{Aut}(I)|}\cdot\frac{1}{(2g+1)!!}\cdot\iota_{g+1}(e_{I})$. Therefore, using (40), all the relations in the image of $\mathsf{R}^{*}(X^{s})$ under the cycle class map come from expanding the basic relation
(41) $\left(\sum_{i=1}^{s}a_{i}^{2}\theta_{i}+\sum_{1\leq i<j\leq s}a_{i}a_{j}\eta_{ij}\right)^{g+1}=0\,$
as a polynomial in the $a_{i}$. The relation (41) is the pullback of the identity $\theta^{g+1}=0\in\mathsf{H}^{*}(X)$ under the morphism
$\varphi:X^{s}\rightarrow X:(x_{1},\ldots,x_{s})\mapsto a_{1}x_{1}+\ldots+a_{s}x_{s}\,$
and, therefore, also holds in the Chow ring of $X^{s}$. The symmetric property of $\theta\in\mathsf{CH}^{1}(X)$ is used here to calculate $\varphi^{*}\theta$ in Chow:
$\varphi^{*}\theta=\sum_{i=1}^{s}a_{i}^{2}\theta_{i}+\sum_{1\leq i<j\leq s}a_{i}a_{j}\eta_{ij}\in\mathsf{CH}^{1}(X^{s})\,.\qed$
Proposition 34.
The ring $\mathsf{R}^{*}(X^{s})$ is Gorenstein, with respect to the integration map, with socle in degree $gs$ generated by
$\frac{\theta_{1}^{g}}{g!}\ldots\frac{\theta_{s}^{g}}{g!}\,.$
Proof.
By Proposition 33, it suffices to show that $\mathfrak{I}_{g,s}$ is Gorenstein. We decompose the cohomology $\mathsf{H}^{*}(X^{s},\mathbb{Q})$ into sums of irreducible representations of $\operatorname{Sp}(V),$ with $\mathfrak{I}_{g,s}$ occurring as the invariant summand. The top degree part $\mathsf{H}^{2gs}(X^{s})$ has rank $1$ and is spanned by the $\operatorname{Sp}(V)$-invariant class
$\frac{\theta_{1}^{g}}{g!}\ldots\frac{\theta_{s}^{g}}{g!}\,.$
Since the intersection pairing on $\mathsf{H}^{*}(X^{s},\mathbb{Q})$ is perfect, the restriction to the invariant summand must also be perfect. Hence, $\mathfrak{I}_{g,s}$ is Gorenstein. ∎
The integration map $\mathsf{R}^{gs}(X^{s})\rightarrow\mathbb{Q}$ has a useful interpretation as follows. Let $\mathsf{Sym}^{2}(s)$ be the set of unordered pairs $I=\{i,j\}$ with $1\leq i,j\leq s$, where repetitions $i=j$ are allowed. For any assignment
$\mathsf{a}:\mathsf{Sym}^{2}(s)\rightarrow\mathbb{Z}_{\geq 0}\,$,
we obtain a tautological monomial in $\mathsf{R}^{*}(X^{s})$:
$\frac{\eta^{\mathsf{a}}}{{\mathsf{a}!}}=\prod_{I\in\mathsf{Sym}^{2}(s)}\frac{{\eta}^{\mathsf{a}(I)}}{\mathsf{a}(I)!}\,$,
where, for $I=\{i,j\}\in\mathsf{Sym}^{2}(s)$, we let $\eta_{I}=\eta_{ij}$ (with the convention $\eta_{ii}=\theta_{i}$). The monomial $\eta^{\mathsf{a}}$ has codimension $\sum_{I\in\mathsf{Sym}^{2}(s)}\mathsf{a}(I).$
Lemma 35.
Let $M$ be an $s\times s$ symmetric matrix with formal coefficients $m_{ij}$, $1\leq i\leq j\leq s$. Then, for any function $\mathsf{a}:\mathsf{Sym}^{2}(s)\to\mathbb{Z}_{\geq 0},$ we have
$\int_{X^{s}}\frac{\eta^{\mathsf{a}}}{\mathsf{a}!}=[m^{\mathsf{a}}]\det(M)^{g}\,$,
where $m^{\mathsf{a}}]$ extracts the coefficient [^23] Here, for $I=\{i,j\}$, we set $m_{I}=m_{ij}$. of $\prod_{I\in\mathsf{Sym}^{2}(s)}m_{I}^{\mathsf{a}(I)}$.
Proof.
Assume $M$ has integer coefficients, and let $\mathcal{L}_{M}\to X^{s}$ denote the line bundle
$\mathcal{L}_{M}=\prod_{i=1}^{s}\theta_{i}^{m_{ii}}\otimes\prod_{1\leq i<j\leq s}\mathcal{P}_{ij}^{m_{ij}}\,.$
The associated morphism
$\varphi_{{\mathcal{L}}_{M}}:X^{s}\rightarrow\widehat{X^{s}}\cong X^{s}\,$
takes the form
$(x_{1},\ldots,x_{s})\mapsto\left(\sum_{j=1}^{s}m_{1j}x_{j},\ldots,\sum_{j=1}^{s}m_{sj}x_{j}\right)\,.$
The degree of $\varphi_{\mathcal{L}_{M}}$ is easily calculated to be $(\det M)^{2g}.$ Therefore, by [5, Corollary 3.6.2],
$\chi(\mathcal{L}_{M})^{2}=\deg(\varphi_{\mathcal{L}_{M}})=\det(M)^{2g}\,.$
By Hirzebruch-Riemann-Roch, we obtain
$\int_{X^{s}}\frac{(\sum_{i<j}m_{ij}\eta_{ij}+\sum_{i}m_{ii}\theta_{i})^{gs}}{(gs)!}=\chi(\mathcal{L}_{M})=\pm\det(M)^{g}\,$,
which we view as a polynomial identity in the variables $m_{ij}$. The sign is easily seen to be positive by comparing the coefficients of $(m_{11}\cdots m_{ss})^{g}$ on both sides. ∎
5.3. Tautological projection
Proposition 34 shows that the pairing
$\mathsf{CH}^{*}(X^{s})\times\mathsf{R}^{gs-*}(X^{s})\rightarrow\mathsf{CH}^{gs}(X^{s})\rightarrow\mathbb{Q}\,$
is nondegenerate when restricted to tautological classes on the leftmost factor. The identification $\left(\mathsf{R}^{gs-*}(X^{s})\right)^{\vee}\cong\mathsf{R}^{*}(X^{s})$ gives rise to a canonical projection operator
$\mathsf{taut}_{X}^{s}:\mathsf{CH}^{*}(X^{s})\rightarrow\mathsf{R}^{*}(X^{s})\,.$
The basic property of the tautological projection is
$\int_{X^{s}}\alpha\cdot\beta=\int_{X^{s}}\mathsf{taut}_{X}^{s}(\alpha)\cdot\beta\,$,
for all $\alpha\in\mathsf{CH}^{*}(X^{s})$ and $\beta\in\mathsf{R}^{*}(X^{s}).$
Proposition 36.
The operator $\mathsf{taut}_{X}^{s}$ is a $\mathbb{Q}$-algebra homomorphism if and only if every class in $\mathsf{CH}^{*}(X^{s})$ is tautological in cohomology. In particular, $\mathsf{taut}_{X}^{s}$ is a homomorphism for a general abelian variety $(X,\theta)\in\mathcal{A}_{g}$, but not for every abelian variety $(X,\theta)$.
Proof.
Assume first that the homomorphism property holds. Let $\alpha,\beta\in\mathsf{CH}^{*}(X^{s})$ be classes of complementary degrees. Since $\alpha\cdot\beta$ is a top degree class, we have
$\int_{X^{s}}\alpha\cdot\beta=\int_{X^{s}}\mathsf{taut}_{X}^{s}(\alpha\cdot\beta)=\int_{X^{s}}\mathsf{taut}_{X}^{s}(\alpha)\cdot\mathsf{taut}_{X}^{s}(\beta)=\int_{X^{s}}\mathsf{taut}_{X}^{s}(\alpha)\cdot\beta\,$,
where in the first and last equality, the definition of the tautological projection is used. Since $\beta$ is arbitrary, the cycle $\alpha-\mathsf{taut}_{X}^{s}(\alpha)$ must be numerically trivial, hence homologically trivial by [45, Theorem 4]. Equivalently, the image of $\alpha$ in cohomology must be tautological.
Conversely, assume that every Chow class on $X^{s}$ is tautological in cohomology. For every class $\alpha\in\mathsf{CH}^{*}(X^{s})$, we can write
$\alpha=\alpha_{0}+\alpha_{1}\,$,
where $\alpha_{0}$ is tautological, and $\alpha_{1}$ is homologically trivial. Having such a decomposition forces $\alpha_{0}=\mathsf{taut}_{X}^{s}(\alpha)$. Indeed, we can easily verify the defining property
$\int_{X^{s}}\alpha\cdot\gamma=\int_{X^{s}}(\alpha_{0}+\alpha_{1})\cdot\gamma=\int_{X^{s}}\alpha_{0}\cdot\gamma$
for all $\gamma\in\mathsf{R}^{*}(X^{s}).$ Therefore, for all $\alpha,\beta\in\mathsf{CH}^{*}(X^{s})$, we have
$\alpha\cdot\beta=(\alpha_{0}+\alpha_{1})\cdot(\beta_{0}+\beta_{1})=\alpha_{0}\cdot\beta_{0}+(\alpha_{0}\cdot\beta_{1}+\alpha_{1}\cdot\beta_{0}+\alpha_{1}\cdot\beta_{1})\,.$
Since the last $3$ terms are homologically trivial, we obtain
$\mathsf{taut}_{X}^{s}(\alpha\cdot\beta)=\alpha_{0}\cdot\beta_{0}=\mathsf{taut}_{X}^{s}(\alpha)\cdot\mathsf{taut}_{X}^{s}(\beta)\,.$
Let $(X,\theta)\in\mathcal{A}_{g}$ be a general abelian variety. The algebra of Hodge classes $\mathsf{Hdg}(X^{s})$ on $X^{s}$ is generated in degree $1$, see [36, Proposition 1.6] which applies thanks to [5, Proposition 17.4.2]. Moreover, for a general $X$, the rational Néron-Severi group of $X^{s}$ is generated by $\theta_{i}$, for $1\leq i\leq s$, and $\eta_{ij}$, for $1\leq i<j\leq s.$  [^24] The Néron-Severi group is the Rosati invariant part of $\text{Hom}(X^{s},\widehat{X}^{s})_{\mathbb{Q}}=\text{Hom}(X^{s},X^{s})_{\mathbb{Q}}=\mathsf{Mat}_{s}(\mathbb{Q})$, which is the space of symmetric matrices, for $X$ general. The standard basis of symmetric matrices corresponds to the classes $\theta_{i}$ and $\eta_{ij}$. Since the image of $\mathsf{CH}^{*}(X^{s})$ lies in $\mathsf{Hdg}(X^{s}),$ and the latter is tautological, every rational Chow class on $X^{s}$ is tautological in cohomology. ∎
5.4. The tautological ring of $\mathcal{X}_{g}^{s}$
We now lift the constructions of Section 5.2 for a fixed abelian variety $(X,\theta)$ to the moduli space $\mathcal{A}_{g}$. Let $\pi:\mathcal{X}_{g}\to\mathcal{A}_{g}$ be the universal abelian variety, and let $\theta\in\mathsf{CH}^{1}(\mathcal{X}_{g})$ be the symmetric theta class normalized by requiring the restriction of $\theta$ to the zero section of $\pi$ to vanish. Over the $s$-fold fibered product $\pi^{s}:\mathcal{X}_{g}^{s}\to\mathcal{A}_{g}$, we have divisor classes
$\big\{\,\theta_{i}\in\mathsf{CH}^{1}(\mathcal{X}_{g}^{s})\,\big\}_{1\leq i\leq s}\ \ \ \text{and}\ \ \ \big\{\,\eta_{ij}\in\mathsf{CH}^{1}(\mathcal{X}_{g}^{s})\,\big\}_{1\leq i\neq j\leq s}\,$,
defined exactly as in Section 5.2 by pullback from the factors. As before, we have $\eta_{ij}=\eta_{ji}$, and we set $\eta_{ii}=\theta_{i}$.
Definition 37.
The tautological ring of $\mathcal{X}_{g}^{s}$ is the subalgebra
$\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\subset\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})$
generated over $\mathsf{R}^{*}(\mathcal{A}_{g})$ by the classes $\theta_{i},$ for $1\leq i\leq s$, and $\eta_{ij}$ for all $1\leq i<j\leq s.$ The subalgebra
$\mathsf{R}^{*}_{\mathsf{vert}}(\mathcal{X}_{g}^{s})\subset\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$
generated over $\mathbb{Q}$ by the classes $\theta_{i}$, for $1\leq i\leq s$, and $\eta_{ij}$ for all $1\leq i<j\leq s$ is the vertical tautological ring. ∎
Let $\overline{\mathcal{X}}_{g}$ be a nonsingular toroidal compactification of $\mathcal{X}_{g}$ lying over a nonsingular toroidal compactification $\overline{\mathcal{A}}_{g}$ of $\mathcal{A}_{g}$. Using the vanishing $\lambda_{g}=0$ on $\overline{\mathcal{X}}_{g}^{s}\smallsetminus\mathcal{X}_{g}^{s}$ established in [9], we can define the $\lambda_{g}$-pairing [^25] We may assume $\overline{\mathcal{X}}_{g}^{s}$ is nonsingular [12, Chapter VI], so that the intersection product is well-defined. on $\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})$ exactly as in (1):
(42) $\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})\times\mathsf{CH}^{gs+\binom{g}{2}-*}(\mathcal{X}_{g}^{s})\rightarrow\mathbb{Q}\,,\quad(\alpha,\beta)=\int_{\overline{\mathcal{X}}_{g}^{s}}\overline{\alpha}\cdot\overline{\beta}\cdot\lambda_{g}\,.$
When $s=1$, we have the isomorphism
$\mathsf{R}^{*}_{\textrm{vert}}(\mathcal{X}_{g})=\mathbb{Q}[\theta]/(\theta^{g+1})\,.$
The vanishing $\theta^{g+1}=0\in\mathsf{CH}^{g+1}(\mathcal{X}_{g})$ holds under our normalization conventions as explained in [34, Theorem 2.1], see also [16, 35]. For $s=2,$ the vertical tautological ring was studied and shown to be Gorenstein in [34, Theorem 3.3]. The tautological ring for families of abelian varieties was also considered in [3, Section 9.4] and [32, Section 6], and $\mathsf{R}^{*}(\mathcal{X}^{s}_{g})$ coincides with the definition of [3] for the universal family.
Theorem 38.
The following isomorphisms determine the tautological rings of $\mathcal{X}_{g}^{s}$ for all $s$:
  • (i)
    $\mathsf{R}^{*}_{\mathsf{vert}}(\mathcal{X}_{g}^{s})\cong\mathsf{R}^{*}(X^{s})$ for every fixed principally polarized abelian variety $(X,\theta)$. Therefore,
    $\mathsf{R}^{*}_{\mathsf{vert}}(\mathcal{X}_{g}^{s})\stackrel{{\scriptstyle\sim}}{{=}}\mathfrak{I}_{g,s}\,.$
  • (ii)
    The natural map $\mathsf{R}^{*}(\mathcal{A}_{g})\otimes\mathsf{R}^{*}_{\mathsf{vert}}(\mathcal{X}_{g}^{s})\longrightarrow\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$ is an isomorphism [^26] All tensor products here are $\otimes_{\mathbb{Q}}$.. Therefore,
    $\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\cong\mathsf{L}_{g-1}\otimes\mathfrak{I}_{g,s}\,$,
    where $\mathsf{L}_{g-1}$ is the Chow ring of the Lagrangian Grassmannian of (g1)(g-1)-dimensional Lagrangian subspaces of $\mathbb{C}^{2g-2}.$
Moreover, the pairing (42) is perfect when restricted to $\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$.
Proof.
There is a natural surjection
(43) $\mathsf{R}^{*}(X^{s})\to\mathsf{R}^{*}_{\textnormal{vert}}(\mathcal{X}_{g}^{s})\,$,
sending the classes $\theta_{i}$ and $\eta_{ij}$ on the left to the analogous classes on the right. The surjection (43) is well-defined because all relations between $\theta_{i}$ and $\eta_{ij}$ on $X^{s}$ also hold on the universal family $\mathcal{X}_{g}^{s}$. Indeed, by the proof of Proposition 33, the ideal of relations in $\mathsf{R}^{*}(X^{s})$ is generated by pullbacks of the identity
$\theta^{g+1}=0\in\mathsf{CH}^{*}(X)$
under multiplication and addition maps. However, the parallel identity
$\theta^{g+1}=0\in\mathsf{CH}^{*}(\mathcal{X}_{g})$
also holds for the normalized theta divisor on the universal abelian variety $\pi:\mathcal{X}_{g}\to\mathcal{A}_{g}$.
Using (43), we obtain a natural map
(44) $\tau:\mathsf{R}^{*}(\mathcal{A}_{g})\otimes\mathsf{R}^{*}(X^{s})\to\mathsf{R}^{*}(\mathcal{A}_{g})\otimes\mathsf{R}_{\textnormal{vert}}^{*}(\mathcal{X}^{s}_{g})\to\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\,$,
which is surjective as well. The tensor product $\mathsf{R}^{*}(\mathcal{A}_{g})\otimes\mathsf{R}^{*}(X^{s})$ is Gorenstein with socle in degree $\binom{g}{2}+gs$, generated by the class
(45) $\lambda_{1}\cdots\lambda_{g-1}\otimes\frac{\theta_{1}^{g}}{g!}\cdots\frac{\theta_{s}^{g}}{g!}\,.$
The image of the socle generator under $\tau$ in $\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$ is non-zero: the $\lambda_{g}$-evaluation (42) of the $\tau$-image (45) is
$\int_{\overline{\mathcal{X}}_{g}^{s}}\lambda_{1}\cdots\lambda_{g}\cdot\frac{\overline{\theta}_{1}^{g}}{g!}\cdots\frac{\overline{\theta}_{s}^{g}}{g!}=\int_{\overline{\mathcal{A}}_{g}}\lambda_{1}\cdots\lambda_{g}=\gamma_{g}\neq 0\,.$
The constant $\gamma_{g}$ is given by (2), and the bars indicate any extensions of the theta classes over the boundary.
On general grounds, if $A$ is Gorenstein with socle in degree $d$, and if $\tau:A\to B$ is a surjective graded morphism which is non-zero in degree $d$, then $\tau$ must be an isomorphism. [^27] If $a\in\text{Ker }\tau$ and $a\neq 0$, we can find $a^{*}\in A$ such that $aa^{*}$ equals the socle generator for $A$. Then $\tau(aa^{*})=\tau(a)\tau(a^{*})=0\,$ which contradicts the non-vanishing of $\tau$ in degree $d$. Item (ii) then follows by applying the general statement to the morphism (44):
$\tau:\mathsf{R}^{*}(\mathcal{A}_{g})\otimes\mathsf{R}^{*}(X^{s})\to\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\,.$
For item (i), we apply the same statement to the morphism (43). Indeed, the non-vanishing of the $\tau$-image of (45) also proves that (43) is nonzero in degree $gs$. ∎
By the same reasoning as in Section 5.3, Theorem 38 yields a projection operator
$\mathsf{taut}^{s}:\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})\to\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\,.$
In addition, we have the projection operators
$\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\to\mathsf{R}^{*}(\mathcal{A}_{g})\,\quad\text{and}\quad\mathsf{taut}^{s}_{X}:\mathsf{CH}^{*}(X^{s})\to\mathsf{R}^{*}(X^{s})\,$,
where $(X,\theta)$ is a fixed abelian variety.
To understand the relationship between the above three projection operators, consider first the following general situation. Let $f:X\rightarrow S$ be a principally polarized abelian scheme of dimension $g$ over an irreducible base $S$, corresponding to a proper morphism $\iota:S\to\mathcal{A}_{g}.$ We have the constructions:
  • (i)
    $X^{s}=X\times_{S}\times\ldots\times_{S}X$ is the $s$-fold fibered product, and $f^{s}:X^{s}\to S$ is the canonical morphism,
  • (ii)
    $j:X^{s}\longrightarrow\mathcal{X}_{g}^{s}$ is the map induced by $\iota$,
  • (iii)
    $X_{\textnormal{gen}}$ is the principally polarized abelian variety over the generic point of $S$.
The geometry is expressed by the diagram:
[…]
Lemma 39.
Let $\alpha\in\mathsf{CH}^{m}(X^{s})$. Assume, for every $\beta\in\mathsf{R}^{k}_{\textnormal{vert}}(\mathcal{X}_{g}^{s})$ with $m+k\neq gs$, the vanishing
(46) $f^{s}_{*}(\alpha\cdot j^{*}\beta)=0\in\mathsf{CH}^{*}(S)\,$
holds. Then, we have
$\mathsf{taut}^{s}(j_{*}\alpha)=\mathsf{taut}_{X_{\textnormal{gen}}}^{s}(\alpha_{\textnormal{gen}})\cdot\left(\pi^{s}\right)^{*}\mathsf{taut}(\iota_{*}[S])\in\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\,$,
where $\alpha_{\textnormal{gen}}\in\mathsf{CH}^{*}(X_{\textnormal{gen}}^{s})$ is the restriction of $\alpha$ to the generic fiber.
Here, to make sense of the right hand side, we use the canonical identification $\mathsf{R}^{*}(X_{\textnormal{gen}}^{s})\cong\mathsf{R}_{\textnormal{vert}}^{*}(\mathcal{X}_{g}^{s})$ established in Theorem 38.
Proof.
The result follows from the definitions via standard arguments, but we provide the details for completeness. It suffices to verify that for all classes $\delta\in\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$, we have
$\int_{\overline{\mathcal{X}}_{g}^{s}}\overline{j_{*}\alpha}\cdot\overline{\delta}\cdot(\overline{\pi}^{s})^{*}\lambda_{g}=\int_{\overline{\mathcal{X}}_{g}^{s}}{\left(\overline{\mathsf{taut}_{X_{\textnormal{gen}}}^{s}(\alpha_{\textnormal{gen}})}\cdot\left(\overline{\pi}^{s}\right)^{*}\mathsf{taut}(\iota_{*}[S])\right)\cdot\overline{\delta}\cdot(\overline{\pi}^{s})^{*}\lambda_{g}}\,.$
The bars indicate extensions to the compactification $\overline{\mathcal{X}}_{g}^{s}$. The tautological classes pulled back from $\mathcal{A}_{g}$ admit canonical extensions to $\overline{\mathcal{A}}_{g}$ by [25], which we assume throughout (and do not indicate explicitly in the notation). With the aid of Theorem 38, we write
$\delta=\beta\cdot(\pi^{s})^{*}\gamma,\quad\beta\in\mathsf{R}^{k}_{\textnormal{vert}}(\mathcal{X}_{g}^{s}),\quad\gamma\in\mathsf{R}^{*}(\mathcal{A}_{g})\,.$
Therefore, we need to show
$\int_{\overline{\mathcal{X}}_{g}^{s}}\overline{j_{*}\alpha}\cdot\overline{\beta}\cdot(\overline{\pi}^{s})^{*}(\gamma\cdot\lambda_{g})=\int_{\overline{\mathcal{X}}_{g}^{s}}{\left(\overline{\mathsf{taut}_{X_{\textnormal{gen}}}^{s}(\alpha_{\textnormal{gen}})}\cdot\left(\overline{\pi}^{s}\right)^{*}\mathsf{taut}(\iota_{*}[S])\right)\cdot\overline{\beta}\cdot(\overline{\pi}^{s})^{*}(\gamma\cdot\lambda_{g}})\,.$
We evaluate both sides by pushing forward under $\overline{\pi}^{s}:\overline{\mathcal{X}}_{g}^{s}\to\overline{\mathcal{A}}_{g}$. We need to establish
(47) $\int_{\overline{\mathcal{A}}_{g}}\overline{\pi}^{s}_{*}\left(\overline{j_{*}\alpha}\cdot\overline{\beta}\right)\cdot\gamma\cdot\lambda_{g}=\int_{\overline{\mathcal{A}}_{g}}\overline{\pi}^{s}_{*}\left(\overline{\mathsf{taut}_{X_{\textnormal{gen}}}^{s}(\alpha_{\textnormal{gen}})}\cdot\overline{\beta}\right)\cdot\mathsf{taut}(\iota_{*}[S])\cdot\gamma\cdot\lambda_{g}\,.$
We first claim that both sides of (47) vanish if $k+m\neq gs$. Indeed, for the left side, the restriction of the class $\overline{\pi}^{s}_{*}\left(\overline{j_{*}\alpha}\cdot\overline{\beta}\right)$ to $\mathcal{A}_{g}$ equals
$\pi^{s}_{*}(j_{*}\alpha\cdot\beta)=\pi^{s}_{*}\,j_{*}(\alpha\cdot j^{*}\beta)=\iota_{*}\,f^{s}_{*}(\alpha\cdot j^{*}\beta)=0\,$,
by the hypothesis of the Lemma. Therefore, the class $\overline{\pi}^{s}_{*}\left(\overline{j_{*}\alpha}\cdot\overline{\beta}\right)$ is supported on the boundary $\overline{\mathcal{A}}_{g}\smallsetminus\mathcal{A}_{g}$. Since $\lambda_{g}$ annihilates the boundary terms, the left side of (47) vanishes. The same argument applies to the right side of (47), but for different reasons. The corresponding vanishing
$\pi^{s}_{*}(\mathsf{taut}_{X_{\textnormal{gen}}}(\alpha)\cdot\beta)=0$
follows by Theorem 38 (i), see also (48) below.
We next consider the case $m+k=gs$. The restriction of the class $\overline{\pi}^{s}_{*}\left(\overline{j_{*}\alpha}\cdot\overline{\beta}\right)$ to $\mathcal{A}_{g}$ equals
$\iota_{*}\,f^{s}_{*}(\alpha\cdot j^{*}\beta)=\mathsf{const}\cdot\iota_{*}([S])\,$,
where the last equality follows for dimension reasons. The constant in the above equation can be computed on the generic fiber of $f^{s}$. Moreover, $\overline{\pi}^{s}_{*}\left(\overline{\mathsf{taut}_{X_{\textnormal{gen}}}^{s}(\alpha_{\textnormal{gen}})}\cdot\overline{\beta}\right)$ restricts on $\mathcal{A}_{g}$ to $\mathsf{const}\cdot[\mathcal{A}_{g}]$, for the same constant (as a consequence of the definition of the tautological projection $\mathsf{taut}_{X_{\textnormal{gen}}}$ on the generic fiber).
Putting everything together, equation (47) becomes
$\int_{\overline{\mathcal{A}}_{g}}\overline{\iota_{*}([S])}\cdot\gamma\cdot\lambda_{g}=\int_{\overline{\mathcal{A}}_{g}}\mathsf{taut}(\iota_{*}[S])\cdot\gamma\cdot\lambda_{g}\,$,
which holds by the definition of the projection operator $\mathsf{taut}$ on $\mathcal{A}_{g}$. ∎
Remark 40.
The assumption in Lemma 39 places a restriction on the class $\alpha$. However, Lemma 39 applies to classes satisfying a natural condition explained here.
For any abelian scheme $p:Y\to S$ of fiber dimension $r$, a class $\alpha\in\mathsf{CH}^{*}(Y)$ has weight $w$ if the multiplication maps $n:Y\to Y$ act as
$n^{*}\alpha=n^{w}\alpha\,.$
If $w\neq 2r$, then
(48) $p_{*}\left(\alpha\right)=0\in\mathsf{CH}^{*}(S)\,$
as a consequence of the calculation
$p_{*}\left(\alpha\right)=\frac{1}{n^{2r}}p_{*}(n_{*}(n^{*}\alpha))=n^{w-2r}p_{*}n_{*}(\alpha)=n^{w-2r}p_{*}(\alpha)\,.$
By a fundamental result of Beauville [4], all classes $\alpha\in\mathsf{CH}^{*}(Y)$ admit a decomposition
(49) $\alpha=\alpha_{1}+\ldots+\alpha_{t}\,$,
where $\alpha_{i}$ has weight $i$.
We apply the weight decomposition to the abelian scheme $f^{s}:X^{s}\to S$ considered in Lemma 39. The classes $\theta_{i}$ and $\eta_{ij}$ on $\mathcal{X}_{g}^{s}$ have weight $2$, so every vertical tautological class $\beta\in\mathsf{R}^{k}_{\textnormal{vert}}(\mathcal{X}_{g}^{s})$ has weight $2k$. If $\alpha\in\mathsf{CH}^{m}(X^{s})$ has a decomposition (49) satisfying
(50) $\alpha_{i}=0\text{ for all }i\leq 2m-1\,$,
then the vanishing (46) holds because of the vanishing (48).
In particular, the vanishing (46) holds when $\alpha$ is the fundamental class of an abelian subscheme of $X^{s}$.
To see an example of a class that does not satisfy (46), let
$\mathcal{C}_{g,1}\rightarrow\mathcal{M}_{g,1}=S$
be the universal curve. Let $\mathcal{X}\rightarrow\mathcal{M}_{g,1}$ be the universal Jacobian, and let $\alpha\in\mathsf{CH}^{g-1}(\mathcal{X})$ be the fundamental class of the universal Abel-Jacobi map $\mathcal{C}_{g,1}\rightarrow\mathcal{X}$ defined via the marking. A straightforward calculation shows
$f_{*}(\alpha\cdot\theta^{2})=(2g-2)\psi+\kappa_{1}\neq 0\,$
whenever $g\geq 2$. ∎
5.5. The homomorphism property for $\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})$
The homomorphism property for the universal fiber product $\mathcal{X}_{g}^{s}$ is formulated by exactly following Definition 3 for $\mathcal{A}_{g}$.
Definition 41.
Let $\alpha$, $\beta\in\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})$. The pair $(\alpha,\beta)$ satisfies the homomorphism property if
(51) $\mathsf{taut}^{s}(\alpha\cdot\beta)=\mathsf{taut}^{s}(\alpha)\cdot\mathsf{taut}^{s}(\beta)\in\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\,.\qed$
Just as for $\mathcal{A}_{g}$, the homomorphism property holds if either $\alpha$ or $\beta$ are tautological. Hence (51) provides an obstruction to classes being tautological. If Conjecture 8 is true, then the homomorphism property (51) holds, by the projection formula, for any pair $(\alpha,\beta)$ with $\alpha\in\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})$ and $\beta$ pulled back from $\mathsf{CH}^{*}(\mathcal{A}_{g})$. We will further investigate the homomorphism property on $\mathcal{X}_{g}^{s}$ in Corollary 50 of Section 6.4.
5.6. Proof of Theorem 13
We compute the tautological projection of the generalized product locus of Definition 12:
$\operatorname{PR}_{g,s}=\underbrace{(\mathcal{A}_{1}\times\mathcal{X}_{g-1})\times_{\mathcal{A}_{g}}\cdots\times_{\mathcal{A}_{g}}(\mathcal{A}_{1}\times\mathcal{X}_{g-1})}_{s}\rightarrow\mathcal{X}_{g}^{s}\,.$
The calculation relies on Lemma 39. More precisely, consider the family of abelian varieties
$f:\mathcal{X}_{1}\times\mathcal{X}_{g-1}\to\mathcal{A}_{1}\times\mathcal{A}_{g-1}$
induced by the product morphism $\iota:S=\mathcal{A}_{1}\times\mathcal{A}_{g-1}\to\mathcal{A}_{g}$. The cycle
$\alpha=[\operatorname{PR}_{g,s}]\in\mathsf{CH}^{s}((\mathcal{X}_{1}\times\mathcal{X}_{g-1})\times_{\mathcal{A}_{1}\times\mathcal{A}_{g-1}}\cdots\times_{\mathcal{A}_{1}\times\mathcal{A}_{g-1}}(\mathcal{X}_{1}\times\mathcal{X}_{g-1}))$
satisfies the vanishing condition (46) by Remark 40. As a result, we have
$\mathsf{taut}^{s}([\operatorname{PR}_{g,s}])$ $=$ $\mathsf{taut}_{X_{\textnormal{gen}}}^{s}(\alpha_{\textnormal{gen}})\cdot\left(\pi^{s}\right)^{*}\mathsf{taut}(\iota_{*}[S])$
$=$ $\mathsf{taut}_{X_{\textnormal{gen}}}^{s}(\alpha_{\textnormal{gen}})\cdot\frac{g}{6|B_{2g}|}\lambda_{g-1}$
in $\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$. Theorem 13 is then a consequence of the following calculation.
Proposition 42.
Let $E$ and $Y$ be principally polarized abelian varieties of dimensions 1 and g1g-1 respectively. Let $X=E\times Y$ with the induced polarization. Then,
$\mathsf{taut}_{X}^{s}\left([(0\times Y)^{s}]\right)=\frac{1}{\kappa_{g,s}}\det\begin{pmatrix}\theta_{1}&\eta_{12}/2&\ldots&\eta_{1s}/2\\ \eta_{12}/2&\theta_{2}&\ldots&\eta_{2s}/2\\ \vdots&\vdots&\ddots&\vdots\\ \eta_{1s}/2&\eta_{2s}/2&\ldots&\theta_{s}\end{pmatrix}\,\in\mathsf{R}^{*}(X^{s})\,$,
where $\kappa_{g,s}=\prod_{k=0}^{s-1}\left(g+\frac{k}{2}\right)$.
Proof.
Let $m_{ij}$ be formal variables such that $m_{ij}=m_{ji}$ for $1\leq i,j\leq s$. Consider the differential operators
$\partial_{ij}=2^{\delta_{ij}-1}\frac{\partial}{\partial m_{ij}}\,$,
where $\delta$ denotes the Kronecker delta. For $\mathsf{p}:\mathsf{Sym}^{2}(s)\to\mathbb{Z}_{\geq 0}$ of codimension $s=\sum_{I\in\textsf{Sym}^{2}(s)}\mathsf{p}(I)$, let $\alpha(\mathsf{p})$ denote the coefficient of
$\prod_{I\in\mathsf{Sym}^{2}(s)}\left(\frac{\partial}{\partial m_{I}}\right)^{\mathsf{p}(I)}$
in the expansion of $\det(\partial_{ij}).$
Fix a function $\mathsf{a}:\mathsf{Sym}^{2}(s)\to\mathbb{Z}_{\geq 0}$. From Lemma 35, we obtain
$\int_{X^{s}}\frac{\eta^{\mathsf{a}}}{{\mathsf{a}}!}\cdot[\operatorname{PR}_{g,s}]=\int_{Y^{s}}\frac{\eta^{\mathsf{a}}}{{\mathsf{a}}!}=[m^{\mathsf{a}}]\det(M)^{g-1}\,$,
for the matrix $M=(m_{ij}).$ We need the following Capelli identity [11] which allows us to differentiate powers of determinants of symmetric matrices:
$\det(\partial_{ij})\det(M)^{g}=\kappa_{g,s}\det(M)^{g-1}\,.$
Therefore,
$\int_{X^{s}}\frac{\eta^{\mathsf{a}}}{{\mathsf{a}}!}\cdot[\operatorname{PR}_{g,s}]$ $=\frac{1}{\kappa_{g,s}}\,[m^{\mathsf{a}}]\det(\partial_{i,j})\det(M)^{g}$
$=\frac{1}{\kappa_{g,s}}\sum_{\mathsf{p}}\frac{(\mathsf{a}+\mathsf{p})!}{{\mathsf{a}}!}\,\alpha(\mathsf{p})\,\,[m^{\mathsf{a}+\mathsf{p}}]\det(M)^{g}$
$=\frac{1}{\kappa_{g,s}}\,\sum_{\mathsf{p}}\alpha(\mathsf{p})\int_{X^{s}}\frac{\eta^{\mathsf{a}+\mathsf{p}}}{{\mathsf{a}}!}$
$=\frac{1}{\kappa_{g,s}}\,\int_{X^{s}}\frac{\eta^{\mathsf{a}}}{\mathsf{a}!}\cdot\left(\sum_{\mathsf{p}}\alpha(\mathsf{p})\cdot\eta^{\mathsf{p}}\right)\,.$
From the definition of $\alpha(\mathsf{p})$, we see
$\sum_{\mathsf{p}}\alpha(\mathsf{p})\cdot\eta^{\mathsf{p}}=\det\begin{pmatrix}\theta_{1}&\eta_{12}/2&\ldots&\eta_{1s}/2\\ \eta_{12}/2&\theta_{2}&\ldots&\eta_{2s}/2\\ \vdots&\vdots&\ddots&\vdots\\ \eta_{1s}/2&\eta_{2s}/2&\ldots&\theta_{s}\end{pmatrix}\,$.
The Proposition then follows from the definition of tautological projection. ∎
Remark 43.
Let $X=Z\times Y$ be a product of principally polarized abelian varieties with $Z$ of dimension $r$ and $Y$ of dimension grg-r. The argument used in the proof of Proposition 42 also yields:
$\mathsf{taut}_{X}^{s}\left([(0\times Y)^{s}]\right)=\frac{1}{\kappa_{g,s}\cdots\kappa_{g-r+1,s}}\det\begin{pmatrix}\theta_{1}&\eta_{12}/2&\ldots&\eta_{1s}/2\\ \eta_{12}/2&\theta_{2}&\ldots&\eta_{2s}/2\\ \vdots&\vdots&\ddots&\vdots\\ \eta_{1s}/2&\eta_{2s}/2&\ldots&\theta_{s}\end{pmatrix}^{r}\in\mathsf{R}^{*}(X^{s})\,.$
Therefore, by Lemma 39, the tautological projection of the cycle
$\underbrace{(\mathcal{A}_{r}\times\mathcal{X}_{g-r})\times_{\mathcal{A}_{g}}\ldots\times_{\mathcal{A}_{g}}(\mathcal{A}_{r}\times\mathcal{X}_{g-r})}_{s}\to\mathcal{X}_{g}^{s}$
is given by
$\frac{1}{\kappa_{g,s}\cdots\kappa_{g-r+1,s}}\det\begin{pmatrix}\theta_{1}&\eta_{12}/2&\ldots&\eta_{1s}/2\\ \eta_{12}/2&\theta_{2}&\ldots&\eta_{2s}/2\\ \vdots&\vdots&\ddots&\vdots\\ \eta_{1s}/2&\eta_{2s}/2&\ldots&\theta_{s}\end{pmatrix}^{r}\cdot\mathsf{taut}([\mathcal{A}_{r}\times\mathcal{A}_{g-r}])\,\in\mathsf{R}^{*}(\mathcal{X}^{s}_{g})\,$,
where $\mathsf{taut}([\mathcal{A}_{r}\times\mathcal{A}_{g-r}])\in\mathsf{R}^{*}(\mathcal{A}_{g})$ was computed in [9, Theorem 1]. ∎
Remark 44.
The class of the generalized product locus $\operatorname{PR}_{g,s}]\in\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})$ will rarely be tautological. Indeed, we have
$\pi_{*}^{s}((\theta_{1}\cdots\theta_{s})^{s(g-1)}\cdot[\operatorname{PR}_{g,s}])=((g-1)!)^{s}\,[\mathcal{A}_{1}\times\mathcal{A}_{g-1}]\,$,
which is non-tautological [^28] The class is expected to be non-tautological for all $g\geq 8$. for $g=6$, $g=12$, and every even $g\geq 16$ [10, 40]. In Section 6, Corollary 48, we will prove that $\operatorname{PR}_{5,1}]$ is not tautological. ∎
6. Generalized products and Abel-Jacobi maps
6.1. Overview
We consider the pullbacks of the generalized product cycles under Abel-Jacobi maps. Using the methods of Sections 3 and 4, we give two proofs that these pullbacks are tautological. We also discuss the relationship between the Abel-Jacobi pullbacks and the kernels of the $\lambda_{g}$-pairing on $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$.
6.2. Abel-Jacobi maps
Let $\mathsf{A}=(a_{ij}\in\mathbb{Z})$ be an $s\times n$ matrix for which the sum of the entries of every row of $\mathsf{A}$ is 0. The matrix $\mathsf{A}$ determines a proper morphism
(52) $\operatorname{aj}_{\mathsf{A}}:\mathcal{M}^{\mathrm{ct}}_{g,n}\to\mathcal{X}_{g}^{s}$
which is defined on $(C,p_{1},\ldots,p_{n})\in\mathcal{M}_{g,n}$ by
$\operatorname{aj}_{\mathsf{A}}(C,p_{1},\ldots,p_{n})\mapsto\left(\mathsf{Jac}(C),\mathcal{O}_{C}\left(\sum_{j=1}^{n}a_{1j}p_{j}\right),\ldots,\mathcal{O}_{C}\left(\sum_{j=1}^{n}a_{sj}p_{j}\right)\right)\,.$
The unique extension of $\operatorname{aj}_{\mathsf{A}}$ over $\mathcal{M}_{g,n}^{\mathrm{ct}}$ involves canonical twists (and lies at the beginning of the theory of the double ramification cycle [33, 35, 42, 57]). The twists are required in order for the corresponding line bundles to have degree zero on every component of the curve.
Two cases are of particular geometric interest. For the $s\times(s+1)$ matrix
$\mathsf{Z}=\begin{pmatrix}-1&1&0&\cdots&0\\ -1&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ -1&0&0&\cdots&1\end{pmatrix}\,$
the induced morphism
(53) $\operatorname{aj}_{\mathsf{Z}}:\mathcal{M}_{g,s+1}^{\mathrm{ct}}\to\mathcal{X}_{g}^{s},\quad(C,p_{0},p_{1},\ldots,p_{s})\mapsto(\mathsf{Jac}(C),\mathcal{O}(p_{1}-p_{0}),\ldots,\mathcal{O}(p_{s}-p_{0}))$
is the $s$-fold product of the classical Abel-Jacobi map constructed via the marking $p_{0}$. A second fundamental case corresponds to the $s\times 2s$ matrix
$\mathsf{B}=\begin{pmatrix}1&-1&0&0&\cdots&0&0\\ 0&0&1&-1&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&0&\cdots&1&-1\end{pmatrix}\,$,
yielding the morphism
(54) $\operatorname{aj}_{\mathsf{B}}:\mathcal{M}_{g,2s}^{\mathrm{ct}}\to\mathcal{X}_{g}^{s},\quad(C,p_{1},p_{2},\ldots,p_{2s-1},p_{2s})\mapsto(\mathsf{Jac}(C),\mathcal{O}_{C}(p_{1}-p_{2}),\ldots,\mathcal{O}_{C}(p_{2s-1}-p_{2s}))\,.$
The Abel-Jacobi map (54) arises geometrically as follows. If $\Gamma$ is the stable graph with one vertex of genus $g$ and $s$ loops, then the composition
$\mathcal{M}^{\mathrm{ct}}_{g,2s}\stackrel{{\scriptstyle\xi_{\Gamma}}}{{\longrightarrow}}\overline{\mathcal{M}}_{g+s}\stackrel{{\scriptstyle{\mathsf{Tor}\ }}}{{\longrightarrow}}\overline{\mathcal{A}}_{g+s}$
can be factorized as
$\mathcal{M}^{\mathrm{ct}}_{g,2s}\rightarrow\mathcal{T}\rightarrow\overline{\mathcal{A}}_{g+s}\,$,
where $\overline{\mathcal{A}}_{g+s}$ is a toroidal compactification for which $\mathsf{Tor}$ extends and $\mathcal{T}$ is a torus-bundle over $\mathcal{X}_{g}^{s}$. The composite $\mathcal{M}^{\mathrm{ct}}_{g,2s}\rightarrow\mathcal{T}\rightarrow\mathcal{X}_{g}^{s}$ is $\operatorname{aj}_{\mathsf{B}}$. A proof of the above factorization for $s=1$ is given in [41, Lemma 5.1], but the argument applies to any $s$.
The pullback along an Abel-Jacobi map $\operatorname{aj}_{\mathsf{A}}$ preserves tautological classes [33, 35]. For a vector $v=(v_{1},\ldots,v_{n})\in\mathbb{Z}^{n}$, let
$\theta(v)=\frac{1}{2}\sum_{i=1}^{n}v_{i}^{2}\psi_{i}-\frac{1}{4}\sum_{h=0}^{g}\sum_{S}v_{S}^{2}\,\delta_{h,S}\,\in\,\mathsf{R}^{1}(\mathcal{M}^{\mathrm{ct}}_{g,n})\,$,
where the last sum is over all subsets $S\subset\{1,\ldots,n\}$. In the summand, $v_{S}=\sum_{i\in S}v_{i}$, and
$\delta_{h,\,S}\in\mathsf{R}^{1}(\mathcal{M}_{g,n}^{\mathrm{ct}})$
is the class of the boundary divisor parameterizing curves with a genus $h$ component which carries exactly the markings of $S$.  [^29] Both $\delta_{h,S}$ and $\delta_{g-h,S^{c}}$ are allowed in the sum. Then, we have
(55) $\operatorname{aj}_{\mathsf{A}}^{*}\theta_{i}$ $=\theta(a_{i1},\ldots,a_{in})\,$,
$\operatorname{aj}_{\mathsf{A}}^{*}\eta_{ij}$ $=\theta(a_{i1}+a_{j1},\ldots,a_{in}+a_{jn})-\theta(a_{i1},\ldots,a_{in})-\theta(a_{j1},\ldots,a_{jn})\,.$
6.3. The Abel-Jacobi pullback of $\operatorname{PR}_{g,s}$
The Abel-Jacobi pullbacks of the generalized product cycles,
$\operatorname{aj}_{{\mathsf{A}}}^{*}\,[\operatorname{PR}_{g,s}]\in\mathsf{CH}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\,$,
can be computed via the methods of either Section 3 or Section 4. These calculations show
$\operatorname{aj}_{{\mathsf{A}}}^{*}\,[\operatorname{PR}_{g,s}]\in\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g,n})\,$,
see Proposition 46 below. More generally, we can consider any class of the form $\mathsf{p}_{*}([\alpha\times\mathcal{X}_{g-1}^{s}])$, where $\alpha\in\mathsf{R}^{*}(\mathcal{X}_{1}^{s})$ and
$\mathsf{p}:(\mathcal{X}_{1}\times\mathcal{X}_{g-1})^{s}=\underbrace{(\mathcal{X}_{1}\times\mathcal{X}_{g-1})\times_{\mathcal{A}_{g}}\cdots\times_{\mathcal{A}_{g}}(\mathcal{X}_{1}\times\mathcal{X}_{g-1})}_{s}\rightarrow\mathcal{X}_{g}^{s}\,$
is the product map. The fibered product over $\mathcal{A}_{g}$ will be understood in the notation throughout Section 6. As a special case,
$\operatorname{PR}_{g,s}]=\mathsf{p}_{*}([\alpha\times\mathcal{X}_{g-1}^{s}])\in\mathsf{R}^{*}(\mathcal{X}_{g}^{s})\text{\ \ for }\alpha=\theta_{1}\cdots\theta_{s}\in\mathsf{R}^{s}(\mathcal{X}_{1}^{s})\,.$
The Abel-Jacobi pullbacks can be expressed in terms of families Gromov-Witten classes for the universal elliptic curve $\pi:\mathcal{E}\to\mathcal{M}_{1,1}.$ To explain the formula, we require the following notation.
  • For $n\geq 1$, let $\mathcal{M}_{g,n}^{\mathrm{ct}}(\pi,1)$ be the space of stable maps to the universal family $\pi$ with compact type domain curves. Stability of the map implies stability of the $n$-pointed domain curve because there are no nonconstant morphisms $\mathbb{P}^{1}\to E$ to any elliptic curve $E$.
  • Let $\tau$ denote the forgetful map
    $\tau:\mathcal{M}_{g,n}^{\mathrm{ct}}(\pi,1)\to\mathcal{M}_{g,n}^{\mathrm{ct}}\,.$
    No stabilization is needed to define $\tau$. The dependence on $n$ is dropped in the notation.
  • There are $n$ evaluation maps
    $\operatorname{ev}_{i}:\mathcal{M}_{g,n}^{\mathrm{ct}}(\pi,1)\longrightarrow\mathcal{E}\,$,
    for $i=1,\ldots,n$, which can be combined as follows. For any $s\times n$ matrix $\mathsf{A}$} as above, we obtain a morphism
    $\operatorname{ev}_{\mathsf{A}}:\mathcal{M}_{g,n}^{\mathrm{ct}}(\pi,1)\longrightarrow\underbrace{\mathcal{E}\times_{\mathcal{M}_{1,1}}\times{\cdots}\times_{\mathcal{M}_{1,1}}\mathcal{E}}_{s}\cong\mathcal{X}_{1}^{s}$
    given by
    (56) $\operatorname{ev}_{\mathsf{A}}(f:C\to E,p_{1},\ldots,p_{n})=\left(E,\sum_{j=1}^{n}a_{1j}f(p_{j}),\ldots,\sum_{j=1}^{n}a_{sj}f(p_{j})\right)\,.$
  • For $n\geq 1$, let $\mathcal{Z}_{n}=\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)$ be the substack corresponding to maps satisfying $f(p_{1})=0$.
  • The moduli space $\mathcal{M}_{g,n}^{\mathrm{ct}}(\pi,1)$ carries a virtual fundamental class of dimension 2g1+n2g-1+n from Gromov-Witten theory. Using the diagram
    $\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)$$\mathcal{M}_{g,n}^{\mathrm{ct}}(\pi,1)$$\mathcal{M}_{1,1}$},{\ $\mathcal{E}\,$$\scriptstyle{\operatorname{ev}_{1}}$$\scriptstyle{\mathsf{z}}$
    we endow the space $\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)$ with a virtual fundamental class of dimension 2g2+n2g-2+n given by the Gysin pullback
    (57) $\left[\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)\right]^{\mathrm{vir}}={\mathsf{z}}^{!}\left[\mathcal{M}_{g,n}^{\mathrm{ct}}(\pi,1)\right]^{\mathrm{vir}},$
    where $\mathsf{z}$ denotes the zero section of the family $\pi$.
The classes (57) are compatible with the forgetful morphisms $q:\mathcal{M}_{g,n}^{\mathrm{ct}}\to\mathcal{M}_{g,n-1}^{\mathrm{ct}},$ for $n\geq 2$. More precisely a formal computation using the definitions and commutativity of Gysin pullbacks shows
(58) $q^{!}\left[\mathcal{M}_{g,n-1}^{\mathrm{ct},0}(\pi,1)\right]^{\mathrm{vir}}=\left[\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)\right]^{\mathrm{vir}}\in\mathsf{CH}_{2g-2+n}(\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1))\,.$
Consider the Torelli map $\mathsf{Tor}_{n}:\mathcal{M}^{\mathrm{ct}}_{g,n}\longrightarrow\mathcal{A}_{g}$ for $n\geq 1$. We have the following diagram
$\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)$$\mathcal{A}_{1}\times\mathcal{A}_{g-1}$$\mathcal{M}_{g,1}^{\mathrm{ct},0}(\pi,1)$$\mathcal{M}_{g,n}^{\mathrm{ct}}$$\mathcal{A}_{g}$$\mathcal{M}_{g,1}^{\mathrm{ct}}$$\scriptstyle{q}$$\scriptstyle{\quad\quad\quad\mathsf{Tor}_{n}}$$\scriptstyle{\mathsf{Tor}_{1}}$
where $q:\mathcal{M}_{g,n}^{\mathrm{ct}}\to\mathcal{M}_{g,1}^{\mathrm{ct}}$ denotes the forgetful morphism. The left front cell is easily seen to be Cartesian. The right front cell is also Cartesian by [10, Proposition 21]. Furthermore,
$\left[\mathcal{M}_{g,1}^{\mathrm{ct}}(\pi,1)\right]^{\mathrm{vir}}=\mathsf{Tor}_{1}^{!}([\mathcal{A}_{1}\times\mathcal{A}_{g-1}])\in\mathsf{CH}_{2g-1}(\mathcal{M}^{\mathrm{ct},0}_{g,1}(\pi,1))\,$,
by [31]. A similar identity holds for all $n\geq 1$:
(59) $\mathsf{Tor}_{n}^{!}([\mathcal{A}_{1}\times\mathcal{A}_{g-1}])=q^{!}\mathsf{Tor}_{1}^{!}([\mathcal{A}_{1}\times\mathcal{A}_{g-1}])=q^{!}\left[\mathcal{M}_{g,1}^{\mathrm{ct},0}(\pi,1)\right]^{\mathrm{vir}}=\left[\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)\right]^{\mathrm{vir}}\,$,
where (58) was used for the last equality.
Lemma 45.
The Abel-Jacobi pullbacks are related to Gromov-Witten theory by:
$\operatorname{aj}_{\mathsf{A}}^{*}\left(\mathsf{p}_{*}([\alpha\times\mathcal{X}_{g-1}^{s}])\right)=\tau_{*}\left(\operatorname{ev}_{\mathsf{A}}^{*}(\alpha)\cap[\mathcal{M}^{\mathrm{ct},0}_{g,n}(\pi,1)]^{\mathrm{vir}}\right)\,\in\mathsf{CH}^{g-1+s}(\mathcal{M}_{g,n}^{\mathrm{ct}})\,.$
Proof.
The proof is a formal consequence of the compatibility between the basic operations in intersection theory. Consider the following fiber diagram:
[…]
Using [29, Theorem 6.2(a), Proposition 6.4], we find
$\operatorname{aj}_{\mathsf{A}}^{*}\left(\mathsf{p}_{*}([\alpha\times\mathcal{X}_{g-1}^{s}])\right)$ $=\tau_{*}\left((\text{pr}\circ j)^{*}\alpha\cap\operatorname{aj}_{\mathsf{A}}^{!}([\mathcal{X}_{1}^{s}\times\mathcal{X}_{g-1}^{s}])\right)$
$=\tau_{*}\left((\text{pr}\circ j)^{*}\alpha\cap\mathsf{Tor}_{n}^{!}([\mathcal{A}_{1}\times\mathcal{A}_{g-1}])\right)$
$=\tau_{*}\left((\text{pr}\circ j)^{*}\alpha\cap[\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)]^{\mathrm{vir}}\right)$
$=\tau_{*}\left(\operatorname{ev}_{\mathsf{A}}^{*}\alpha\cap[\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)]^{\mathrm{vir}}\right)\,.$
Here, we have used (59) and $\text{pr}\circ j=\operatorname{ev}_{\mathsf{A}}$. ∎
Proposition 46.
For every $\alpha\in\mathsf{R}^{*}(\mathcal{X}_{1}^{s})$, the Abel-Jacobi pullback $\operatorname{aj}_{\mathsf{A}}^{*}\left(\mathsf{p}_{*}\left([\alpha\times\mathcal{X}_{g-1}^{s}]\right)\right)$ belongs to the tautological ring of $\mathcal{M}^{\mathrm{ct}}_{g,n}$.
First Proof of Proposition 46. We use the analysis of the Torelli pullback in Section 3. Consider the fiber diagram
(61) $\mathcal{Z}_{n}$$\mathcal{A}_{1}\times\mathcal{A}_{g-1}$$\mathcal{Z}$$\mathcal{M}_{g,n}^{\mathrm{ct}}$$\mathcal{A}_{g}$$\mathcal{M}_{g}^{\mathrm{ct}}$$\scriptstyle{\tau}$$\scriptstyle{\tau}$$\scriptstyle{q}$$\scriptstyle{\quad\quad\quad\mathsf{Tor}_{n}}$$\scriptstyle{\mathsf{Tor}}$
where $q$ is the forgetful map. By Theorem 5, the results of [10] show
$\mathsf{Tor}^{!}[\mathcal{A}_{1}\times\mathcal{A}_{g-1}]=\sum_{\mathsf{T}}\frac{1}{|\operatorname{{\mathsf{Aut}}}(\mathsf{T})|}\left(\iota_{\mathsf{T}}\right)_{*}\mathsf{Cont}_{\mathsf{T}}\,$,
where $\iota_{\mathsf{T}}:\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}}\to\mathcal{Z}$ are the strata of the fibered product $\mathcal{Z}$ indexed by extremal rooted trees with a genus $1$ root, and
$\mathsf{Cont}_{\mathsf{T}}\in\mathsf{R}^{*}(\mathcal{M}_{\mathsf{T}}^{\mathrm{ct}})\,.$
Of course, $\tau\circ\iota_{\mathsf{T}}=\xi_{\mathsf{T}}$ is the boundary map in $\mathcal{M}_{g}^{\mathrm{ct}}$.
The strata of the fibered product $\mathcal{Z}_{n}$ are indexed by extremal rooted trees $\overline{\mathsf{T}}$ with $n$ leaves and a genus $1$ root, whose stabilization after forgetting the leaves is a rooted tree $\mathsf{T}$ as before. Then
(62) $\mathsf{Tor}_{n}^{!}[\mathcal{A}_{1}\times\mathcal{A}_{g-1}]$ $=q^{!}\mathsf{Tor}^{!}[\mathcal{A}_{1}\times\mathcal{A}_{g-1}]=\sum_{\mathsf{T}}\frac{1}{|\operatorname{{\mathsf{Aut}}}(\mathsf{T})|}q^{!}\left(\iota_{\mathsf{T}}\right)_{*}\mathsf{Cont}_{\mathsf{T}}$
$=\sum_{\overline{\mathsf{T}}}\frac{1}{|\operatorname{{\mathsf{Aut}}}(\overline{\mathsf{T}})|}\left(\iota_{\overline{\mathsf{T}}}\right)_{*}\mathsf{Cont}_{\overline{\mathsf{T}}}\,$,
for tautological classes $\mathsf{Cont}_{\overline{\mathsf{T}}}\in\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{\overline{\mathsf{T}}})$ obtained by pulling back $\mathsf{Cont}_{\mathsf{T}}\in\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{\mathsf{T}})$ under the tautological maps. Here, $\iota_{\overline{\mathsf{T}}}:\mathcal{M}_{\overline{\mathsf{T}}}\to\mathcal{Z}_{n}$ is the natural inclusion, so that
$\tau\circ\iota_{\overline{\mathsf{T}}}=\xi_{{\overline{\mathsf{T}}}}$
is the boundary map in $\mathcal{M}_{g,n}^{\mathrm{ct}}.$  [^30] By (59), we see that for $n\geq 1$, the left hand side of (62) equals $\mathcal{Z}_{n}]^{\mathrm{vir}},$ under the identification $\mathcal{Z}_{n}\cong\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1).$
Fix a graph $\overline{\mathsf{T}}$, and let $m$ be the valence of the root (the number of edges and leaves). The root corresponds to a genus $1$ curve with $m$ markings. The marking set $\{1,2,\ldots,n\}$ is canonically partitioned into subsets $N_{1},\ldots,N_{m}$, where $N_{j}$ is the singleton set corresponding to a leaf attached to the root or $N_{j}$ is the set of markings which lie on components attached to the root elliptic curve at a given node. Let ${\mathsf{A}}_{\overline{\mathsf{T}}}$ be the $s\times m$ matrix with entries
(63) $\left(\mathsf{A}_{\overline{\mathsf{T}}}\right)_{ij}=\sum_{k\in N_{j}}a_{ik}\,.$
Since $\mathsf{A}$ has vanishing row sums, $\mathsf{A}_{\mathsf{\overline{T}}}$ also has vanishing row sums. We construct the associated Abel-Jacobi map $\mathcal{M}_{1,m}^{\mathrm{ct}}\to\mathcal{X}_{1}^{s}.$ Composing with the natural projection $\mathcal{M}^{\mathrm{ct}}_{\overline{\mathsf{T}}}\to\mathcal{M}_{1,m}^{\mathrm{ct}}$, we obtain a morphism
$\operatorname{aj}_{\overline{\mathsf{T}}}:\mathcal{M}^{\mathrm{ct}}_{\overline{\mathsf{T}}}\to\mathcal{X}_{1}^{s}.$
Finally, consider the diagram obtained by putting together (60) and (61):
[…]
We intersect (62) with $\operatorname{ev}_{\mathsf{A}}^{*}\alpha=j^{*}\text{pr}^{*}\alpha$. Using Lemma 45, equation (59), and the projection formula, we obtain
$\operatorname{aj}_{\mathsf{A}}^{*}\left(\mathsf{p}_{*}\left([\alpha\times\mathcal{X}_{g-1}^{s}]\right)\right)$ $=\tau_{*}(\operatorname{ev}_{\mathsf{A}}^{*}\alpha\cap\mathsf{Tor}_{n}^{!}([\mathcal{A}_{1}\times\mathcal{A}_{g-1}]))=\sum_{\overline{\mathsf{T}}}\frac{1}{|\operatorname{{\mathsf{Aut}}}(\overline{\mathsf{T}})|}\tau_{*}\left(\iota_{\overline{\mathsf{T}}}\right)_{*}\left(\iota_{\overline{\mathsf{T}}}^{*}\operatorname{ev}_{\mathsf{A}}^{*}(\alpha)\cap\mathsf{Cont}_{\overline{\mathsf{T}}}\right)$
$=\sum_{\overline{\mathsf{T}}}\frac{1}{|\operatorname{{\mathsf{Aut}}}(\overline{\mathsf{T}})|}\left(\xi_{\overline{\mathsf{T}}}\right)_{*}\left(\operatorname{aj}_{\overline{\mathsf{T}}}^{*}(\alpha)\cap\mathsf{Cont}_{\overline{\mathsf{T}}}\right)\,.$
The latter expression is tautological on $\mathcal{M}_{g,n}^{\mathrm{ct}}$ since the Abel-Jacobi pullbacks preserve tautological classes by (55).
Second Proof of Proposition 46. The starting point of the second proof is Lemma 45. The wall-crossing formula established in [51] is also valid for the class
(64) $\tau_{*}\left(\operatorname{ev}_{\mathsf{A}}^{*}(\alpha)\cap[\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)]^{\mathrm{vir}}\right)\,$
with insertions. The resulting expression is a sum over star-shaped graphs as in Section 4.5, where the root has genus $g_{0}=1$, and the partitions $\mu^{i}=(1)$ for all $i$. The difference is that the graphs are also allowed to carry leaves on non-root vertices, corresponding to the marked points $p_{1},\ldots,p_{n}$. We denote the set of such graphs by $\mathsf{Star}(g,n)$. For each star graph $\mathsf{S}$ with $m$ non-root vertices, let
$N_{j}\subset\{1,\dots,n\}\,,\quad j\in\{1,\ldots,m\},$
be the (possibly empty) set of leaves on the $i$-th vertex. Define the $s\times m$ matrix ${\mathsf{A}}_{\mathsf{S}}$ by
(65) $\left({\mathsf{A}}_{\mathsf{S}}\right)_{ij}=\sum_{k\in N_{j}}a_{ik}\,.$
We have an isomorphism
$\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))\cong\mathcal{M}^{\mathrm{ct}}_{1,m}\,.$
Under the above identification, the evaluation morphism from the root component,
$\operatorname{ev}_{\mathsf{S}}\colon\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))\longrightarrow\mathcal{E}^{s}=\mathcal{X}_{1}^{s},$
is easily seen to be the Abel-Jacobi map $\mathcal{M}_{1,m}^{\mathrm{ct}}\to\mathcal{X}_{1}^{s}$ associated to the matrix $\mathsf{A}_{\mathsf{S}}$}. We have $\operatorname{ev}_{\mathsf{S}}=\xi_{\mathsf{S}}\circ\operatorname{ev}_{\mathsf{A}}.$
The analogue of Theorem 6 here takes the form
$\operatorname{ev}_{\mathsf{A}}^{*}(\alpha)\cap[\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)]^{\mathrm{vir}}\\ =\sum_{\mathsf{S}\in\mathsf{Star}(g,n)}\frac{1}{|\operatorname{Aut}({\mathsf{S}})|}(\xi_{\mathsf{S}})_{*}\left(\mathrm{ev}_{\mathsf{S}}^{*}(\alpha)\cap[\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))]^{\mathrm{vir}}\boxtimes\prod^{m}_{i=1}I_{g_{i},N_{i}}(-\Psi_{i})\right)\,$,
where $\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))]^{\mathrm{vir}}$ equals the fundamental class $\mathcal{M}_{1,m}^{\mathrm{ct}}]$, and $I_{g_{i},N_{i}}(z)$ is the marked $I$-function, see [49, Proposition 1] for the exact expression in terms of tautological classes. As proven in Section 6.2, the Abel-Jacobi pullbacks preserve tautological classes
$\operatorname{ev}_{\mathsf{S}}^{*}\alpha\in\mathsf{R}^{*}(\mathcal{M}_{1,m}^{\mathrm{ct}})\,.$
Therefore, after pushing forward under $\tau:\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)\to\mathcal{M}^{\mathrm{ct}}_{g,n}$, the class (64) is tautological on $\mathcal{M}_{g,n}^{\mathrm{ct}}$. ∎
6.4. The Gorenstein kernels and generalized products
The $\lambda_{g}$-pairing on $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ is
$\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\times\mathsf{R}^{2g-3+n-*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\to\mathbb{Q}\,,\,\quad(\alpha,\beta)\mapsto\int_{\overline{\mathcal{M}}_{g,n}}\overline{\alpha}\cdot\overline{\beta}\cdot\lambda_{g}\,.$
Contrary to the situation on $\mathsf{R}^{*}(\mathcal{A}_{g})$ and $\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$, the $\lambda_{g}$-pairing is not in general perfect on $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$. The first failure was found in $\mathsf{R}^{5}(\mathcal{M}_{2,8}^{\mathrm{ct}})$ by Petersen [58], and failures in $\mathsf{R}^{5}(\mathcal{M}_{6}^{\mathrm{ct}})$ and $\mathsf{R}^{5}(\mathcal{M}^{\mathrm{ct}}_{5,2})$ were predicted by Pixton in [59]. By recent results of Canning-Larson-Schmitt [8], the $\lambda_{g}$-pairing fails to be perfect whenever $2g+n\geq 12$ and $g\geq 2$. The strategy of [8] is to establish the non-vanishing of the Gorenstein kernels in
(66) $\mathsf{R}^{5}(\mathcal{M}^{\mathrm{ct}}_{6}),\,\,\,\mathsf{R}^{5}(\mathcal{M}^{\mathrm{ct}}_{5,2}),\,\,\,\,\mathsf{R}^{5}(\mathcal{M}^{\mathrm{ct}}_{4,4}),\,\,\,\mathsf{R}^{5}(\mathcal{M}^{\mathrm{ct}}_{3,6}),\,\,\,\,\mathsf{R}^{5}(\mathcal{M}^{\mathrm{ct}}_{2,8})\,$,
and then to propagate these Gorenstein kernels to the other pairs $(g,n)$ using properties (i)-(v) stated in Section 1.11.
An explicit geometric class in the Gorenstein kernel was constructed in [10]: the non-zero class
$\mathsf{Tor}^{*}\Big([\mathcal{A}_{1}\times\mathcal{A}_{5}]-\mathsf{taut}([\mathcal{A}_{1}\times\mathcal{A}_{5}])\Big)\in\mathsf{R}^{5}(\mathcal{M}_{6}^{\mathrm{ct}})$
generates $\mathsf{K}_{6}\subset\mathsf{R}^{*}(\mathcal{M}_{6}^{\mathrm{ct}})$.
We would like to explain geometrically the failure of the $\lambda_{g}$-pairing to be perfect for the remaining cases in (66) using generalized products $\operatorname{PR}_{g,s}$ and their Abel-Jacobi pullbacks. Consider the Abel-Jacobi map $\operatorname{aj}_{\mathsf{B}}$ defined by (54), and let
$\Delta_{g,s}=\operatorname{aj}_{\mathsf{B}}^{*}\Big([\operatorname{PR}_{g,s}]-\mathsf{taut}^{s}([\operatorname{PR}_{g,s}])\Big)\in\mathsf{R}^{g-1+s}(\mathcal{M}^{\mathrm{ct}}_{g,2s})\,.$
Theorem 47.
The class $\Delta_{g,s}$ satisfies the following properties:
  • (i)
    $\Delta_{g,s}\in\mathsf{K}_{g,2s}$ for all $g$ and $s\in\{0,1\}$,
  • (ii)
    $\Delta_{g,s}\neq 0$ for $(g,s)\in\{{(6,0)},(5,1)\}\,.$
By [8, Theorem 7], $\mathsf{K}_{5,2}$ is 1-dimensional. Therefore, Theorem 11 is a consequence of parts (i) and (ii) of Theorem 47 for $\Delta_{5,1}\in\mathsf{K}_{5,2}$. The only obstruction to cover all the cases of (66) is computational. At the moment, the remaining cases require too much time and memory to compute.
As a consequence of the non-vanishings in part (ii) of Theorem 47, we obtain the following result.
Corollary 48.
The class $\operatorname{PR}_{5,1}]\in\mathsf{CH}^{5}(\mathcal{X}_{5})$ is not tautological.
We have also checked the vanishings
$\Delta_{2,2}=0\in\mathsf{CH}^{3}(\mathcal{M}^{\mathrm{ct}}_{2,4})\,,\quad\Delta_{2,3}=0\in\mathsf{CH}^{4}(\mathcal{M}^{\mathrm{ct}}_{2,6})\,,\quad\Delta_{3,2}=0\in\mathsf{CH}^{4}(\mathcal{M}^{\mathrm{ct}}_{3,4})\,$,
which are consistent with the vanishings of the Gorenstein kernels in these cases proven in [8].
For every $s\times n$ matrix $\mathsf{A}$}, we define
$\Delta_{g,\mathsf{A}}=\operatorname{aj}_{\mathsf{A}}^{*}\Big([\operatorname{PR}_{g,s}]-\mathsf{taut}^{s}([\operatorname{PR}_{g,s}])\Big)\in\mathsf{R}^{g-1+s}(\mathcal{M}^{\mathrm{ct}}_{g,n})\,.$
Based on the calculations of Theorem 47, we propose the following general statement about the Gorenstein kernels of $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$.
Conjecture 49.
The Gorenstein kernels $\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ satisfy:
  • $(\mathsf{P1})$
    $\Delta_{g,{\mathsf{A}}}\in\mathsf{K}_{g,n}$ for all $g$ and all $s\times n$ matrices $\mathsf{A}$}.
  • $(\mathsf{P2})$
    $\{\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\}_{g,n}$ is the smallest system of ideals which is closed under properties (i)-(v) of Section 1.11 and for which $(\mathsf{P1})$ holds.
We now turn to the proof of Theorem 47 (i). The verification of (ii) is through $\mathsf{admcycles}$, and the corresponding code can be found in [27].
Proof of Theorem 47 (i).
The case $s=0$ corresponds to [10, Theorem 4], so we need only consider $s=1$. Then
$\Delta_{g,1}=\operatorname{aj}^{*}\Big([\operatorname{PR}_{g,1}]-\frac{1}{6|B_{2g}|}\lambda_{g-1}\cdot\theta\Big)\in\mathsf{R}^{g}(\mathcal{M}_{g,2}^{\mathrm{ct}})\,$,
for the Abel-Jacobi map
$\operatorname{aj}:\mathcal{M}_{g,2}^{\mathrm{ct}}\to\mathcal{X}_{g},\quad(C,p_{1},p_{2})\mapsto(\mathsf{Jac}(C),\mathcal{O}(p_{1}-p_{2}))\,.$
By Lemma 45 and the definition of $\mathcal{M}_{g,2}^{\mathrm{ct},0}(\pi,1)]^{\mathrm{vir}}$, we obtain that
$\operatorname{aj}^{*}[\operatorname{PR}_{g,1}]$ $=\tau_{*}(\operatorname{ev}_{2}^{*}(0)\cap[\mathcal{M}_{g,2}^{\mathrm{ct},0}(\pi,1)]^{\mathrm{vir}})$
$=\tau_{*}(\operatorname{ev}_{1}^{*}(0)\cdot\operatorname{ev}_{2}^{*}(0)\cap[\mathcal{M}^{\mathrm{ct}}_{g,{2}}(\pi,1)]^{\mathrm{vir}})\in\mathsf{R}^{g}(\mathcal{M}_{g,2}^{\mathrm{ct}})\,.$
By [39, Proposition 7], we have
$\lambda_{g}\cap[\overline{\mathcal{M}}_{g,2}(\pi,1)]^{\mathrm{vir}}=\frac{1}{24}\lambda_{g-1}\cap[\overline{\mathcal{M}}_{g,2}(E,1)]^{\mathrm{vir}}\,$,
where $E$ is a fixed elliptic curve. We conclude
$\overline{\operatorname{aj}^{*}([\operatorname{PR}_{g,1}])}\cdot\lambda_{g}=\frac{1}{24}\tau_{*}(\operatorname{ev}_{1}^{*}(0)\cdot\operatorname{ev}_{2}^{*}(0)\cap[\overline{\mathcal{M}}_{g,{2}}(E,1)]^{\mathrm{vir}})\cdot\lambda_{g-1}=\frac{1}{12|B_{2g}|}(\psi_{1}+\psi_{2})\lambda_{g}\lambda_{g-1}\,$,
where the last equality is a special case of [53, Proposition 6.8].
On the other hand,
$\overline{\operatorname{aj}^{*}\left(\frac{1}{6|B_{2g}|}\cdot\theta\cdot\lambda_{g-1}\right)}\cdot\lambda_{g}=\frac{1}{6|B_{2g}|}\cdot\frac{1}{2}\left(\psi_{1}+\psi_{2}-\sum_{h=1}^{g-1}\delta_{h,\{p_{1}\}}\right)\lambda_{g}\lambda_{g-1}\,$,
where (55) was used in the last equality. The class $\lambda_{g}\lambda_{g-1}$ annihilates the boundary contributions $\delta_{h,\{p_{1}\}}$ as can be seen from the Mumford relations $\lambda^{2}_{h}=0$ and $\lambda_{g-h}^{2}=0$.
We therefore obtain
(67) $\lambda_{g}\cdot\overline{\Delta}_{g,1}=0\in\mathsf{R}^{2g}(\overline{\mathcal{M}}_{g,2})\,$,
which immediately implies that the class $\Delta_{g,1}\in\mathsf{R}^{g}(\mathcal{M}_{g,2}^{\mathrm{ct}})$ is in the kernel of the $\lambda_{g}$-pairing. ∎
Theorem 47 can be used to obtain cycles on the universal abelian variety that satisfy the homomorphism property (51). The Abel-Jacobi map
$\operatorname{aj}:\mathcal{M}_{g,2}^{\mathrm{ct}}\to\mathcal{X}_{g},\quad(C,p_{1},p_{2})\to(\textsf{Jac}(C),\mathcal{O}(p_{1}-p_{2}))$
is proper. The Abel-Jacobi cycle is defined as the pushforward
${\textnormal{AJ}}_{g}]=\operatorname{aj}_{*}([\mathcal{M}_{g,2}^{\mathrm{ct}}])\in\mathsf{CH}^{\binom{g-1}{2}}(\mathcal{X}_{g})\,.$
Corollary 50.
The pair $([{\textnormal{AJ}}_{g}],[\operatorname{PR}_{g,1}])$ satisfies the homomorphism property on $\mathcal{X}_{g}$.
We expect this result to hold more generally for pairs of the form
$([{\textnormal{AJ}}_{g,\mathsf{A}}],[\operatorname{PR}_{g,s}])$
on $\mathcal{X}_{g}^{s}$, where ${\textnormal{AJ}}_{g,\mathsf{A}}]\in\mathsf{CH}^{*}(\mathcal{X}_{g}^{s})$ denotes the Abel-Jacobi cycle for arbitrary $g$ and $\mathsf{A}.$
Proof.
We must verify the equality
$\mathsf{taut}([{\textnormal{AJ}}_{g}]\cdot[\operatorname{PR}_{g,1}])=\mathsf{taut}([{\textnormal{AJ}}_{g}])\cdot\mathsf{taut}([\operatorname{PR}_{g,1}])\in\mathsf{R}^{*}(\mathcal{X}_{g})\,.$
For brevity, we omit here the superscripts, writing $\mathsf{taut}$ for the tautological projection on the universal abelian variety $\mathcal{X}_{g}$. Equivalently, we will prove
$\mathsf{taut}\left([{\textnormal{AJ}}_{g}]\cdot\big([\operatorname{PR}_{g,1}]-\mathsf{taut}([\operatorname{PR}_{g,1}])\big)\right)=0\in\mathsf{R}^{*}(\mathcal{X}_{g})\,.$
By definition of tautological projection, the above vanishing amounts to showing
(68) $\int_{\overline{\mathcal{X}}_{g}}\overline{[{\textnormal{AJ}}_{g}]}\cdot\overline{[\operatorname{PR}_{g,1}]-\mathsf{taut}([\operatorname{PR}_{g,1}])}\cdot\overline{\alpha}\cdot\lambda_{g}=0$
for all tautological classes $\alpha\in\mathsf{R}^{*}(\mathcal{X}_{g})$. We fix here any smooth compactification $\overline{\mathcal{X}}_{g}$ of $\mathcal{X}_{g}$.
The crux of the argument is the vanishing
(69) $\overline{\operatorname{aj}^{*}([\operatorname{PR}_{g,1}]-\mathsf{taut}([\operatorname{PR}_{g,1}]))}\cdot\lambda_{g}=0\in\mathsf{R}^{2g}(\overline{\mathcal{M}}_{g,2})$
established in (67). A minor complication is that the map
$\operatorname{aj}:\mathcal{M}_{g,2}^{\mathrm{ct}}\to\mathcal{X}_{g}\hookrightarrow\overline{\mathcal{X}_{g}}$
does not extend to $\overline{\mathcal{M}}_{g,2}.$ However, such an extension always exists after suitable blowups, yielding morphisms
$\pi:\widetilde{\mathcal{M}}\to\overline{\mathcal{M}}_{g,2}\,,\quad\widetilde{\operatorname{aj}}:\widetilde{\mathcal{M}}\to\overline{\mathcal{X}}_{g}\,.$
We may assume $\pi$ is an isomorphism over the compact type locus $\mathcal{M}_{g,2}^{\mathrm{ct}}$, which we view as embedded in $\widetilde{\mathcal{M}}$. Furthermore, we may take $\widetilde{\mathcal{M}}$ to be a nonsingular proper Deligne-Mumford stack.
[…]
$\overline{[{\textnormal{AJ}}_{g}]}=\widetilde{\operatorname{aj}}_{*}([\widetilde{\mathcal{M}}])+\text{classes supported on }\overline{\mathcal{X}}_{g}\smallsetminus\mathcal{X}_{g}\,.$
Since $\lambda_{g}$ annihilates the boundary terms, we can rewrite the left hand side of (68) as
(70) $\int_{\overline{\mathcal{X}}_{g}}\widetilde{\operatorname{aj}}_{*}([\widetilde{\mathcal{M}}])\cdot\overline{[\operatorname{PR}_{g,1}]-\mathsf{taut}([\operatorname{PR}_{g,1}])}\cdot\overline{\alpha}\cdot\lambda_{g}=\int_{\widetilde{\mathcal{M}}}\widetilde{\operatorname{aj}}^{*}\overline{[\operatorname{PR}_{g,1}]-\mathsf{taut}([\operatorname{PR}_{g,1}])}\cdot\widetilde{\operatorname{aj}}^{*}{\overline{\alpha}}\cdot\lambda_{g}\,.$
The cycles
$\widetilde{\operatorname{aj}}^{*}\overline{[\operatorname{PR}_{g,1}]-\mathsf{taut}([\operatorname{PR}_{g,1}])}\ \text{ and }\pi^{*}\overline{\operatorname{aj}^{*}([\operatorname{PR}_{g,1}]-\mathsf{taut}([\operatorname{PR}_{g,1}]))}$
agree on $\mathcal{M}_{g,2}^{\mathrm{ct}}\hookrightarrow\widetilde{\mathcal{M}}$. Using again that $\lambda_{g}=0$ kills the boundary contributions, we rewrite (70) as
$\int_{\widetilde{\mathcal{M}}}\pi^{*}\overline{\operatorname{aj}^{*}([\operatorname{PR}_{g,1}]-\mathsf{taut}([\operatorname{PR}_{g,1}]))}\cdot\widetilde{\operatorname{aj}}^{*}{\overline{\alpha}}\cdot\lambda_{g}=0\,$,
as needed. The last equality follows by pulling back (69) to $\widetilde{\mathcal{M}}$. ∎

### v2

@@
 $\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\rightarrow\mathsf{R}^{*}(\mathcal{A}_{g})$
-was constructed in [9]. We present here new calculations of intersection products of the Torelli locus in $\mathcal{A}_{g}$ with the product loci
-$\mathcal{A}_{r}\times\mathcal{A}_{g-r}\rightarrow\mathcal{A}_{g}$
-for $r\leq 3$.
+was constructed in [9].
+We present here new calculations of intersection products of the Torelli locus in $\mathcal{A}_{g}$ with the product loci $\mathcal{A}_{r}\times\mathcal{A}_{g-r}\rightarrow\mathcal{A}_{g}$ for $r\leq 3$.
 The results suggest that $\mathsf{taut}$ is a $\mathbb{Q}$-algebra homomorphism (at least for special cycles).
@@
 Using Abel-Jacobi pullbacks of product cycles on $\mathcal{X}_{g}^{s}$ and their projections, we construct a new family of classes which we conjecture to lie in the Gorenstein kernels of the tautological rings $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g,n})$.
-In particular, we construct a nontrivial element of the Gorenstein kernel of $\mathsf{R}^{5}(\mathcal{M}_{5,2}^{\mathrm{ct}})$.
-Whether these classes generate the Gorenstein kernels of all $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ is an intriguing open question.
+In particular, we construct nontrivial elements of the Gorenstein kernels of $\mathsf{R}^{5}(\mathcal{M}_{5,2}^{\mathrm{ct}})$ and $\mathsf{R}^{5}(\mathcal{M}_{4,4}^{\mathrm{ct}})$.
+Whether all of these classes generate the Gorenstein kernels of all $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ is an intriguing open question.
 1. Introduction
@@
 By a result of Borel [6], the rational group cohomology of $\operatorname{Sp}_{2g}(\mathbb{Z})$ stabilizes to a polynomial algebra in the odd lambda classes:
-$\lim_{g\to\infty}\mathsf{H}^{*}(\operatorname{Sp}_{2g}(\mathbb{Z}),\mathbb{Q})=\mathbb{Q}[\lambda_{1},\lambda_{3},\ldots]\,.$
+$\lim_{g\to\infty}\mathsf{H}^{*}(\operatorname{Sp}_{2g}(\mathbb{Z}),\mathbb{Q})=\mathbb{Q}[\lambda_{1},\lambda_{3},\lambda_{5},\ldots]\,.$
 Motivated by Borel’s calculation, van der Geer [64] defined the tautological ring
@@
 (2) $\epsilon^{\mathsf{ab}}\left(\prod_{i=1}^{g-1}\lambda_{i}\right)=\int_{\overline{\mathcal{A}}_{g}}\,\prod_{i=1}^{g}\lambda_{i}=\prod_{i=1}^{g}\frac{|B_{2i}|}{4i}\,\,$,
-see [63, page 71].
+see [64, page 71].
 Definition 2 ([9]).
@@
 $-\frac{12533760}{691}\lambda_{2}\lambda_{3}\lambda_{4}\lambda_{6}+\frac{552960}{691}\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}\lambda_{5}\,.$
-Formulas for $\mathsf{taut}([\mathcal{J}_{g}])$ for $g\leq 11$ can be found at [60].
+Formulas for $\mathsf{taut}([\mathcal{J}_{g}])$ for $g\leq 11$ are available at the link [61].
 Since Faber’s seminal paper [24], many closed formulas for Hodge integrals have been found (especially when the integrand includes the top Chern class $\lambda_{g}$), see [19, 20, 21].
@@
 By equation (3), if either $\alpha\in\mathsf{R}^{*}(\mathcal{A}_{g})$ or $\beta\in\mathsf{R}^{*}(\mathcal{A}_{g})$, then the homomorphism property holds for $(\alpha,\beta)$.
-The homomorphism property is therefore an obstruction to being tautological.
+Failure of the homomorphism property is therefore an obstruction to being tautological.
 The homomorphism property has been recently proven to hold in the following nontrivial cases:
@@
   • (iii)
-    If $\beta$ is supported on the Noether-Lefschetz locus of abelian varieties that decompose (up to isogeny) as a product of $m$ elliptic curves and a (gm)(g-m)-dimensional abelian variety, then the homomorphism property holds for $(\alpha,\beta)$ for all $\alpha\in\mathsf{CH}^{*}(\mathcal{A}_{g})$, see [26].
+    If $\beta$ is supported on the Noether-Lefschetz locus of abelian varieties that decompose (up to isogeny) as a product of $m>0$ elliptic curves and a (gm)(g-m)-dimensional abelian variety, then the homomorphism property holds for $(\alpha,\beta)$ for all $\alpha\in\mathsf{CH}^{*}(\mathcal{A}_{g})$, see [26].
 In particular, the homomorphism property holds for $(\alpha,[\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{1}\times\mathcal{A}_{g-m}])$.
@@
 The first is a direct study of the excess intersection geometry of the fiber product of the Torelli and product morphisms following [10].
-The second approach is by recasting the intersection geometry in terms of families Gromov-Witten classes, which are computed by a wall-crossing formula related to unramified maps [51, 49].
+The second approach is by recasting the geometry in terms of families Gromov-Witten classes, which are computed by a wall-crossing formula related to unramified maps [52, 50].
 1.5. The Torelli fiber product
@@
 under the isomorphism (11).
-The wall-crossing formula established in [51, Theorem 6.21] relates the virtual fundamental classes of moduli spaces of stable and unramified maps. Let
+The wall-crossing formula established in [52, Theorem 6.21] relates the virtual fundamental classes of moduli spaces of stable and unramified maps. Let
 $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r})=\mathcal{M}_{g_{0}}^{\mathrm{ct},\mathrm{un}}(\pi_{r},\beta_{\mathrm{min}})/\mathcal{X}_{r}\,$
@@
 $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{1}(\pi_{1},(1),\ldots,(1))^{\circ},$
-where $\mu^{i}=(1)$ for all $i$.
+where $\mu^{i}=(1)$ for all $i$, see [50].
 $\bullet$ For $r=2$, the only moduli spaces of unramified maps which occur are
 (14) $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{2}(\pi_{2},(1),\ldots,(1))^{\circ}\ \ \ \text{and}\ \ \ \mathcal{M}^{\mathrm{ct},\mathrm{un}}_{1}(\pi_{2},(1,1),(1),\ldots,(1))^{\bullet}\,$,
-The partition $\mu=(1,1)$ arises for unramified maps to fibers of $\pi_{2}$ given by products $E_{1}\times E_{2}$ of elliptic curves: the associated domains curves are two disconnected genus 1 curves mapping to a nodal genus 2 curve consisting of two elliptic components.
+The partition $\mu=(1,1)$ arises for unramified maps to fibers of $\pi_{2}$ given by products $E_{1}\times E_{2}$ of elliptic curves: the associated domain curves are two disconnected genus 1 curves mapping to a nodal genus 2 curve consisting of two elliptic components.
 The point on the target corresponding to the node of the genus 2 curve has multiplicity $(1,1)$.
+Since curves in the minimal class $\beta_{\mathrm{min}}$ are at worst nodal, maps to an abelian surface can only have multiplicities $(1)$ and $(1,1)$.
+By admissibility, the curves have degree at most $2$ over rational components of Fulton-MacPherson degenerations of abelian surfaces.
+Degree $2$ curves in $\mathbb{P}^{2}$ are also at worst nodal.
+Hence, the same multiplicities $(1)$ and $(1,1)$ occur for maps to Fulton-MacPherson degenerations.
 A result of [26] exhibits the spaces (14) as an iterated blowup of $\mathcal{M}^{\mathrm{ct}}_{2,n}$ and $\mathcal{M}^{\mathrm{ct}}_{1,n}$ along boundary strata in both the connected and disconnected cases.
@@
 A full analysis is provided in Section 4 with examples.
-$\bullet$ For $r\geq 3$, only partitions of the type $\mu=(1,\ldots,1)$ can occur since the minimal curves can have at worst ordinary multiple point singularities.
-The analysis for $r\geq 3$ is more complex and will not be treated here.
 1.7. Results about the homomorphism property
@@
 $\mathsf{cod}_{\mathcal{J}}=\binom{g+1}{2}-(3g-3)\,.$
-The codimension of the product cycle $\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}$ in $\mathcal{A}_{g}$ is
+The codimension of the product cycle $\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}$ in $\mathcal{A}_{g}$ is
 $\mathsf{cod}_{\underline{g}}=\binom{g+1}{2}-\sum_{i=1}^{k}\binom{g_{i}+1}{2}=\sum_{1\leq i<j\leq k}g_{i}g_{j}\,.$
 The homomorphism property
-(15) $\mathsf{taut}([\mathcal{J}_{g}]\cdot[\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}])\stackrel{{\scriptstyle?}}{{=}}\mathsf{taut}([\mathcal{J}_{g}])\cdot\mathsf{taut}([\mathcal{A}_{g_{1}}\times\ldots\times\mathcal{A}_{g_{k}}])\,\in\,\mathsf{R}^{\mathsf{cod}_{\mathcal{J}}+\mathsf{cod}_{\underline{g}}}(\mathcal{A}_{g})$
+(15) $\mathsf{taut}([\mathcal{J}_{g}]\cdot[\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])\stackrel{{\scriptstyle?}}{{=}}\mathsf{taut}([\mathcal{J}_{g}])\cdot\mathsf{taut}([\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}])\,\in\,\mathsf{R}^{\mathsf{cod}_{\mathcal{J}}+\mathsf{cod}_{\underline{g}}}(\mathcal{A}_{g})$
 is trivial (and true) if
@@
 The widest interpretation of our calculations is that the homomorphism property holds in general.
-Conjecture 8 (Strong form).
+Conjecture 8.
 The $\mathbb{Q}$-linear map $\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\rightarrow\mathsf{R}^{*}(\mathcal{A}_{g})$ is a homomorphism of $\mathbb{Q}$-algebras.
@@
 Unfortunately, the Chow groups of $V_{0}$ have not yet been determined.
-Nevertheless, the existence of $V_{0}$ in characteristic $p$ may be viewed as motivation for the following weakening of Conjecture 8:
-Conjecture 9 ($\overline{\mathbb{Q}}$ form).
-The homomorphism property holds for all classes $\alpha,\beta\in\mathsf{CH}^{*}(\mathcal{A}_{g})$ defined over $\overline{\mathbb{Q}}$.
-Both Conjectures 8 and 9 are rather bold interpretations of the current data.
+A specialization argument shows that the homomorphism property for cycles on $\mathcal{A}_{g}(\overline{\mathbb{F}}_{p})$ for all $p$ implies the homomorphism property for cycles on $\mathcal{A}_{g}(\overline{\mathbb{Q}})$.
+Since every cycle over $\mathbb{C}$ is homologous to a cycle defined over $\overline{\mathbb{Q}}$, Conjecture 8 would then follow.
+Conjecture 8 is a rather bold interpretation of the current data.
 Another possibility is that the homomorphism property holds in the restricted case where at least one of the classes $\alpha$ or $\beta$ is a Noether-Lefschetz cycle or, more generally, a special cycle  [^9] See [47] for a discussion of special cycles in the context of $\mathcal{A}_{g}$..
@@
 If we further assume that the image $\mathsf{Tor}(\overline{\mathcal{M}}_{g})$ lies in the nonsingular locus of $\overline{\mathcal{A}}_{g}$, a property that holds for both perfect cone and second Voronoi [46, Section 5.2], then the following question is well-defined.
-Question 10.
+Question 9.
 Compute $\mathsf{Tor}^{*}([\,\overline{\mathcal{A}_{g_{1}}\times\cdots\times\mathcal{A}_{g_{k}}}\,])\in\mathsf{CH}^{*}(\overline{\mathcal{M}}_{g})\,.$
-At the moment, Question 10 is open.
+At the moment, Question 9 is open.
 In principle, both excess intersection theory and the wall-crossing approach can be applied.
@@
 1.10. The geometry of the universal fiber product $\mathcal{X}^{s}_{g}$
-A difficulty in studying Question 10 and extending Theorem 11 to higher torus rank loci is the complexity of the boundary geometry of toroidal compactifications $\mathcal{A}_{g}\subset\overline{\mathcal{A}}_{g}$.
+A difficulty in studying Question 9 and extending Theorem 10 to higher torus rank loci is the complexity of the boundary geometry of toroidal compactifications $\mathcal{A}_{g}\subset\overline{\mathcal{A}}_{g}$.
 To generate elements of the Gorenstein kernel of $\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g-s,2s})$ for general $s$, a different approach can be pursued using the $s$-fold fiber product
@@
 The construction of $\operatorname{aj}^{*}(\Delta_{g,s})$ for $s>1$ may be viewed as probing the higher torus rank loci of toroidal compactifications.
-That $\operatorname{aj}^{*}(\Delta_{g,s})$ lies in the Gorenstein kernel was proven for $s=0$ in [10] and is proven for $s=1$ in Theorem 47, Section 6.
+That $\operatorname{aj}^{*}(\Delta_{g,s})$ lies in the Gorenstein kernel was proven for $s=0$ in [10] and is proven for $s=1$ in Theorem 48 of Section 6.
+The class $\operatorname{aj}^{*}(\Delta_{4,2})\in\mathsf{R}^{5}(\mathcal{M}_{4,4}^{\mathsf{ct}})$ is proven to be a nontrivial element of the kernel of the $\lambda_{4}$-pairing on $\mathsf{R}^{*}(\mathcal{M}_{4,4}^{\mathrm{ct}})$ in Theorem 48.
 The non-vanishing part of Conjecture 15 is open for all $s\geq 0$.
@@
   2. (ii)
-    $\pi_{i*}:\mathsf{K}_{g,n}\rightarrow\mathsf{K}_{g,n-1}$, where $\pi_{i}:\mathcal{M}_{g,n}^{\mathrm{ct}}\rightarrow\mathcal{M}_{g,n-1}^{\mathrm{ct}}$ is the map forgetting the $i^{th}$ marking,
+    $\rho_{i*}:\mathsf{K}_{g,n}\rightarrow\mathsf{K}_{g,n-1}$, where $\rho_{i}:\mathcal{M}_{g,n}^{\mathrm{ct}}\rightarrow\mathcal{M}_{g,n-1}^{\mathrm{ct}}$ is the map forgetting the $i^{th}$ marking,
   3. (iii)
-    $\pi^{*}_{i}:\mathsf{K}_{g,n-1}\rightarrow\mathsf{K}_{g,n}$,
+    $\rho^{*}_{i}:\mathsf{K}_{g,n-1}\rightarrow\mathsf{K}_{g,n}$,
   4. (iv)
@@
 To state the compatibility with respect to pullback along the gluing map $\xi$, we first define the correspondence map
-$\eta:\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1}+g_{2},n_{1}+n_{2}})\otimes\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})\,\,,\ \ \ \eta(\alpha\otimes\beta)=\pi_{1*}\left(\xi^{*}(\alpha)\cup\pi_{2}^{*}(\beta)\right)\,$,
-where $\pi_{i}:\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1}\times\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1}\rightarrow\mathcal{M}^{\mathrm{ct}}_{g_{i},n_{i}+1}$ are the projections and
-$\pi_{1*}:\mathsf{R}^{*}(\mathcal{M}_{g_{1},n_{1}+1}^{\mathrm{ct}})\times\mathsf{R}^{*}(\mathcal{M}_{g_{2},n_{2}+1}^{\mathrm{ct}})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})$
+$\eta:\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1}+g_{2},n_{1}+n_{2}})\otimes\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})\,\,,\ \ \ \eta(\alpha\otimes\beta)=\rho_{1*}\left(\xi^{*}(\alpha)\cup\rho_{2}^{*}(\beta)\right)\,$,
+where $\rho_{i}:\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1}\times\mathcal{M}^{\mathrm{ct}}_{g_{2},n_{2}+1}\rightarrow\mathcal{M}^{\mathrm{ct}}_{g_{i},n_{i}+1}$ are the projections and
+$\rho_{1*}:\mathsf{R}^{*}(\mathcal{M}_{g_{1},n_{1}+1}^{\mathrm{ct}})\times\mathsf{R}^{*}(\mathcal{M}_{g_{2},n_{2}+1}^{\mathrm{ct}})\rightarrow\mathsf{R}^{*}(\mathcal{M}^{\mathrm{ct}}_{g_{1},n_{1}+1})$
 is defined by the $\lambda_{g_{2}}$-evaluation on the second factor. Then, we have:
@@
 Conjecture 15 motivates a natural question.
-Question 17.
+Question 16.
 Is $\big\{\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\big\}_{g,n}$ the smallest system of ideals that satisfies properties (i)-(v) and the further condition $\operatorname{aj}^{*}(\Delta_{g,s})\in\mathsf{K}_{g,2s}$ ?
@@
 1.12. Acknowledgments
-The calculation of $\mathsf{Tor}^{*}[\mathcal{A}_{2}\times\mathcal{A}_{g-2}]\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$ with the goal of determining whether the homomorphism property holds for $([\mathcal{J}_{g}],[\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ was a group project at ETH Zürich in the Fall term of 2025 with both theoretical and computational sides.
+The calculation of $\mathsf{Tor}^{*}([\mathcal{A}_{2}\times\mathcal{A}_{g-2}])\in\mathsf{R}^{*}(\mathcal{M}_{g}^{\mathrm{ct}})$ with the goal of determining whether the homomorphism property holds for $([\mathcal{J}_{g}],[\mathcal{A}_{2}\times\mathcal{A}_{g-2}])$ was a group project at ETH Zürich in the Fall term of 2025 with both theoretical and computational sides.
 We dedicate our paper to Carel Faber: his work [24] is the starting point of several of the themes developed here, and he pioneered the computational exploration of the tautological ring of the moduli space of curves.
@@
 $\Psi_{i}\in\mathsf{CH}^{1}(\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\bullet})\,.$
-Up to certain normalization terms, these $\Psi$-classes can also be defined via vector bundles associated with the cotangent spaces at the markings, see [50, Section 3.2] for more details.
+Up to certain normalization terms, these $\Psi$-classes can also be defined via vector bundles associated with the cotangent spaces at the markings, see [51, Section 3.2] for more details.
 The same definition applies to $\mathcal{M}^{\mathrm{ct},\mathrm{un}}_{g_{0}}(\pi_{r},\mu^{1},\ldots,\mu^{m})^{\circ}$.
@@
 The following result is proven using classical invariant theory.
-Theorem 32 ([62, Theorem 3.4]).
+Theorem 31 ([63, Theorem 3.4]).
 The ring of invariants has the following presentation
@@
 and, therefore, also holds in the Chow ring of $X^{s}$. The symmetric property of $\theta\in\mathsf{CH}^{1}(X)$ is used here to calculate $\varphi^{*}\theta$ in Chow:
-$\varphi^{*}\theta=\sum_{i=1}^{s}a_{i}^{2}\theta_{i}+\sum_{1\leq i<j\leq s}a_{i}a_{j}\eta_{ij}\in\mathsf{CH}^{1}(X^{s})\,.\qed$
+(42) $\varphi^{*}\theta=\sum_{i=1}^{s}a_{i}^{2}\theta_{i}+\sum_{1\leq i<j\leq s}a_{i}a_{j}\eta_{ij}\in\mathsf{CH}^{1}(X^{s})\,.\qed$
 Proposition 33.
@@
 By Proposition 32, it suffices to show that $\mathfrak{I}_{g,s}$ is Gorenstein.
-We decompose the cohomology $\mathsf{H}^{*}(X^{s},\mathbb{Q})$ into sums of irreducible representations of $\operatorname{Sp}(V),$ with $\mathfrak{I}_{g,s}$ occurring as the invariant summand.
+Since $\operatorname{Sp}(V)$ is reductive, we can decompose the cohomology $\mathsf{H}^{*}(X^{s},\mathbb{Q})$ into sums of irreducible representations, with $\mathfrak{I}_{g,s}$ occurring as the invariant summand.
 The top degree part $\mathsf{H}^{2gs}(X^{s})$ has rank $1$ and is spanned by the $\operatorname{Sp}(V)$-invariant class
@@
 ∎
+5.4. Properties of tautological rings
+By definition, tautological rings are preserved under pullbacks via projections forgetting a factor:
+(43) $\text{pr}:X^{n+1}\to X^{n}\,,\ \ \ \ \text{pr}^{*}:\mathsf{R}^{*}(X^{n})\rightarrow\mathsf{R}^{*}(X^{n+1})\,.$
+By [3, Theorem 9.10], tautological rings are also preserved under the corresponding pushforwards,
+(44) $\text{pr}_{*}:\mathsf{R}^{*}(X^{n+1})\rightarrow\mathsf{R}^{*}(X^{n})\,.$
+In fact, much more general versions of the properties (43) and (44) hold.
+Let $\mathsf{A}=(a_{ij})$ be a matrix of size $t\times s$ with integer coefficients. The matrix $\mathsf{A}$ induces a natural morphism
+(45) $f_{\mathsf{A}}:X^{s}\to X^{t},\quad(x_{1},\ldots,x_{s})\to\left(\sum_{j=1}^{s}a_{1j}x_{j},\ldots,\sum_{j=1}^{s}a_{tj}x_{j}\right)\,.$
+The diagonal inclusion $X\rightarrow X\times X$ is obtained from the case $\mathsf{A}=\begin{pmatrix}1\\ 1\end{pmatrix}\,$.
+Proposition 36.
+The tautological rings are preserved under pushforward,
+$(f_{\mathsf{A}})_{*}:\mathsf{R}^{*}(X^{s})\rightarrow\mathsf{R}^{*}(X^{t})\,.$
+Proof.
+For all classes $\alpha\in\mathsf{CH}^{*}(X^{s})$, we have
+$(f_{\mathsf{A}})_{*}\alpha=\widehat{\text{pr}}_{*}\left({\text{pr}}^{*}\alpha\cdot\left[\Gamma_{\mathsf{A}}\right]\right)\,$,
+where $\Gamma_{\mathsf{A}}\subset X^{s+t}$ is the graph of $f_{\mathsf{A}}$, and $\text{pr},$ $\widehat{\text{pr}}$ are the projections to $X^{s}$ and $X^{t}$ respectively.
+By the pushforward property (44), if $\left[\Gamma_{\mathsf{A}}\right]\in\mathsf{R}^{*}(X^{s+t})$, then $(f_{\mathsf{A}})_{*}\alpha\in\mathsf{R}^{*}(X^{t})$.
+We may further assume $t=1$ by using the following equation:
+$\left[\Gamma_{\mathsf{A}}\right]=\widehat{\text{pr}}_{1}^{*}\left[\Gamma_{1}\right]\cdots\widehat{\text{pr}}_{t}^{*}\left[\Gamma_{t}\right]\,$,
+where $\Gamma_{i}\subset X^{s+1}$ is the graph of the morphism
+$(x_{1},\ldots,x_{s})\to\sum_{j=1}^{s}a_{ij}x_{j}\,.$
+Here, for $1\leq i\leq t$, the projection $\widehat{\text{pr}}_{i}:X^{s+t}\to X^{s+1}$ records the first $s$ coordinates as well as the coordinate in position $s+i$.
+For $t=1$, we write $(a_{1},\ldots,a_{s})$ for the entries of $\mathsf{A}$, and we will prove $\left[\Gamma_{\mathsf{A}}\right]\in\mathsf{R}^{g}(X^{s+1})$. Setting
+$\varphi:X^{s+1}\to X,\quad(x_{1},\ldots,x_{s},x_{s+1})\to x_{s+1}-a_{1}x_{1}-\ldots-a_{s}x_{s}\,$,
+we have
+$\left[\Gamma_{\mathsf{A}}\right]=\varphi^{*}([0])=\varphi^{*}\left(\frac{\theta^{g}}{g!}\right)\,.$
+Finally, to show $\varphi^{*}\theta\in\mathsf{CH}^{1}(X^{s+1})$ is tautological, we use the identity
+$\varphi^{*}\theta=\theta_{s+1}+\sum_{i=1}^{s}a_{i}^{2}\theta_{i}+\sum_{1\leq i<j\leq s}a_{i}a_{j}\eta_{ij}-\sum_{i=1}^{s}a_{i}\eta_{i,s+1}\,$,
+which is a particular case of (42). ∎
+Proposition 37.
+The tautological rings are preserved under pullback,
+$(f_{\mathsf{A}})^{*}:\mathsf{R}^{*}(X^{t})\rightarrow\mathsf{R}^{*}(X^{s})\,.$
+Proof.
+By the definition of $\mathsf{R}^{*}(X^{t})$, we can reduce to the case where $t=1$ or $t=2$. The $t=1$ case is a consequence of (42). For $t=2$, we write
+$\mathsf{A}=\begin{pmatrix}a_{1}&\ldots&a_{s}\\ b_{1}&\ldots&b_{s}\end{pmatrix}\,$,
+and we use (42) and the equation
+$(f_{\mathsf{A}})^{*}(\eta_{12})=\sum_{1\leq i<j\leq s}(a_{i}b_{j}+a_{j}b_{i})\,\eta_{ij}+2\sum_{i=1}^{s}a_{i}b_{i}\,\theta_{i}\,.$
+The latter is easily verified by induction on $s$ and the See-Saw Theorem. ∎
 5.5. The tautological ring of $\mathcal{X}_{g}^{s}$
@@
 generated over $\mathbb{Q}$ by the classes $\theta_{i}$, for $1\leq i\leq s$, and $\eta_{ij}$ for all $1\leq i<j\leq s$ is the vertical tautological ring. ∎
+The proofs of Propositions 36 and 37 carry over to families of abelian varieties, hence the same pushforward and pullback properties also hold for the tautological rings $\mathsf{R}^{*}(\mathcal{X}_{g}^{s})$ and $\mathsf{R}^{*}_{\mathsf{vert}}(\mathcal{X}_{g}^{s})$.
 Let $\overline{\mathcal{X}}_{g}$ be a nonsingular toroidal compactification of $\mathcal{X}_{g}$ lying over a nonsingular toroidal compactification $\overline{\mathcal{A}}_{g}$ of $\mathcal{A}_{g}$.
@@
 For $s=2,$ the vertical tautological ring was studied and shown to be Gorenstein in [34, Theorem 3.3].
-The tautological ring for families of abelian varieties was also considered in [3, Section 9.4] and [32, Section 6], and $\mathsf{R}^{*}(\mathcal{X}^{s}_{g})$ coincides with the definition of [3] for the universal family.
+The tautological ring for families of abelian varieties was also considered in [3, Section 9.4], [32, Section 6] and [48].
+The tautological ring $\mathsf{R}^{*}(\mathcal{X}^{s}_{g})$ coincides with the definition of [3] for the universal family.
 Theorem 39.
@@
   • (i)
-    $X^{s}=X\times_{S}\times\ldots\times_{S}X$ is the $s$-fold fibered product, and $f^{s}:X^{s}\to S$ is the canonical morphism,
+    $X^{s}=X\times_{S}\cdots\times_{S}X$ is the $s$-fold fibered product, and $f^{s}:X^{s}\to S$ is the canonical morphism,
   • (ii)
@@
 Therefore, by Lemma 40, the tautological projection of the cycle
-$\underbrace{(\mathcal{A}_{r}\times\mathcal{X}_{g-r})\times_{\mathcal{A}_{g}}\ldots\times_{\mathcal{A}_{g}}(\mathcal{A}_{r}\times\mathcal{X}_{g-r})}_{s}\to\mathcal{X}_{g}^{s}$
+$\underbrace{(\mathcal{A}_{r}\times\mathcal{X}_{g-r})\times_{\mathcal{A}_{g}}\cdots\times_{\mathcal{A}_{g}}(\mathcal{A}_{r}\times\mathcal{X}_{g-r})}_{s}\to\mathcal{X}_{g}^{s}$
 is given by
@@
 $\mathcal{M}^{\mathrm{ct}}_{g,2s}\rightarrow\mathcal{T}\rightarrow\overline{\mathcal{A}}_{g+s}\,$,
-where $\overline{\mathcal{A}}_{g+s}$ is a toroidal compactification for which $\mathsf{Tor}$ extends and $\mathcal{T}$ is a torus-bundle over $\mathcal{X}_{g}^{s}$.
+where $\overline{\mathcal{A}}_{g+s}$ is a toroidal compactification for which $\mathsf{Tor}$ extends and $\mathcal{T}$ is a torus bundle over $\mathcal{X}_{g}^{s}$.
 The composite $\mathcal{M}^{\mathrm{ct}}_{g,2s}\rightarrow\mathcal{T}\rightarrow\mathcal{X}_{g}^{s}$ is $\operatorname{aj}_{\mathsf{B}}$.
@@
 $\operatorname{ev}_{\mathsf{A}}^{*}(\alpha)\cap[\mathcal{M}_{g,n}^{\mathrm{ct},0}(\pi,1)]^{\mathrm{vir}}\\ =\sum_{\mathsf{S}\in\mathsf{Star}(g,n)}\frac{1}{|\operatorname{Aut}({\mathsf{S}})|}(\xi_{\mathsf{S}})_{*}\left(\mathrm{ev}_{\mathsf{S}}^{*}(\alpha)\cap[\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))]^{\mathrm{vir}}\boxtimes\prod^{m}_{i=1}I_{g_{i},N_{i}}(-\Psi_{i})\right)\,$,
-where $\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))]^{\mathrm{vir}}$ equals the fundamental class $\mathcal{M}_{1,m}^{\mathrm{ct}}]$, and $I_{g_{i},N_{i}}(z)$ is the marked $I$-function, see [49, Proposition 1] for the exact expression in terms of tautological classes.
+where $\mathcal{M}_{1}^{\operatorname{ct,un}}(\pi_{1},(1),\ldots,(1))]^{\mathrm{vir}}$ equals the fundamental class $\mathcal{M}_{1,m}^{\mathrm{ct}}]$, and $I_{g_{i},N_{i}}(z)$ is the marked $I$-function, see [50, Proposition 1] for the exact expression in terms of tautological classes.
 As proven in Section 6.2, the Abel-Jacobi pullbacks preserve tautological classes
@@
   • (ii)
-    $\Delta_{g,s}\neq 0$ for $(g,s)\in\{{(6,0)},(5,1)\}\,.$
+    $\Delta_{g,s}\in\mathsf{K}_{g,2s}$ for all $g\leq 4$ and $s=2$,
+  • (iii)
+    $\Delta_{g,s}\neq 0$ for $(g,s)\in\{{(6,0)},(5,1),(4,2)\}\,.$
 By [8, Theorem 7], $\mathsf{K}_{5,2}$ is 1-dimensional.
-Therefore, Theorem 11 is a consequence of parts (i) and (ii) of Theorem 47 for $\Delta_{5,1}\in\mathsf{K}_{5,2}$.
-The only obstruction to cover all the cases of (66) is computational.
+Therefore, Theorem 10 is a consequence of parts (i) and (iii) of Theorem 48 for $\Delta_{5,1}\in\mathsf{K}_{5,2}$.
+Theorem 48 covers the first three cases listed in (70).
+The only obstruction to covering all the cases of (70) is computational.
 At the moment, the remaining cases require too much time and memory to compute.
-As a consequence of the non-vanishings in part (ii) of Theorem 47, we obtain the following result.
+As a consequence of the non-vanishings in part (iii) of Theorem 48, we obtain the following results.
 Corollary 49.
-The class $\operatorname{PR}_{5,1}]\in\mathsf{CH}^{5}(\mathcal{X}_{5})$ is not tautological.
+The classes $\operatorname{PR}_{5,1}]\in\mathsf{CH}^{5}(\mathcal{X}_{5})$ and $\operatorname{PR}_{4,2}]\in\mathsf{CH}^{5}(\mathcal{X}^{2}_{4})$ are not tautological.
 We have also checked the vanishings
@@
 Based on the calculations of Theorem 48, we propose the following general statement about the Gorenstein kernels of $\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$.
-Conjecture 49.
+Speculation 50.
 The Gorenstein kernels $\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})$ satisfy:
@@
     $\{\mathsf{K}_{g,n}\subset\mathsf{R}^{*}(\mathcal{M}_{g,n}^{\mathrm{ct}})\}_{g,n}$ is the smallest system of ideals which is closed under properties (i)-(v) of Section 1.11 and for which $(\mathsf{P1})$ holds.
-We now turn to the proof of Theorem 47 (i). The verification of (ii) is through $\mathsf{admcycles}$, and the corresponding code can be found in [27].
+Calculations suggest that Speculation 50 is correct at least for $\mathsf{K}_{g\leq 7,0}$.
+Pixton suggests that the first interesting case of Speculation 50 is for codimension 7 cycles in $\mathsf{K}_{8,0}$.
+6.5. Proof of Theorem 48
+The verifications of Thereom 48 (ii) and (iii) are through $\mathsf{admcycles}$, and the corresponding code can be found in [27].
+In these cases, we know the 3-spin relations are complete [8, Theorem 7, proof of Proposition 33].
+With the current computational tools, the proof that $\Delta_{4,2}\neq 0$ requires additional optimization to produce the matrix of 3-spin relations, which is explained in [8, Section 7] and was used for [8, proof of Proposition 33].
 Proof of Theorem 48 (i).
@@
 $\overline{\operatorname{aj}^{*}([\operatorname{PR}_{g,1}])}\cdot\lambda_{g}=\frac{1}{24}\tau_{*}(\operatorname{ev}_{1}^{*}(0)\cdot\operatorname{ev}_{2}^{*}(0)\cap[\overline{\mathcal{M}}_{g,{2}}(E,1)]^{\mathrm{vir}})\cdot\lambda_{g-1}=\frac{1}{12|B_{2g}|}(\psi_{1}+\psi_{2})\lambda_{g}\lambda_{g-1}\,$,
-where the last equality is a special case of [53, Proposition 6.8].
+where the last equality is a special case of [54, Proposition 6.8].
 On the other hand,